### v2

# secReferences in the Online Appendix

Abstract
We analyze the timing of NGO campaigns to shed light on NGOs’ objectives and how their strategies evolve over time. Using data from 2,500 campaigns, we find that NGOs are six times more likely to launch campaigns on their target’s Annual General Meeting (AGM) date. Although this strategy increases media exposure and stakeholder scrutiny, resulting in consumer boycotts and related shareholder proposals at the following AGM of the targeted firm, it has no impact on current AGM votes. As NGOs build reputational capital, they adjust their timing to influence AGM votes, revealing the trade-offs they face to drive corporate change.
1 Introduction
Over recent decades, non-governmental organizations (NGOs) have emerged as influential social actors, shaping regulations and corporate strategies outside traditional political channels. Through tactics of “private politics,” such as boycotts and shaming campaigns, NGOs have effectively pressured companies to reform their practices (Baron, 2001; Egorov and Harstad, 2017). Many studies examine how campaigns by both large and small NGOs impact targeted firms, often showing enduring effects (e.g., Spar and La Mure, 2003; Yaziji and Doh, 2009). For instance, just weeks before McDonald’s 2006 Annual General Meeting (AGM), Greenpeace launched a campaign branding the company ‘McAmazon‘ for alleged links to Amazon deforestation, prompting McDonald’s to adopt a new sustainable sourcing policy within days (Goldberg and Droste Yagan, 2007). Similar campaigns have targeted companies across various sectors, including Nike, Shell, Monsanto, and BNP Paribas. Despite the significance of these campaigns in shaping NGO strategies and influencing corporate behavior, we know relatively little about NGOs’ objectives, trade-offs, and how these evolve over the course of an NGO’s life.
Researchers addressing these questions face several challenges, particularly due to the scarcity of data on NGOs. For instance, donation data for non-political nonprofits is difficult to access because of donor privacy concerns. [^1] Data on donations to political parties and charities are more readily available due to disclosure rules and surveys (e.g., Cagé and Rouzet, 2015; Perez-Truglia and Cruces, 2017; Bouton et al., 2022; Guillot and Cagé, 2023). This paper differs from this literature in that it does not focus on the connection between NGO and politics, but rather on the incentives that NGOs face when deciding to campaign against a firm. NGOs are also subject to much less stringent transparency requirements than other firms such as publicly listed corporations, which complicates the tracking of their fund usage. Moreover, the absence of a quantitative framework to systematically analyze NGO objectives presents another significant challenge, as NGOs’ diverse missions cannot easily be reduced to profit maximization (Maxwell, 2012; Fioretti, 2022). To obviate to this problem, the literature emphasizes extreme behaviors of NGOs, with some contributing to the greater good (e.g., Daubanes and Rochet, 2019), while others are captured by corporate interests (e.g., Bertrand et al., 2021, 2023).
The starting point of our analysis is that NGO campaigns are strategically timed. Drawing on campaign data from a wide range of NGOs—spanning smaller entities like the ‘Physicians Committee for Responsible Medicine‘ and the ‘European Center for Digital Rights‘ to major players like ‘PETA‘ and the ‘Sierra Club‘—we identify two key factors that influence an NGO’s ability to capture stakeholders’ attention. First, media coverage amplifies NGO campaigns, making efforts like shaming and boycotting more effective. [^2] Consistent with this, Hawn (2021) finds that media coverage of a firm’s irresponsible behavior can significantly amplify negative public perceptions (Shea and Hawn, 2019), potentially resulting in the failure of planned merger. Second, NGOs with stronger reputations are better positioned to influence target firms, as they can leverage broader support from stakeholders such as workers, consumers, shareholders, suppliers, and the public. We demonstrate that these two factors create a trade-off regarding the strategic decision of campaign timing. NGOs must choose between maximizing influence on firms and enhancing their reputation.
Employing over ten years of campaign data against the largest US corporations, we find that NGOs can improve their reputation by launching campaigns near the target firm’s AGM. The AGM attracts the highest media coverage, as media outlets engage with managers and key shareholders to discuss the firm’s current achievements and future outlook (Dyck et al., 2008). This heightened media attention motivates NGOs to campaign during this period, increasing their visibility and positively influencing the perception held by the public and the target firm regarding the NGO’s stance on the issue, as well as the NGO’s commitment to driving change, thereby accumulating reputation (Bar-Isaac et al., 2008; Gent et al., 2015; Mitchell and Stroup, 2017). Since our estimation routine compares the timing of the campaigns of NGOs campaigning against multiple firms in a year, we can account for any confounding factors varying within an NGO (and target firms) across years through fixed effects, such as their size.
A late campaign is unlikely to influence shareholder votes at the AGM, as most shareholders have already voted by proxy (Bebchuk et al., 2020). Nevertheless, we find that the heightened visibility of campaigns timed with the AGM increases the likelihood of (i) a shareholder proposal on a related topic the following year and (ii) a boycott campaign in the subsequent months, compared to campaigns launched at other times. Campaigns initiated one to two months before the AGM are more likely to lead to successful votes on related proposals at the current AGM, revealing a new trade-off between influence and exposure.
We argue that NGOs navigate this tradeoff differently depending on their reputation. We first show that NGOs are more likely to campaign on AGM dates early in their history but gradually shift to launching campaigns two or three months prior the AGM to impact votes directly. Our analysis rules out learning this aspect of proxy voting as an explanation for our finding, as we observe the same trend for NGOs whose first campaign occurred just after an AGM, meaning they already knew they could not influence the vote by campaigning that late. We also find no evidence that NGOs select different target firms as they gain more campaigning experience.
Second, we demonstrate that the success of a shareholder proposal at a current AGM is linked to how many times the NGO backing the proposal has campaigned on AGM dates in the past. We interpret this as a reduced-form proxy for the reputation built by the NGO for fighting over issues. As we document diminishing returns to targeting AGM dates, once an NGO has accumulated sufficient reputation, it is more effective at influencing shareholder votes through early campaigns. Conversely, we find that accumulating experience at campaigning outside AGM dates does not help NGOs at swaying shareholder proposal votes.
Empirically, our approach leverages exogenous variation in the timing of AGM meetings, which in the U.S. we show to be relatively stable within firms over time due to regulations by the Securities and Exchange Commission (SEC) designed to give shareholders adequate time to review proposals before voting them. Across analyses, our approach also allows us to exclude time-varying confounding factors at the firm or industry, and campaign topic levels. In particular, while current events and news may influence an NGO’s decision to launch a campaign, they cannot influence the AGM calendar. As a result, objectives are identified by the tradeoff between targeting a firm’s AGM date to increase their exposure and impacting the firm’s decisions by campaigning early in support of the news or event.
Although we do not structurally estimate stakeholders’ beliefs about an NGO’s type (reputation)—an approach unfeasible with campaign data alone—our analysis indicates that the success of an NGO campaign in supporting a shareholder proposal at an upcoming AGM relies on the NGO’s past activism, highlighting the importance of reputation as a signaling mechanism (Bar-Isaac et al., 2008).
We interpret our results through the lens of firm life cycle models, where consumers initially lack information about a firm’s product quality, spurring firms to invest in reputation, especially when firms are young (e.g., Wernerfelt, 1988; Cabral, 2000, 2009). In this context, waiting for the target firm’s AGM is a costly investment for NGOs, as they must weigh the immediate recognition gained from protesting as soon as discovering a project they oppose against the benefits of waiting to protest when a firm is under substantial media scrutiny at a time that is salient for shareholders and decision makers.
Over time, NGOs may recognize that their campaigns garner significant support from key stakeholders like consumers and shareholders, reflecting a better reputation. Once established, they can leverage this reputation to directly impact firms’ decisions, engage in negotiations with firms, or spend their image to directly support shareholder proposals.
Taken together, our results describe a novel trade-off between exposure and (immediate) impact as a key driver of the societal consequences of NGO activism. Unlike our paper, these activists are often assumed to pursue social objectives, which may stem from prosocial behavior as described by Bénabou and Tirole (2006) or from moral preferences, according to Baron (2010). Their level of commitment can also depend on how closely they align with the cause, as noted by Besley and Ghatak (2005). Baron (2010) further illustrates that NGOs can emerge when socially motivated individuals coordinate efforts, drawing on frameworks outlined by Dixit (2004) and Tabellini (2008). When firms become targets of activist campaigns, firms may engage in “self-regulation” to preempt opposition and reduce potential harm. As firms’ capture of regulators has become an unavoidable feature of our economies, NGO activism surges to rebalance public and private interests (Glaeser and Shleifer, 2003; Egorov and Harstad, 2017; Daubanes and Rochet, 2019). [^3] A related literature highlights several instances in which NGOs can become subject to firm capture (e.g., Grossman and Helpman, 1994, 2001; Bertrand et al., 2014; Drutman, 2015; Bertrand et al., 2021). For example, firms may strategically exploit their relationships with NGOs—through donations or partnerships—to further political agendas (Bertrand et al., 2020) or to regain consumer and investor trust after reputational damages (Akey et al., 2021).
This paper also contributes to the growing body of literature on firms’ objectives and their societal impact (Flammer, 2018; Fioretti, 2022; George et al., 2023; Fangwa et al., 2024). Firms can help address significant global challenges such as pollution and inequality by internalizing their externalities. However, the mechanisms by which firms do so without explicit public policies, like Pigouvian taxes, remain unclear, even when CSR investments increase market value (e.g., Hawn and Ioannou, 2016; Hartzmark and Sussman, 2019; Hasan et al., 2022). [^4] A recent literature also suggests that government initiatives might even hinder private ones (Akey et al., 2024). The literature emphasizes that shareholders can influence firms through either voicing concerns—especially through voting at AGMs (Cuñat et al., 2016; Gine et al., 2017)—or selling shares (Broccardo et al., 2022; Bonnefon et al., 2022; Hart et al., 2022; Saint-Jean, 2024), whereas consumers exert influence through their willingness to pay (e.g., Gneezy et al., 2010; Elfenbein and McManus, 2010). However, other agents may play a pivotal role. Among them, NGOs can serve as a coordination mechanism among stakeholders with potentially unaligned preferences and empower them to demand change (Fioretti et al., 2023; Chuah et al., 2024; Koenig and Poncet, 2019).
Therefore, this short paper sheds new light on NGOs as coordinating various stakeholders such as shareholders and consumers, empirically highlighting a set of tradeoffs that arise from stakeholder involvement (Porter and Kramer, 2011; Kaplan, 2020). We provide new insights into the incentives NGOs face and how they can help align a firm’s action with the interests of its stakeholders (McGahan and Pongeluppe, 2023). In this direction, the results also suggest that other stakeholders can gain privately (i.e., reputation) by engaging with a firm, bringing empirical evidence to theories suggesting that stakeholders would not contribute resources to an organization without the potential to claim at least some of the value created (cf., Barney, 2018; McGahan, 2023).
Finally, our findings contribute to a growing body of literature that exploits the exogenous occurrence of NGO campaigns to asses changes in the performances and prosocial outcomes of targeted firms (e.g., Koenig and Poncet, 2022; Mazet-Sonilhac and Mésonnier, 2023; Conway and Boxell, 2024). Our study turns the table by focusing on the incentives that non-profit firms face when campaigning against for-profit firms.
The paper is organized as follows. Section 2 outlines the data construction and provides summary statistics. The empirical approach is described in Section 3. Sections 4 and 5 examine media exposure and influence at the current AGM as potential drivers of AGM actions. Section 6 explores the role of NGO reputation in supporting related shareholder proposals. Section 7 discusses the results and Section 8 concludes.
2 Data and Summary Statistics
Our dataset integrates three sources. First, we use data from Sigwatch, a consulting firm tracking global NGO campaigns against corporations (Hatte and Koenig, 2020). The data includes the campaign’s public release date and cause (e.g., “GMOs in food”), the targeted company and its prominence (e.g., if mentioned in the headline), and the sentiment (varying from negative, -2, to positive, +2), which is generally negative. We focus on campaigns against US-listed firms with negative sentiment, where only one firm is mentioned, and it appears in the headline. [^5] Since our analysis links campaign timing with a firm’s AGM date, we exclude campaigns against multiple firms as their AGM dates may differ. The dataset spans 2,512 NGO campaigns launched by 731 unique NGOs and targeting 523 unique firms. The most active NGOs in our sample include well-known organizations such as the Sierra Club, PETA, and Greenpeace. The most targeted companies are Monsanto, McDonald’s, Facebook, Coca-Cola, Apple, and Exxon Mobil. The most common topics are Environment/Pollution, Human & Labor Rights, Animal Rights, and Health.
Second, we use data from Institutional Shareholders Service (ISS), a key provider of corporate voting data, on shareholder proposals at the largest (by market capitalization) 3,000 US-listed firms’ Annual General Meetings (AGMs). For those firms, ISS data report detailed information on shareholder proposals for these firms, including proposals withdrawn before the AGM.
We select environmental and social proposals (4,483 in total) from the complete set of proposals filed at US firms between 2010 and 2021 (837 firms in total). Around a third of these E&S proposals are successful: either withdrawn from the voting agenda before the AGM or passed after going through a shareholder vote. Evidence shows that shareholders agree to withdraw the proposal only if the company commits to its implementation (Mueller et al., 2023). We document two facts: the number of E&S shareholder proposals has increased by 40% between 2010 and 2021, and their success rate has increased from 35% in 2010 to almost 50% in 2021. Appendix Figure B1 illustrates these trends.
Lastly, we source firm-level news data from Ravenpack. RavenPack is a data analytics platform that transforms unstructured news content into structured data, offering sentiment analysis and event detection. The data allows us to count the news related to every firm headquartered in the US. [^6] We only retain news with a Relevance Score of at least 75%, which ensures the relevance of the news for the firm. We include only US-listed firms from 2010 to 2020. The sample includes 5,212 US firms.
We link these three datasets using standard identifiers, including year, firms’ CUSIP, and ISIN numbers. Appendix Table A1 presents summary statistics for NGO-related variables (Panel a). On average, NGOs launch 3.40 campaigns during the sample period, but the distribution is skewed, with over 50% of NGOs initiating only one campaign. Most campaigns focus on social issues rather than environmental ones. Panel (b) displays the variation in key variables for the targeted firms, including market capitalization, assets, and news coverage.
The following three sections examine the payoffs accruing to an NGO in connection with the timing of the campaigns it launched, as well as the reactions to the campaigns of central stakeholders such as consumers and shareholders. We present the key takeaways from each analysis at the end of each section. Next, we discuss the empirical approach in detail.
3 Empirical Strategy
Our empirical approach investigates the time gap between a firm’s AGM and the actions taken by NGOs or other stakeholders, while controlling for time-varying confounding factors at the firm, industry, and campaign topic levels. Importantly, our identification strategy relies on the independence of the corporate calendar from events that may motivate NGOs to launch campaigns. The timing of AGMs is largely exogenous, as SEC regulations mandate a stable schedule across years. Specifically, Rule 14a-8 ensures shareholders have ample time to review proposals before voting, requiring that shareholder proposals be filed at least 120 days before the proxy statement is released, based on the previous year’s AGM timing. This stabilizes AGM dates across consecutive years (Glac, 2014), as shown in Appendix Figure B2, which plots AGM months across adjacent years. Most variation outside the diagonal line—indicating significant fluctuations in AGM months—is attributed to the pandemic, as the SEC allowed firms to adjust AGM dates in 2020 to prevent COVID-19 spread and facilitate shareholder attendance (SEC, 2020). Our results remain robust even when excluding data from 2020 and 2021.
While current events may influence an NGO’s decision to launch a campaign against a firm, they are unlikely to affect the timing of AGMs. These external factors may shift an NGO’s objectives by altering incentives to target a firm. Our analysis leverages these shifts to identify a key trade-off for NGOs: whether to influence a firm’s strategies by launching campaigns early or to gain greater visibility by timing campaigns around the firm’s AGM.
Our results highlight NGO experience as a crucial factor in this trade-off. Before interpreting our findings, we present results in Appendix Table A1 suggesting that NGOs do not select targets based on their experience, namely the number of campaigns they launched in the past. Panel (c) reveals that both small and large NGOs—in terms of their number of campaigns in the sample—target firms with similar market capitalization and total assets. Targets also have similar numbers of yearly news mentions, environmental and social scores, and suffered a similar number of environmental and social campaigns. Panel (d) examines target selection by regressing each one of these firm-level variables against the number of past campaigns launched by the campaigning NGO, excluding (Column 1) or including (Column 2) fixed effects at the firm-by-year, NGO-by-year, and topic-by-year levels. This correlation analysis shows that NGOs’ campaign targets do not change as NGOs gain experience, although they are more likely to focus on environmental issues rather than social ones in later years due to the increasing prominence of environmental challenges over social ones.
4 Timing Campaigns for Media Exposure
We start our analysis by investigating when NGOs are most likely to launch their campaigns relative to a firm’s corporate calendar. Figure 1(a) shows that NGOs are significantly more likely to launch campaigns on a firm’s AGM date. Campaigns are about 4 to 6 times more common on AGM weeks compared to other weeks; while a typical week sees 28 campaigns, AGM weeks have more than 100 campaigns. [^7] Appendix Figure B3 confirms the result showing a significant drop in campaign density immediately after AGM dates (p-value $\leq 1\%$), following Cattaneo et al. (2020).
Figure 1: Timing of NGO campaigns and media attention around AGMs
(a) Timing of NGO campaigns
(b) Media coverage in the AGM week (gray)
Note: Panel (a) displays the number of NGO campaigns launched relative to the date of the Annual General Meeting of the targeted firm (week 0), based on 1,489 NGO campaigns. Panel (b) compares the weekly average count of news during the AGM week (gray bars) and outside of the AGM week (white bars), based on weekly observations for 5,211 US firms. The NGO data come from Sigwatch; the news data are from Ravenpack. Calendar weeks are re-centered around the firm’s AGM.
This rise could be related to the heightened media coverage during AGMs (see also Fioretti et al., 2023), as illustrated in Figure 1(b), where news coverage peaks in the 10 days around an AGM (gray bars) compared to other periods (white bars). To test if media exposure of NGO campaigns is highest around AGMs, we aggregate news by firm ($n$) and month-by-year ($m,y$) and run the following event study:
$\text{News}_{n,m,y}$ $=\sum^{\tau=6}_{\begin{subarray}{c}\tau=-6\\ \tau\neq-3\end{subarray}}\left(\beta_{\tau}\cdot\mathbf{1}_{\{m=\tau\}}\cdot\text{NGO Campaign}_{n,m,y}\right)+\mathbf{X}_{n,m,y}\mathbf{\gamma}^{\prime}+\alpha_{n,y}+\alpha_{m,y,s(n)}+u_{n,m,y}.$ (1)
First, note that we adjusted the calendar year to align with the corporate year so that the year starts six months before the AGM month (month 0) and ends six months later for each firm. Importantly, since firms are unlikely to change their AGM month over time due to SEC regulations as outlined in Section 3, moving from the calendar year to the corporate year does not affect the distribution of months by firms (i.e., months in following years are not double counted). Thus, the recalendarization is a mere relabeling of the time variation within a firm. [^8] In each year, we define months -6 and +6 as half months, so the calendar year has twelve full months.
Second, Equation (1) interacts the event-study dummy, $\mathbf{1}_{\{m=\tau\}}$, with the variable $\text{NGO Campaign}_{n,m,y}$, which is 1 if firm $n$ experienced an NGO campaign in month $m$ of year $y$, and 0 otherwise. We use $\tau=-3$ as the reference period because shareholders typically receive voting materials within 90 days of the AGM, marking a date when the AGM’s visibility in the news could increase (ISS, 2022).
The vector $\mathbf{X}_{n,m,y}$ contains the direct effects of $\mathbf{1}_{\{m=\tau\}}$ and $\text{NGO Campaign}_{n,m,y}$ on the dependent variable. Since (1)—as well as the following regressions—involves stacking multiple difference-in-differences regressions across years, we also include firm-by-year ($\alpha_{n,y}$) and industry-by-month-by-year fixed effects ($\rho_{m,y,s(n)}$) to control for yearly changes within firms as well as time-varying industry-relevant changes (e.g., policies or tariffs). Standard errors are clustered at the firm level.
Figure 2 reports the estimated coefficients, revealing that media coverage peaks during the month of the AGM. To quantify its magnitude, an NGO campaign launched in the AGM month increases media coverage by 80% compared to a campaign started 6 months earlier, on average. After the AGM, coverage returns to average levels, indicating that NGOs gain more media attention by campaigning close to the AGM date.
Result 1.
Media coverage of NGO campaigns is substantially greater during the target’s AGM.
Figure 2: Media exposure of NGO campaigns around an AGM
Note: We regress the number of news about firm $f$ in month $t$ on a dummy indicating the publication of a negative NGO report about firm $f$ in month $t$, a dummy equal to one when month $t$ is $k$ months away from the nearest AGM, where we take $k$ to vary between -6 and +6 months, the interaction between these two dummy variables, and firm and month by industry fixed effects. Month -6 and +6 are “half” months, so the year still consists of 12 full months. The figure also reports the 95% confidence interval when the standard errors are clustered at the level of the firm. The sample period goes from January 2010 to December 2019 and covers 4,676 firms. NGO campaign data from Sigwatch. AGM data from ISS. News data from Ravenpack. Industry data from Compustat. The regression also includes firm and year-by-industry fixed effects. Standard errors are clustered at the firm level.
4.1 Stakeholders’ Reactions
In this section, we explore how campaign timing affects critical stakeholders of the targeted firms, such as shareholders and consumers. Therefore, timing a campaign on an AGM date has significant implications with consequences that go beyond media coverage. We report the exhibits related to the analysis in this section in the Online Appendix.
4.1.1 Shareholders
Shareholders can be seen as primary proponent of firm’s social values (e.g., Naaraayanan et al., 2021; Baker et al., 2022). Shareholders may respond more to causes campaigned during AGMs, when media coverage is highest. We define as “related” a shareholder proposal on the same topic of the campaign in the previous year. Appendix C describes the matching procedure.
For each campaign in our sample, we look at the probability that the next AGM has a related shareholder proposal, depending on the timing of the campaign:
[…]
Figure B4 shows that only NGO campaigns timed twelve months prior to the next AGM increase the likelihood of a shareholder proposal on the same topic at the AGM. This substantial increase is as large as 12 percentage points, or 44%, above the average probability of observing a proposal. Thus, NGO campaigns from the previous AGMs likely lead to a shareholder proposal at the current AGM, potentially due to heightened media attention.
Importantly, we can rule out the possibility that NGOs partnering with shareholders drive this outcome. If that were the case, submitting a shareholder proposal followed by an early campaign would allow the partnering shareholders and NGOs to immediately gain broad support on the shareholder’s proposal, without having to wait a year to advocate for change at the next AGM.
4.1.2 Consumers
Drawing from quarterly data from Compustat, we investigate the impact of NGO campaigns on consumers as the stakeholder consuming a firm’s values (Conway and Boxell, 2024) by studying how the Sales to Asset ratio changed around the time of an NGO campaign to the AGM date through
$\frac{\text{Sales}_{n,q}}{\text{Assets}_{n,q}}$ $=\sum^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}\cdot\text{NGO Campaign}_{n}\cdot\mathbf{1}_{\{\text{distance}=\tau\}}\right)+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$ (3)
where $q$ indicates a quarter and $\alpha_{n}$ and $\alpha_{s(n),q}$ are fixed effects at either the firm level or the industry-by-quarter-by-year level, respectively. The dependent value rebases $n$’s sales by its asset to account for time variation in relative firm sizes, which is large in our sample (10 years, largest 3,000 firms). That is, we investigate potential consumer boycotts in the months following an NGO campaign. Standard errors are clustered at the firm level.
Appendix Figure B5 reports the estimated $\beta_{\tau}$ coefficients separately for two subsets of the data: for NGO campaigns (i) not planned on an AGM date (blue dots) and (ii) those planned on an AGM date (red triangles). We find that consumers do not respond to NGO campaigns not timed on an AGM day, but they do so two quarters after a campaign targeted on the AGM day, implying a loss of sales to total assets of 0.7pp, or about 5%. Similar sale losses as a result of negative ESG corporate events have been documented in the literature (Conway and Boxell, 2024; Saint-Jean, 2024).
Result 2.
NGO campaigns on an AGM date impact stakeholders. (i) Shareholders are more likely to make related proposals in future years. (ii) Consumers are more likely to boycott the firm in the following months.
5 Timing Campaigns to Influence AGMs
Early campaigns could benefit an NGO aiming to influence current AGM outcomes. Since most votes are cast by proxy before the AGM, targeting the AGM date may limit an NGO’s ability to sway a significant share of the votes. For instance, over 85% of vote-eligible shares for Microsoft’s 2023 AGM were cast by proxy before the meeting (Microsoft, 2023), suggesting that a campaign close to the AGM date may have less impact on the voting results.
We define a proposal as successful if either it is withdrawn by the shareholder before the AGM or it secures at least $50\%+1$ vote at the AGM. We first look at the influence of NGO campaigns on the success of related shareholder proposals using the following specification on the whole sample of shareholder proposals:
$\text{Success}_{i,n,m,y}=\sum^{0}_{\begin{subarray}{c}l=-6\end{subarray}}\beta_{l}\cdot\text{Campaign}_{i,n,m+l,y}+\alpha_{n,y}+\alpha_{i,y}+\alpha_{m}+u_{i,n,m,y},$ (4)
where $\text{Success}_{i,n,m,y}$ is a dummy indicating the success of proposal $i$ at firm $n$ in month $m$ in year $t$, and $\text{Campaign}_{i,n,m+l,t}$ is a dummy indicating the occurrence of an NGO campaign on the same topic as proposal $i$, $l$ months before the AGM. The reference period covers the months that occur more than six months before the AGM of year $y$—i.e., $m\in[-11,-7]$. We account for firm-by-year, and topic-by-year fixed effects, as well as seasonality with month fixed effects, thereby exploiting only variation across campaigns and topics within firm-year pairs. The standard errors are clustered by firm.
Figure 3(a) plots the estimated $\hat{\beta}_{l}$. We find that NGO campaigns happening 6 months and 1 month before the AGM positively influence the outcome of related shareholder proposals. They increase, respectively, the probability of success by 2pp (6%) and 3pp (8%) more than a campaign launched more than seven months before.
To understand why campaigns 6 months before the AGM are also somewhat successful, we estimate (4) on the subset of voted proposals, with the percentage of votes in favor as the dependent variable. Note that every proposal to be voted on at the AGM is accompanied by a vote recommendation by the management (“For”, “Against”, or “Abstain”). Institutional investors can also follow the recommendations of proxy advisors, the most popular one being ISS (Shu, 2024). Those recommendations have a critical influence on the vote decisions. We thus include management and ISS’s recommendations as covariates in (4).
Figure 3(b) shows that only NGO campaigns occurring a month or three months before the AGM increases support for related shareholder proposals. This increase is by nearly 2 percentage points. Given the generally low baseline support for these proposals (24%), this translates to a 8% increase in shareholder support when an NGO campaign addresses the same topic, on average. In contrast to Figure 3(a), there is no significant increase in support for campaigns initiated as early as six months before the AGM. This suggests that the difference between the two figures is due to shareholder proposals being withdrawn after reaching an agreement with the firm.
Figure 3: NGOs’ influence on the vote of E&S shareholder proposals
(a) Proposals accepted or withdrawn
(b) Shareholder support for proposal
Note: E&S denote environmental and social proposals. The figure plots the estimated coefficients of regressions of a dummy indicating that a shareholder proposal is withdrawn or passes if it goes to a vote (left panel, based on 911 proposals), or the percentage of votes in favor of the proposal when it goes to the vote (right panel, based on 2,078 proposals), on dummies indicating the occurrence of an NGO campaign on the same topic $k$ months before the AGM, for $k$ varying from -6 to 0. The regressions also include month, firm-by-year, and topic-by-year fixed effects. Standard errors are clustered at the firm level.
Note that the subset of shareholder proposals going to a vote is not random; an argument raised, for instance, by Flammer (2015). For this reason, we investigate all E&S proposals in Figure 3(a). Given the low support of proposals going to vote, managers may let proposals that are the least likely to succeed go to a vote. However, those votes can still be helpful in disciplining firms (Levit and Malenko, 2011) as large shareholder support to a cause may be hard to conceal for managers.
Result 3.
NGO campaigns are more effective at influencing a firm’s decisions when launched shortly before the AGM.
6 Exposure vs. Influence
Therefore, Sections 4 and 5 outline a tradeoff: late campaigns impact society by boosting an NGO’s visibility among the media and key stakeholders, while early ones are more effective at influencing a firm’s decisions. How do NGOs navigate this tradeoff?
In this section, we first explore whether NGOs, after gaining visibility by targeting AGMs in the past, switch to campaigning before the AGM. We then examine whether this “visibility capital” from previous campaigns increases the likelihood of success at the AGM for shareholder proposals related to the NGO campaign.
6.1 Campaign Timing and Accumulated NGO Visibility
Section 4 suggests that targeting a firm’s AGM increases an NGO’s visibility in the media and among stakeholders. This visibility could enhance an NGO’s reputation for commitment to a cause. In this section, we explore whether NGOs, after establishing such a reputation, begin launching campaigns earlier—before the target firm’s AGM—to improve their chances of influencing the outcome as suggested in Section 5.
We investigate this mechanism by:
$\text{Distance to AGM}_{c(i,n,g),y}$ $=\sum_{\begin{subarray}{c}\tau>1\end{subarray}}\beta_{\tau}\cdot\mathbf{1}_{\{\text{Campaign \#}_{c(\cdot),g,y}=\tau\}}+\alpha_{g,y}+\alpha_{n,y}+\alpha_{i,y}+u_{c(g,n,i),y},$ (5)
which examines $\text{Distance to AGM}_{c(i,n,g),y}$, the distance (in months) of campaign $i$ to the AGM of targeted firm $n$, as a function of the number of past campaigns by NGO $g$, using a dummy, $\text{Campaign \#}_{c,g,y}$, that is one if the number of NGO campaigns launched by NGO $g$ until year $y$ is equal to $\tau$ and zero otherwise.
Since our campaign data start in 2010, we do not know how many campaigns an NGO launched before 2010. To control for this truncation as well as time varying confounders—e.g., the size of the NGO social media network across years (Tufekci, 2017)—we include NGO-by-year fixed effects. We also include firm-by-year and topic-by-year fixed effects to account for firm and topic-specific occurrences that, varying over time, may incentivize an NGO to launch a campaign. Hence, the identification of $\beta_{\tau}$ comes from multiple campaigns $c(\cdot)$ by the same NGO $g$ within year $y$ [^9] In Appendix Figure B6, we present similar results when controlling only for firm and NGO fixed effects, so that the estimation of $\hat{\beta}_{\tau}$ come from variation across both campaigns and years. Because of the scope of this analysis, we only consider NGOs with at least two campaigns, but the results still hold in the full sample. The standard errors are clustered at the NGO level.
To interpret the estimated $\beta_{\tau}$ coefficients, consider that, on average, the first NGO campaign is exactly on the month of the AGM. Thus, Figure 4 indicates that as the NGO’s experience increases, the new campaigns are launched before the AGM date. [^10] The large negative coefficients estimated for NGOs with many campaigns—e.g., more than 9—are driven by outlier NGOs with several campaigns per year. We postpone the interpretation of this result to Section 7, which sums up and discuss the results.
Alternative mechanism—learning. This result is unlikely due to NGOs learning that campaigning on the AGM date does not influence voting decisions, as most shareholders have already voted by proxy. Appendix Figure B7 shows the same downward trend observed in Figure 4 when subsetting to NGOs whose first campaign took place within 2 weeks after the AGM, when the NGO surely knew that it could not affect votes.
Alternative mechanism—selection. NGO do not increasingly launch campaigns against firms that may be more easily influenced over time. Appendix Figures B8 and B9 exclude selection on observables. For instance, target sizes do not covary (market capitalization) with NGO experience. In addition, environmental (social) scores do not covary with an NGO’s experience in environmental (social) campaigns.
Result 4.
NGOs initially target AGM dates to gain visibility. They schedule subsequent campaigns a few months before AGM dates to influence the outcome of the votes.
Figure 4: NGO experience and the timing of NGO campaigns
Note: This figure plots the estimated coefficients from the regression of the distance between NGO campaigns and the targeted firm’s next AGM (in months) on dummies indicating the experience of the NGO, measured by the number of campaigns launched before. Regressions are based on a sample of 1,136 NGO campaigns. Standard errors are clustered at the NGO level.
6.2 NGO Success and Reputation
We investigate whether NGO reputation, as proxied by past campaigns on AGM dates, makes NGOs more effective at producing positive corporate change. We assess this channel with the following regression:
$\text{Success}_{i,n,y}$ $=\beta_{1}\text{Campaign on Same Topic}_{c(i,n,g),y}+\beta_{2}\text{Experience}_{g,y}$ (6)
$+\beta_{3}\text{Campaign on Same Topic}_{c(i,n,g),y}\cdot\text{Experience}_{g,y}$
$+\alpha_{n,y}+\alpha_{g,y}+\alpha_{i,y}+\alpha_{m(c(i,n,g))}+u_{i,n,y},$
where $\text{Success}_{i,n,y}$ is defined as in (4)—namely, either a successful vote or a withdrawn proposal. On the right-hand side, $\text{Campaign on Same Topic}_{i,n,y}$ is one if there has been a campaign by NGO $g$ against firm $n$ on topic $i$ in the ten months preceding the AGM and zero otherwise, and $\text{Experience}_{g,y}$ indicates the number of campaigns NGO $g$ had launched in the past. We control for all confounding variations at the month, firm-by-year, NGO-by-year, and topic-by-year through fixed effects. Standard errors are clustered by firm.
Table 1: NGO experience at targeting AGMs and success at supporting current shareholder proposals
Dependent Variable: Proposal Passes (0/1)
Model: (1) (2) (3) (4)
Main Variables
Campaign on Same Topic 0.008 0.076 0.032 0.070
(0.064) (0.039) (0.056) (0.041)
Experience 0.018
Campaign on Same Topic $\times$ Experience 0.026
AGM Experience (0/1) -0.042 -0.041
(0.070) (0.069)
Campaign on Same Topic $\times$ AGM Experience (0/1) 0.169∗∗ 0.150∗∗
(0.074) (0.073)
Non-AGM Experience (0/1) 0.005
Campaign on Same Topic $\times$ Non-AGM Experience (0/1) 0.065
AGM Experience -0.018
Squared AGM Experience 0.002∗∗
Campaign on Same Topic $\times$ AGM Experience 0.074∗∗
Campaign on Same Topic $\times$ Squared AGM Experience -0.005∗∗∗
Fixed-effects
Firm x Year
Topic x Year
NGO x Year
Month
Fit statistics
Observations 38,613 38,613 38,613 38,613
R2 0.32100 0.32065 0.32101 0.32260
*–p$<0.1$; **–p$<0.05$; ***–p$<0.01$.
  • Note: This table reports the estimated coefficients from the regression of a dummy indicating whether an E&S shareholder proposal is withdrawn before voting or passes after voting (Success) on a dummy indicating the occurrence of an NGO campaign on the same topic as the proposal in the 10 months preceding the AGM (Campaign on Same Topic), a dummy indicating whether the NGO has launched a campaign on a firm’s AGM day in the past (AGM Experience (0/1)), and their interaction. In Column 2, we also consider the number of campaigns the NGO has launched in the past, which were not on the firm’s AGM day (Non-AGM Experience), and its interaction with Campaign on Same Topic. In Column 3 consider the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (AGM Experience), its square (Squared AGM Experience), and their interaction with Campaign on Same Topic. All regressions also include Firm x Year, NGO x Year, and Topic × Year fixed effects. Clustered (firm) standard errors are reported between parentheses.
Column (1) of Table 1 reports the estimated $\{\beta_{1},\beta_{2},\beta_{3}\}$. Campaigning on the same topic as a shareholder proposal only weakly influences the outcome of the proposal, and the $\beta_{1}$ coefficient is not significantly different from zero. Reassuringly, the number of past campaigns affects a proposal’s success only if the NGO campaigns on a topic being discussed at the AGM ($\beta_{2}=0$ but $\beta_{3}>0$): such campaigns increase a proposal’s success by 8% on average.
Column (2) unpacks Experienceg,y by considering the timing of an NGO’s past campaigns. We replace $\text{Experience}_{g,y}$ for the variable $\text{AGM Experience}_{g,y}$, which is 1 if NGO $g$ launched at least one campaign on the day of the AGM in past years and zero otherwise. The interaction coefficient is significant and economically important. While a campaign on the same proposal topic increases its success at the AGM by 8pp (26%), the effect increases by 14pp (45%) when that NGO has a history of targeting AGMs.
[…]
Finally, as in signaling games where there are decreasing returns to signal one’s type if the agent has already signaled the type in the past, we estimate decreasing returns from launching campaigns on AGM dates. Column (4) replaces $\text{Experience}_{g,y}$ for the cumulative count of the campaigns when NGO $g$ targeted an AGM date ($\text{AGM Experience}_{g,y}$) and its squared value: the last two rows of Column (4) find that the coefficients of the interactions between those variables and Campaign on Same Topic imply a concave relationship between AGM experience and AGM success. The results are consistent with Figure 4, which shows that only the very first campaigns happen close to an AGM date.
Result 5.
Targeting AGMs helps an NGO driving support for shareholder proposals in future periods, but display decreasing returns.
7 Discussion
Our results demonstrate that NGOs strategically use the timing of their campaigns to influence firms’ decisions. In particular, as media coverage is higher during AGMs, NGOs are likely to wait for the AGM date to launch their campaigns against target firms’ practices or strategies (Result 1). By doing so, NGOs prompt reactions from key stakeholders, such as shareholders, who are more likely to file a related proposal at the following AGM, and consumers, as the likelihood of boycotts increases if NGOs campaign on the AGM date (Result 2). The goal of these campaigns is to alter a firm’s strategy by damaging its reputation and showing stakeholders’ disappointment (Bakos and Dellarocas, 2011; Kitzmueller and Shimshack, 2012).
However, campaigning before the AGM offers direct advantages. Specifically, when a shareholder proposal has already been filed at a firm, targeting that firm on the same issue increases the likelihood that the firm accepts the proposal outright—potentially leading to its withdrawal from the AGM ballot—or, at the very least, raises shareholder support for the proposal by an average of 8% (Result 3). Since shareholder proposals generally receive low support, early NGO campaigns can play a role in disciplining managers (Levit and Malenko, 2011). Importantly, these analyses control for NGO-by-year fixed effects and topic-by-year fixed effects, accounting for any time-varying confounding factors at the NGO level—such as experience or reputation—or at the cause level—such as specific news events—thus ensuring that globally renown NGOs (e.g., WWF) do not solely drive these results. Also, we provide evidence suggesting that NGOs campaigns on AGMs do not target different firms compared to other campaign times in Appendix Table A2.
Therefore, NGOs face a trade-off, which they navigate by adapting their strategies over time. NGOs with limited activism histories tend to campaign on the AGM date, while more established NGOs prefer targeting firms just before the AGM to directly influence shareholder votes (Result 4). This pattern is consistent with the idea of using campaigning time as a signaling device. One interpretation of this result is that, as NGOs gain more experience with campaigning—and potentially run multiple campaigns in a year—they accumulate exposure through standard news media and may already have a large social media following. As a result, there may be less pressure to campaign specifically on the AGM date. Another interpretation is that, as NGOs become more complex organizations over time, media exposure becomes less vital for influencing decision-making at target firms. Disentangling these interpretations suggest several potential avenues for future research, contributing to the broader conversation on civil society lobbying and activism (Bertrand et al., 2021, 2023).
The public may also expect mature NGOs not only to generate negative publicity but also to produce immediate results. Thus, there may be diminishing returns to reputation from consistently targeting AGMs. We show that this practice enables NGOs with a history of targeting AGMs to help related shareholder proposals pass. However, an excessively long history of targeting AGMs, which according to our estimates happens after 13 campaigns on average, decreases support for shareholder proposals (Result 5).
In conclusion, NGOs play a crucial role in increasing information and aligning corporate interests with societal needs, and their growth in recent years is well-documented (e.g., Egorov and Harstad, 2017; Daubanes and Rochet, 2019). However, our understanding of the incentives they face, their relationships with firms and stakeholders, and the strategies available to them remains limited due to the scant data available on NGOs and models to interpret their actions. This paper offers a detailed analysis of NGO campaigns to uncover these trade-offs and incentives by leveraging the exogeneity of the corporate calendar to NGO’s goals. Similar approaches have been successfully used to identify the preferences of other stakeholders and may help reframe the role of AGMs not just as voting venues but also as forums where firms engage with and negotiate alongside outside stakeholders in ways that are not often observed by researchers but can be inferred from the underlying system of incentives (Fioretti et al., 2023). We also acknowledge that NGOs not only campaign but also collaborate with corporations (Dahan et al., 2010). Future research could capitalize on similar exogenous variation to that employed in this paper to quantitatively study the incentives for collaboration.
By focusing on the timing of NGO campaigns, we demonstrate how this seemingly innocuous decision has significant implications for both targeted firms and the success of future initiatives. We show that, similar to for-profit organizations, NGOs are influenced by their reputation, which affects the timing of the information presented to stakeholders and, consequently, their responses. Thus, systematically studying NGOs’ objectives and how they evolve over time is essential for understanding whether and how their activism can help firms internalize firms’ externalities.
8 Conclusion
This paper examines the timing and motives behind NGO campaigns, revealing that these campaigns are not randomly timed but often strategically target firms’ Annual General Meetings. While this timing does not affect votes on environmental and social proposals, which are cast beforehand, it enhances the NGO’s visibility, leading to greater consumer reactions and increasing the likelihood of related shareholder proposals in the following year. Mature NGOs utilize the reputation similarly built over the years to influence current firm decisions by supporting relevant shareholder proposals. Thus, this paper contributes by empirically showing that an NGO’s objectives vary throughout its life, indicating that NGOs pursue different outcomes for targeted firms, their stakeholders, and society at large as they gain visibility, thereby fostering diverse discussions within companies and, potentially, in the broader public arena.
Appendix A Omitted Tables
Table A1: Summary Statistics
Panel a: Variables varying by NGO
Mean 25% 50% 75% 95%
Number of campaigns by NGO 3.20 1 1 3 13
      – E campaigns 1.51 0 1 1 5
      – S campaigns 1.70 0 1 1 6
      – On AGM date 0.24 0 0 0 1
Campaign sentiment -1.72 -2 -2 -1 -1
Panel b: Variables varying by target firm
Mean 25% 50% 75% SD
Market Cap (USD bn) 175.66 21.71 75.05 209.47 263.66
Total Assets (USD bn) 98.81 21.80 59.27 138.68 124.36
Yearly Sales (USD bn) 37.26 6.85 17.35 42.41 150.58
Yearly Number of News 1,183 328 616 1,046 1,695.25
E score 5.66 4 5.58 7.3 2.22
S score 4.44 3.30 4.70 5.60 1.58
E campaign 0.47 0 0 1 0.50
S campaign 0.53 0 1 1 0.50
Panel c: Variables varying by target firm and NGO experience
High Experience Low Experience Test difference
Mean S.E. Mean S.E. p-value
Market Cap (decile) 9.23 0.10 9.07 0.04 0.13
Total Assets (decile) 9.26 0.08 9.23 0.04 0.72
Sales (decile) 9.31 0.09 9.14 0.04 0.07
Yearly Number of News (lag) 1,132.37 114.74 1,368.67 47.67 0.06
E score 5.64 0.12 5.72 0.06 0.58
S score 4.44 0.09 4.45 0.04 0.89
E campaign 0.46 0.03 0.5 0.01 0.15
S campaign 0.54 0.03 0.49 0.01 0.12
Panel d: OLS Regression of $y_{ny}=\beta_{0}+\beta_{1}\cdot\textit{NGO Experience}_{g(n)y}+\varepsilon_{ny}$
$\hat{\beta}_{1}$ (SE) w/o FEs $\hat{\beta}_{1}$ (SE) w/ FEs N
Dependent Variable ($y_{ny}$): (1) (2) (3) (4)
Market Cap Decile -0.002 (0.003) 0.003 (0.016) 1,338
Total Assets Decile 0.003 (0.004) -0.002 (0.012) 1,338
Sales Decile 0.001 (0.004) -0.002 (0.03) 1,337
Yearly Number of News 0.996 (12.4) -4.12 (16.5) 1,323
E score 0.004 (0.009) 0.041 (0.041) 1,338
S score -0.0006 (0.006) 0.019 (0.032) 1,338
E campaign 0.006∗∗∗ (0.002) 0.021∗∗ (0.010) 1,338
S campaign -0.006∗∗∗ (0.002) -0.018 (0.012) 1,338
  • Note: This table reports summary statistics on NGO campaigns (Panel a, NGO campaign data from Sigwatch and AGM dates from ISS), firms targeted by the campaigns (Panel b, accounting data from Compustat, ESG scores from MSCI, and news count based on RavenPack data), as well as an analysis of the relationship between NGO experience and targeted firms’ characteristics (Panels c and d). In Panel c, the Market Cap, Total Assets, and Sales deciles are based on the universe of US firms in Compustat on the quarter of the campaign. We use deciles instead of dollar values because highly ranked firms saw a large expansion in market cap and assets over the last years. Ranking measures like deciles avoid such a mechanical correlation with NGO Experience, which, being defined as the number of campaigns launched by NGO g in previous years, also grows over time. In Panel d, the fixed-effects are: NGO×$\times$×Year, Firm×$\times$×Year and Topic×$\times$×Year, except for the last two rows where Topic×$\times$×Year fixed effects are not included because of perfect collinearity with the dependent variables. Standard errors are clustered by firm.
Table A2: Campaign timing and characteristics of target firms
Characteristics of Target Firm
Market Cap Assets Sales Yearly Environmental Social Campaign
Dependent Variables: (decile) (decile) (decile) News Score Score Sentiment
Model: (1) (2) (3) (4) (5) (6) (7)
Variables
Campaign on AGM day (0/1) 0.469 0.495 0.374 83.2 0.469 -0.522 0.022
(0.628) (0.695) (0.550) (185.7) (0.919) (0.715) (0.383)
Fixed-effects
Industry-Year
NGO-Year
Topic-Year
Fit statistics
Observations 1,338 1,338 1,337 1,299 1,338 1,338 1,338
R2 0.93364 0.92669 0.90388 0.99232 0.95420 0.92167 0.84525
*–p$<0.1$; **–p$<0.05$; ***–p$<0.01$.
  • Note: This table reports the estimated coefficients from regressions of different target firm characteristics on a dummy indicating whether the campaign took place on the target firm’s AGM day. All regressions also include Industry-by-Year, NGO-by-Year, and Topic-by-Year fixed effects. The Market Cap, Assets, and Sales deciles are based on the entire universe of US firms in Compustat in the quarter of the campaign. Yearly News refers to the four-quarter moving average of news mentioning the target firm. Accounting data from Compustat. NGO campaign (including Sentiment) data from Sigwatch. News data from RavenPack. Environmental and Social scores are from MSCI. Standard errors are clustered by firm.
Appendix B Omitted Figures
Figure B1: E&S shareholder proposals, and their success
Note: This figure plots the number of E&S proposals at Russell 3,000 firms from 2006 to 2021 (solid line) and the number of successful E&S proposals (blue area) in the same period. A successful proposal is defined as a proposal that is either withdrawn before the AGM, or that goes to vote and passes. Data from ISS. Based on 5,937 shareholder proposals filed at 979 firms.
Figure B2: Persistence of the AGM month over time
Note: This figure plots the percentage of AGMs happening in each month of the year (y-axis), given the month of the previous year’s AGM (x-axis). A larger and darker circle indicates a greater proportion of meetings. Plot based on 38,471 meetings of 5,745 US companies, from 2010 to 2020.
Figure B3: Testing NGO campaign timing around the AGM date
Note: This figure plots the density test for the occurrence of NGO campaigns around firms’ AGM, following Cattaneo et al. (2020). The test is based on 521 NGO campaigns launched within 50 days of the firm’s closest AGM. The NGO data come from Sigwatch; the AGM dates are from ISS.
Figure B4: NGO campaigns and shareholder proposals
Note: This figure plots the estimated coefficients of the regression of a dummy indicating the occurrence of a shareholder proposal on month $t$ at firm $f$, relative to an NGO campaign targeting firm $f$ on the same topic, $k$ months ahead. The regression also includes month-by-year-by-industry-by-topic fixed effects. The regression is based on 1,227 NGO campaigns. Standard errors are clustered at the firm level.
Figure B5: NGO campaigns and consumer reactions
Note: This figure plots the estimated coefficients of the regression of sales over total assets, on a set of time dummies interacted with a dummy indicating the occurrence of an NGO campaign targeting firm $n$. The regressions also include firm and quarter-by-industry fixed effects. The number of observations is 351,351 and 10,188, respectively. Standard errors are clustered at the firm level.
Figure B6: NGO experience and the timing of NGO campaigns – Alternative set of Fixed Effects
Note: This figure plots the estimated coefficients from the regressions of the distance between NGO campaigns and the targeted firm’s closest AGM (in months) on dummies indicating the experience of the NGO, measured by the number of campaigns launched before. The regression includes replaces NGO-by-year fixed effects with separate fixed effects for each NGO and each year. Hence the coefficients $\{\hat{\beta}_{\tau}\}_{\tau}$ in (5) is also identified from variation in NGO experience over years. Standard errors are clustered at the NGO level.
Figure B7: Campaign timing for NGOs whose first campaign is after the AGM
Note: This figure plots the estimated coefficients from the regressions (5) of the distance between NGO campaigns and the targeted firm’s closest AGM (in months) on dummies indicating the experience of the NGO, measured by the number of campaigns launched before. The panel focuses on NGOs whose first campaign was within the two weeks following the AGM. Standard errors are clustered at the NGO level.
Figure B8: NGO experience and target firm’s size
Note: This figure plots the estimated coefficients from regressing the target firm’s market cap (in deciles to control for the fact that the tails of the market cap distribution have thickened in recent years) on dummies indicating the experience of the NGO, measured by the number of campaigns it had launched in previous year as in (5). Standard errors are clustered at the firm level. The number of observations is 1,338.
Figure B9: NGO experience and target firm’s E&S characteristics
(a) Target firm’s E score for Environmental campaigns
(b) Target firm’s S score for Social campaigns
Note: This figure plots the estimated coefficients from regressing either a target firm’s environmental score (top panel) or its social score (bottom panel) on dummies indicating the experience of the NGO, measured by either the number of environmental (top panel) or social (bottom panel) campaigns it had launched in previous year as in (5). Standard errors are clustered at the firm level. The number of observations is 636 and 698, respectively.
Appendix C Matching Proposals And NGO Campaigns
This section outlines the process of matching each shareholder proposal in the ISS dataset with a relevant NGO campaign topic.
Sigwatch assigns a“Generic Topic” to each campaign in their dataset. To streamline the analysis, we combine certain topics when they are difficult to distinguish or when one category has few campaigns and can reasonably fit within a broader category. Specifically, we merge “Labor Rights” with “Human & Labor Rights,” “Children” with ”Human & Labor Rights,” “CSR / ESG” with “CSR,” and “Biodiversity” and “Wildlife” with ”Environment.” For campaigns under the “Safety” topic, which includes only 12 entries, we examine each campaign’s specific purpose and classify it as either “Product Safety” or “Food Safety.”
Using these consolidated “Generic Topic Names”, we then categorize each shareholder proposal based on its short description. When a shareholder proposal is submitted, it includes a brief, one-sentence description, which is provided in the ISS dataset and serves as the foundation for topic classification. We start by assigning proposals to topics based on keywords in their descriptions and then manually classify any remaining proposals.
Table C1: Proposal Topics
NGO Campaign Category Example of Shareholder Proposals
Topic Classification (E or S)
Climate Change E Encourage and Support Carbon Tax to Address Climate Change
Environment E Stop financing mountain top removal coal mining
Pollution E Report on Efforts to Reduce Pollution From Products and Operations
Rainforest E Report on impact on rainforest
Sustainability E Establish board committee on sustainability
Waste Generation E Do not increase radioactive waste production
Water E Report on supply chain water impacts
Animal Rights & Welfare S Review suppliers’ animal slaughter methods
CSR S Disclose Charitable Contributions
Food Safety S Label Genetically Modified Ingredients
Health S Reduce Tobacco Harm to Health
Human & Labor Rights S End Sri Lanka Trade Partnerships Until Human Rights Improve
Indigenous People S Adopt Global Policy Regarding the Rights of Indigenous People
Intellectual Property Rights S Implement a Process to Fully Vet Whether a
Supervisor is a Bona Fide Inventor Before a Patent is Filed
Politics S Affirm Political Nonpartisanship
Product Safety S Report on Nanomaterial Product Safety
Worker Safety S Report on Reducing Risks of Accidents
  • Note: This Table reports the 17 topics assigned to each shareholder proposal in the sample. Shareholder proposals data from ISS. The topics are from Sigwatch’s categorization of campaigns.