### v4

# secReferences in the Online Appendix

Abstract
We analyze the timing of NGO campaigns to shed light on NGOs’ strategies and how they evolve over time. Data from 2,500 campaigns show that NGOs are six times more likely to launch campaigns on their target’s Annual General Meeting (AGM) date. Although this strategy increases media exposure and stakeholder scrutiny, resulting in consumer backlash and related shareholder proposals at the following AGM of the targeted firm, it has no impact on current AGM votes. As NGOs build reputational capital, they adjust their timing to influence AGM votes, revealing the trade-offs they face to drive corporate change.
JEL classifications: L21, L31
Keywords: Non-Governmental Organizations, campaigns, social and environmental responsibility, ESG, firm dynamics, non-profit objectives
1 Introduction
Over recent decades, non-governmental organizations (NGOs) have emerged as influential actors shaping regulations and corporate strategies outside traditional political channels (Baron, 2001; Egorov and Harstad, 2017). However, their actions are difficult to observe, and evidence on their strategic incentives remains limited (Yaziji and Doh, 2009). As a result, NGOs are often viewed either as principled actors promoting social good (Daubanes and Rochet, 2019) or as dysfunctional entities vulnerable to corporate capture (Bertrand et al., 2020).
This paper provides new insights into how NGO campaigns coordinate stakeholder actions and ultimately shape the creation of corporate social value. We focus on campaign timing—one of the few observable aspects of NGO activity, which remains far more opaque than that of for-profit firms. [^1] A parallel literature uses administrative data to study charitable donations to NGOs (e.g., Cagé et al., 2023; Cagé and Guillot, 2025), offering valuable insights into their funding base. Our focus is instead on NGOs’ strategic choices—an important aspect of their behavior that remains relatively underexplored. While timing may seem like a minor feature, our approach reveals a key trade-off NGOs face: gaining media exposure, which can be leveraged in future campaigns, versus exerting direct influence on firm behavior. We identify the main determinants of how NGOs resolve this trade-off and show that their choices shape which stakeholders are mobilized and how firms ultimately respond.
Our findings reveal a “life cycle” in NGO strategies (Wernerfelt, 1988; Cabral, 2000). Early-stage NGOs prioritize visibility by campaigning on a target firm’s Annual General Meeting (AGM) date, where executives face shareholder scrutiny and heightened media attention, triggering consumer backlash and prompting future shareholder proposals. However, these high-visibility campaigns rarely improve firms’ environmental and social outcomes directly. Over time, NGOs shift to campaigning before a target’s AGM date to exert direct influence, increasing the success rate of current shareholder proposals and leading to measurable improvements in corporate behavior.
Consistent with reputation-building models, shareholder proposals are more likely to succeed when accompanied by NGO campaigns on the same issue—especially when the NGO has a history of targeting AGM dates (Bar-Isaac and Tadelis, 2008). However, this effect diminishes with the number of past AGM-date campaigns, suggesting decreasing returns. In contrast, past non-AGM campaigns have no influence on proposal outcomes. NGOs thus strategically use AGM-date campaigns to signal toughness and build their reputation, initially benefiting from substantial media attention but paying the cost of limited direct influence on AGM voting (Kreps and Wilson, 1982; Milgrom and Roberts, 1982). As a result, more mature NGOs shift toward earlier, pre-AGM campaigns, aiming to build credibility through constructive stakeholder engagement (Baur and Palazzo, 2011; Mitchell and Stroup, 2017).
Our identification strategy leverages the timing of firms’ AGMs, which is exogenously fixed by SEC regulations and stable within firms over time. This enables us to analyze how NGOs respond to predictable opportunities for visibility, holding firm strategies constant. Our main endogenous variable, therefore, is NGO campaign timing, which we explain through NGO experience: more experienced NGOs campaign earlier, trading visibility for impact. While unobservable factors might also influence timing decisions, we verify that key characteristics of the target firms—such as size and ESG scores—do not systematically correlate with NGO experience. This reduces concerns about selection bias, for instance, ruling out that less experienced NGOs strategically target smaller or less prosocial firms for easier gains.
By comparing NGO campaign outcomes across firms and issues within years, our analysis accounts for factors that vary across years within NGOs, firms, and issues, such as organizational competence, learning, moral authority, resources, and the salience of specific issues. NGOs also engage firms through partnerships and other channels, but since such activities are voluntarily disclosed—and thus subject to selection biases (Burger and Owens, 2010)—we rather focus on campaign timing, which must be public to be effective and, thus, observable to researchers.
Our findings open new avenues for research, as future work may empirically explore how NGOs choose between campaigning and other activist tactics or whether these different approaches act as complements or substitutes. The exogenous timing of AGMs provides a useful identification strategy for such analyses, as it holds the target firm’s strategic choices constant, allowing researchers to isolate NGOs’ choices and their incentives.
Our results reveal a novel tradeoff between exposure and direct impact as a key driver of the societal consequences of NGO activism (Briscoe and Gupta, 2016). Unlike our framework, much of the existing literature assumes that activists primarily pursue social objectives, either stemming from prosocial behavior (Besley and Ghatak, 2005; Bénabou and Tirole, 2006) or moral preferences (Baron, 2010), without discussing their incentives.
At the same time, several studies highlight instances in which NGOs themselves are subject to firm capture (e.g., Grossman and Helpman, 1994, 2001; Bertrand et al., 2014; Drutman, 2015; Bertrand et al., 2021). Firms may also strategically leverage their relationships with NGOs—through donations or partnerships—to advance political agendas (Bertrand et al., 2020), or to restore consumer and investor trust following reputational damage (Akey et al., 2021).
Our main contribution in this literature is to show that NGOs are not passive actors. Rather, their tactics evolve over time, shaping firms and their stakeholders’ responses in different ways depending on the strategies they employ (e.g., Hendricks and Sorensen, 2009; Cabral and Hortacsu, 2010). Thus, understanding NGOs’ incentives is essential to learn why firms internalize their externalities.
We also contribute to the growing body of literature on firms’ objectives and their societal impact (Flammer, 2018; Fioretti, 2022; Fangwa et al., 2024). Firms can help address significant global challenges such as pollution and inequality by internalizing their externalities. However, the mechanisms by which firms do so without explicit public policies, like Pigouvian taxes, remain unclear (e.g., Hawn and Ioannou, 2016). The literature emphasizes that shareholders can influence firms through either voicing concerns—especially through voting at AGMs (Cuñat et al., 2016)—or selling shares (Broccardo et al., 2022; Bonnefon et al., 2022; Hart et al., 2022; Saint-Jean, 2024), whereas consumers exert influence through their willingness to pay (e.g., Gneezy et al., 2010). However, other agents may play a pivotal role. Among them, we highlight that NGOs can serve as a coordination mechanism among stakeholders with potentially unaligned preferences and empower them to demand change (Koenig and Poncet, 2019; Fioretti et al., 2023; Chuah et al., 2024; Mazet-Sonilhac and Mésonnier, 2023; McGahan, 2023).
The paper is organized as follows. Section 2 outlines the data construction and provides summary statistics. The empirical approach is described in Section 3. Sections 4 and 5 examine media exposure and influence at the current AGM as potential drivers of AGM actions. Section 6 explores the role of NGO reputation in supporting related shareholder proposals. Section 7 discusses the results and concludes.
2 Data
Our dataset integrates three sources. First, we use data from Sigwatch, a consulting firm tracking global NGO campaigns against corporations (Hatte and Koenig, 2020). The data includes the campaign’s public release date and cause (e.g., “GMOs in food”), the targeted company and its prominence (e.g., if mentioned in the headline), and the sentiment (varying from negative, -2, to positive, +2), which is generally negative. We focus on campaigns against US-listed firms with negative sentiment, where only one firm is mentioned, and it appears in the headline. [^2] Since our analysis links campaign timing with a firm’s AGM date, we exclude campaigns against multiple firms as their AGM dates may differ. The dataset spans 2,512 NGO campaigns launched by 731 unique NGOs and targeting 523 unique firms. The most active NGOs in our sample include well-known organizations such as the Sierra Club, PETA, and Greenpeace. The most targeted companies are Monsanto, McDonald’s, Facebook, Coca-Cola, Apple, and Exxon Mobil. The most common causes are “environment/pollution,” “human & labor rights,” “animal rights,” and “health.”
Second, we use data from Institutional Shareholders Service (ISS), a key provider of corporate voting data, on shareholder proposals at the Annual General Meetings (AGMs) of the 3,000 largest US-listed firms by market capitalization. For those firms, ISS data report detailed information on shareholder proposals for these firms, including proposals withdrawn before the AGM.
We select environmental and social (E&S) proposals (4,483 in total) from the complete set of proposals filed at US firms between 2010 and 2021 (837 firms in total). Around a third of these E&S proposals are successful: either withdrawn from the voting agenda before the AGM or passed after going through a shareholder vote. [^3] Evidence shows that shareholders agree to withdraw the proposal only if the company commits to its implementation (Mueller et al., 2023). We document two facts: the number of E&S shareholder proposals has increased by 40% between 2010 and 2021, and their success rate has increased from 35% in 2010 to almost 50% in 2021. Appendix Figure B1 illustrates these trends.
Lastly, we source firm-level news data from Ravenpack. RavenPack is a data analytics platform that transforms unstructured news content into structured data, offering sentiment analysis and event detection. The data allows us to count the news related to every firm headquartered in the US. [^4] We only retain news with a Relevance Score of at least 75%, which ensures the relevance of the news for the firm. We include only US-listed firms from 2010 to 2020. The sample includes 5,212 US firms.
We link these three datasets using standard identifiers, including year, firms’ CUSIP, and ISIN numbers. Appendix Table A1 presents summary statistics for NGO-related variables (Panel a). On average, NGOs launch 3.40 campaigns during the sample period, but the distribution is skewed, with over 50% of NGOs initiating only one campaign. Most campaigns focus on social issues rather than environmental ones. Panel (b) displays the variation in key variables for the targeted firms, including market capitalization, assets, and news coverage.
The following three sections examine an NGO’s payoffs in connection with the timing of its campaigns, and the reactions of central stakeholders. We present the key takeaways from each analysis at the end of each section. Before that, we detail our empirical approach.
3 AGM Dates
Our empirical approach examines the timing of NGO campaigns relative to AGMs, leveraging the exogeneity of AGM dates, which remain stable within firms across years due to SEC regulations. Rule 14a-8 mandates shareholder proposals be filed 120 days before the proxy statement, anchoring AGM dates across years (Glac, 2014). Appendix Figure B2 confirms this stability, with exceptions largely due to pandemic-related adjustments allowed by the SEC (2020) to facilitate shareholder attendance at AGMs. [^5] Results remain robust excluding the covid period (2020–2021). While external events may influence NGOs’ campaign decisions, they do not affect AGM timing. Thus, AGMs allow us to isolate NGOs’ strategic trade-off from firms’ responses.
Target selection. Our analysis focuses on campaign timing, abstracting from NGOs’ selection of a target, which we leave for future research. However, Appendix Table A1 shows that NGOs’ targets does not depend on prior campaign experience, a key variable of our analysis. Panel (c) finds no differences in market capitalization, assets, news mentions, ESG scores, or campaign history between targets of small and large NGOs. Panel (d) regresses firm characteristics on NGOs’ past campaigns, showing no correlation between target selection and NGO experience. However, we see that NGOs increasingly focus on environmental over social issues in later years, reflecting the rising importance of environmental concerns.
4 Timing Campaigns for Media Exposure
NGOs are substantially more likely to launch their campaigns around AGM dates, on average. According to Figure 1(a), NGO campaigns are about 4 to 6 times more common on AGM weeks compared to other weeks; while a typical week sees 28 campaigns, AGM weeks have more than 100.666Appendix Figure B3 confirms this finding showing a significant drop in campaign density immediately after AGM dates (p-value $\leq 1\%$) (Cattaneo et al., 2020).
Figure 1: Timing of NGO campaigns and media attention around AGMs
(a) Campaigns peak during AGMs
(b) Media coverage rises during AGMs
Note: Panel (a) displays the number of NGO campaigns launched relative to the date of the Annual General Meeting of the targeted firm (week 0), based on 1,489 NGO campaigns. Panel (b) compares the weekly average count of news during the AGM week (gray bars) and outside of the AGM week (white bars), based on weekly observations for 5,211 US firms. The NGO data come from Sigwatch; the news data are from Ravenpack. Calendar weeks are re-centered around the firm’s AGM.
This finding could be related to the heightened media coverage during AGMs (Fioretti et al., 2023), as illustrated in Figure 1(b), where news coverage peaks in the 10 days around an AGM (gray bars) compared to other periods (white bars). To test if media exposure of NGO campaigns is highest during the AGM week, we aggregate negative E&S news by firm ($n$) and week-by-year ($w,y$) and perform the event study:
$\text{News}_{n,w,y}$ $=\sum^{\tau=5}_{\begin{subarray}{c}\tau=-5\\ \tau\neq-4\end{subarray}}\left(\beta_{\tau}\cdot\mathbf{1}{\{M_{n,w,y}=\tau\}}\,\cdot\,\text{NGO Campaign}_{n,w,y}\right)$ (1)
$+\mathbf{X}_{n,w,y}\mathbf{\gamma}^{\prime}+\alpha_{n,y}+\alpha_{w,y,s(n)}+u_{n,w,y}\,\,,$
where “$\text{News}_{n,w,y}$” is an indicator for whether firm $n$ in week $w$ of year $y$ experiences at least one negative E&S news event. The independent variables include a set of dummies “$\mathbf{1}\{M_{n,w,y}=\tau\}$” indicating the distance in months from the closest AGM. The function $M_{n,w,y}$ assigns week $w$ to each month; the indicator is 1 if the week falls in month $\tau$, where a month is made up of 4 weeks, and zero otherwise.777 The analysis aligns the calendar year with the corporate year, starting five months before the AGM month (month 0) and ending five months later, with months -5 and +5 slightly longer. Since SEC regulations make AGM months stable over time (Section 3), this relabeling does not alter the distribution of months by firms or double-count adjacent years. We interact these event-study indicators, with the variable “$\text{NGO Campaign}_{n,m,y}$,” an indicator variable equal to 1 if firm $n$ experienced an NGO campaign in week $w$ of year $y$, and 0 otherwise. We set the reference month to four months before the AGM, following SEC rules that require the AGM date to be set at least four months in advance.
Figure 2: Firms’ negative media coverage increases if NGOs campaign during AGMs
Note: The figure presents estimates from (1). The dependent variable is an indicator that a firm receives at least one negative E&S news in a week. Standard errors are clustered at the firm level.
The vector “$\mathbf{X}_{n,w,y}$” includes “$\mathbf{1}{\{M_{n,w,y}=\tau\}}$” and “$\text{NGO Campaign}_{n,m,y}$” separately, and controls for (i) the occurrence of negative news and (ii) all E&S news over the previous four weeks to account for a firm’s recent media coverage. [^8] Adding controls for the current count of E&S news to account for NGOs targeting AGMs of high-coverage firms does not qualitatively change the results. Since (1)—as well as the following regressions—involves stacking multiple difference-in-differences regressions across years, we also include firm-by-year ($\alpha_{n,y}$) and industry-by-week-by-year fixed effects ($\rho_{w,y,s(n)}$) to control for yearly changes within firms as well as time-varying industry-relevant changes (e.g., policies or tariffs). Standard errors are clustered at the firm level.
Figure 2 reports the estimated $\{\hat{\beta}_{\tau}\}_{\tau}$ coefficients. Negative news follow a bell-shaped pattern around the AGM, peaking significantly above zero in the AGM month, highlighting the benefits of targeting AGMs. [^9] In section 6.2, we show that the spike on AGM date is particularly strong for an NGO’s first few campaigns but it fades away for later ones (Appendix Figure B4). To quantify the magnitude of the AGM peak, an NGO campaign launched in the AGM week doubles the probability of a negative E&S news compared to a campaign initiated four months earlier, on average. After the AGM, coverage returns to baseline levels, indicating that NGOs attract more media attention when campaigning closer to the AGM.
Result 1.
Media coverage of NGO campaigns is substantially greater during the target’s AGM.
Importantly, we rule out the hypothesis that NGOs select more negative causes when targeting AGMs. Without an adequate ranking of causes by negativity, we estimate (1) with firm $n$’s average E&S news sentiment in week $w$ as the dependent variable. [^10] RavenPack scores the tone of a firm’s news coverage between -1 (adverse) and 1 (favorable). Examples of positive and negative news can be found in Appendix Figure B5. The estimated coefficients show that campaigns launched during AGM months are no different than other campaigns (Appendix Figure B6).
4.1 Stakeholders’ Reactions
This section explores how campaign timing affects critical stakeholders of a targeted firm, such as shareholders and consumers, while Section 4.2 examines its ramifications for the target’s operations. The relevant exhibits are in the Online Appendix.
4.1.1 Shareholders
Shareholders can be seen as primary proponent of firm’s social values (e.g., Naaraayanan et al., 2021; Baker et al., 2022). Shareholders may respond more to causes campaigned during AGMs, when media coverage is highest. We define as “related” a shareholder proposal on the same cause of the campaign in the previous year. Appendix C describes the matching procedure.
For each campaign in our sample, we look at the probability that the next AGM has a related shareholder proposal, depending on the timing of the campaign:
[…]
$+\alpha_{i,s(n),g,y}+u_{i,n,y}$
[…]
Figure B7 shows that only campaigns launched twelve months before the next AGM increase the likelihood of a shareholder proposal on the same issue. The effect is substantial—an increase of 12 percentage points, or 44% relative to the baseline. Since SEC 14a8(e)(2) rule prevents shareholders from submitting proposals within 120 days of the proxy statement’s release, the x-axis is capped at four months before the previous year’s AGM. This timing suggests that campaigns shortly after an AGM may spur proposals at the next one, possibly due to heightened media attention.
Importantly, we can rule out the possibility that NGOs partnering with shareholders drive this outcome. If that were the case, submitting a shareholder proposal followed by an early campaign would allow the partnering shareholders and NGOs to immediately gain broad support on the shareholder’s proposal, without having to wait a year to advocate for change at the next AGM.
4.1.2 Consumers
Drawing from quarterly data from Compustat, we investigate the responses of consumers to the timing of NGO campaigns by studying the relation between the target’s “Sales to Asset ratio” and the distance between the campaign and the target’s AGM through
$\frac{\text{Sales}_{n,q}}{\text{Assets}_{n,q-1}}$ $=\sum^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}\cdot\text{NGO Campaign}_{n}\cdot\mathbf{1}{\{\text{distance}=\tau\}}\right)+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$ (3)
where $q$ indicates a quarter and $\alpha_{n}$ and $\alpha_{s(n),q}$ are fixed effects at either the firm level or the industry-by-quarter-by-year level, respectively. The dependent value normalizes $n$’s sales by its asset to account for time variation in relative firm sizes, which is large in our ten-year sample. Standard errors are clustered at the firm level.
Appendix Figure B8 shows the estimated ${\hat{\beta}\tau}\tau$ coefficients separately for NGO campaigns (i) planned on an AGM date (Panel a) and (ii) not planned on an AGM date (Panel b). We find no consumer response to the latter, but sales drop by 0.7pp (about 5%) for the former—a magnitude consistent with Conway and Boxell (2024); Saint-Jean (2024). In Panel (c), the sales gap between AGM and non-AGM campaigns emerges one quarter before the event and continues to widen, possibly due to the greater media coverage of AGM dates.
4.2 Does AGM-Targeting Create Positive Change?
If public shaming were an effective strategy to influence firm behavior, targeting the AGM month should yield the largest improvements in E&S scores. However, Appendix Figure B9 shows that only campaigns launched 3 to 6 months before the AGM lead to subsequent increases in firms’ E&S scores—rising by 2–3% over the following year. By contrast, campaigns launched at other times, including during the AGM month, have no economically or statistically significant effect.
Although ESG scores are an imperfect measure of a firm’s social impact (e.g., Kotsantonis and Serafeim, 2019; Allcott et al., 2022), this finding suggests that NGOs’ preference for targeting AGM dates may be driven by factors beyond their direct influence on firms. Therefore,
Result 2.
NGO campaigns on an AGM date impact stakeholders but do not lead to improvements in E&S scores. (i) Shareholders are more likely to make related proposals the following year. (ii) The target’s sales drop in the months after a campaign.
5 Timing Campaigns to Influence AGM Votes
Early campaigns could benefit an NGO aiming to influence current AGM outcomes. Since most votes are cast by proxy before the AGM, targeting the AGM date may limit an NGO’s ability to sway a significant share of the votes. For instance, over 85% of vote-eligible shares for Microsoft’s 2023 AGM were cast by proxy before the meeting, indicating that campaigning around AGM dates may have little impact on shareholders’ votes (Microsoft, 2023).
We define a proposal as successful if either it is withdrawn by the shareholder before the AGM or it secures at least $50\%+1$ vote at the AGM. We first look at the influence of NGO campaigns on the success of related shareholder proposals using the following specification on the whole sample of shareholder proposals:
$\text{Success}_{i,n,m,y}=\sum^{0}_{\begin{subarray}{c}l=-6\end{subarray}}\beta_{l}\cdot\text{Campaign}_{i,n,m+l,y}+\alpha_{n,y}+\alpha_{i,y}+\alpha_{m}+u_{i,n,m,y},$ (4)
where “$\text{Success}_{i,n,m,y}$” is an indicator for the success of proposal $i$ at firm $n$ in month $m$ in year $t$, and “$\text{Campaign}_{i,n,m+l,t}$” is an indicator for the occurrence of an NGO campaign on the same cause as proposal $i$, $l$ months before the AGM. The reference period covers the months that are more than six months before the AGM of year $y$—i.e., $m\in[-11,-7]$. We account for firm-by-year, and cause-by-year fixed effects, as well as seasonality with month fixed effects, thereby exploiting only variation across campaigns and causes within firm-year pairs. The standard errors are clustered by firm.
Figure 3: NGOs’ campaigns before the AGM help E&S shareholder proposals succeed
(a) Proposals accepted or withdrawn
(b) Vote share in favor of proposal
Note: E&S denote environmental and social proposals, respectively. The figure plots the estimated coefficients from (4). The dependent variables are an indicator variable that is one if a shareholder proposal is withdrawn or successful if voted (911 proposals) in Panel (a), and the percentage of votes in favor of a proposal (2,078 proposals) in Panel (b). Standard errors are clustered at the firm level.
Figure 3(a) plots the estimated $\{\hat{\beta}_{l}\}_{l}$. We find that NGO campaigns happening 6 months and 1 month before the AGM positively influence the outcome of related shareholder proposals. They increase, respectively, the probability of success by 2pp (6%) and 3pp (8%) more than a campaign launched more six months earlier.
To understand why campaigns 6 months before the AGM are also somewhat successful, we estimate (4) on the subset of voted proposals, with the percentage of votes in favor as the dependent variable. Note that every proposal to be voted on at the AGM is accompanied by a vote recommendation by the management (“For”, “Against”, or “Abstain”). Institutional investors often follow proxy advisor recommendations, particularly from ISS (Shu, 2024). We include management and ISS recommendations as covariates in (4).
Figure 3(b) shows that only NGO campaigns launched one or three months before the AGM increase support for related shareholder proposals, with an increase of nearly 2 percentage points. Given the generally low baseline support for these proposals (24%), this corresponds to an 8% increase in shareholder support when an NGO campaign addresses the same cause, on average.
In contrast to Figure 3(a), Figure 3(b) detects no significant increase in support for campaigns initiated as early as six (or five) months before the AGM. This suggests that the difference between the two figures arises from shareholder proposals being withdrawn after reaching an agreement with the firm. As a result of these early campaigns, the target firm may take actions to resolve the issue highlighted in both the campaign and the shareholder proposal, leading to an improvement in its E&S score, which supports the finding in Section 4.2 that only early campaigns result in better E&S scores.
Moreover, the fact that campaigns launched one to three months before an AGM help secure more votes aligns with the timing of proxy voting. Shareholders typically receive voting materials within 90 days of the AGM, a period when AGM visibility in the news may increase as well as the salience of the campaign for shareholders (ISS, 2022).
Note that the subset of shareholder proposals going to a vote is not random; an argument raised, for instance, by Flammer (2015). For this reason, we investigate all E&S proposals in Figure 3(a), which also include withdrawn proposals. Given the low support of proposals going to vote, managers may let proposals that are the least likely to succeed go to a vote. However, those votes can still be helpful in disciplining firms as large shareholder support to a cause may be hard to conceal for managers (Levit and Malenko, 2011).
Result 3.
NGO campaigns are more effective at influencing a firm’s decisions when launched before the AGM.
6 Exposure vs. Influence
Sections 4 and 5 outline a tradeoff: late campaigns expose a firm’s wrongdoing to its stakeholders and boost the NGO’s visibility, while early ones are more effective at directly influencing a firm’s decisions. How do NGOs navigate this tradeoff?
In this section, we first explore whether NGOs, after gaining visibility by targeting AGMs in the past, switch to campaigning before the AGM. We then examine whether this “visibility capital” from previous campaigns increases the likelihood of success at the AGM for shareholder proposals related to the NGO campaign.
6.1 Accumulating Visibility
Targeting AGM-dates could enahnce an NGO’s visibility and its reputation for commitment to a cause. In this section, we explore whether NGOs, after accumulating some visibility, begin launching campaigns earlier—before the target firm’s AGM—to improve their chances of influencing the outcome as suggested in Section 5.
We investigate this mechanism by:
$\text{Distance to AGM}_{c(i,n,g),y}$ $=\sum_{\begin{subarray}{c}\tau>1\end{subarray}}\beta_{\tau}\cdot\mathbf{1}{\{\text{Campaign \#}_{c(\cdot),g,y}=\tau\}}$ (5)
$+\alpha_{g,y}+\alpha_{n,y}+\alpha_{i,y}+u_{c(g,n,i),y},$
which examines “$\text{Distance to AGM}_{c(i,n,g),y}$,” the distance (in months) of campaign $i$ to the AGM of targeted firm $n$, as a function of the number of past campaigns by NGO $g$, using an indicator, “$\text{Campaign \#}_{c,g,y}$,” that is one if the number of NGO campaigns launched by NGO $g$ until year $y$ is equal to $\tau$ and zero otherwise.
Since our campaign data start in 2010, we do not know how many campaigns an NGO launched before 2010. To control for this truncation as well as time varying confounders such as NGO size, we include NGO-by-year fixed effects. We also include firm-by-year and cause-by-year fixed effects to account for firm and cause-specific occurrences that, varying over time, may incentivize an NGO to launch a campaign. Hence, the identification of $\{\beta_{\tau}\}_{\tau}$ comes from multiple campaigns $c(\cdot)$ by the same NGO $g$ within year $y$ [^11] In Appendix Figure B10, we present similar results when controlling only for firm and NGO fixed effects, so that the estimation of $\hat{\beta}_{\tau}$ come from variation across both campaigns and years. Because of the scope of this analysis, we only consider NGOs with at least two campaigns, but the results still hold in the full sample. The standard errors are clustered at the NGO level.
To interpret the estimated $\{\hat{\beta}_{\tau}\}_{\tau}$ coefficients, consider that, on average, the first NGO campaign is exactly on the month of the AGM. Thus, Figure 4 indicates that as the NGO’s experience increases, the new campaigns are launched before the AGM date. [^12] The large negative coefficients estimated for NGOs with many campaigns—e.g., more than 9—are driven by outlier NGOs with several campaigns per year. We postpone the interpretation of this result to Section 7, which sums up and discusses the results.
Alternative mechanism—learning. This result is unlikely due to NGOs learning that campaigning on the AGM date does not influence voting decisions, as most shareholders have already voted by proxy. Appendix Figure B11 shows the same downward trend observed in Figure 4 when subsetting to NGOs whose first campaign took place within 2 weeks after the AGM, when the NGO surely knew that it could not affect votes.
Alternative mechanism—selection. NGO do not increasingly launch campaigns against firms that may be more easily influenced over time. Appendix Figures B12 and B13 exclude selection on observables. For instance, target sizes do not covary (market capitalization) with NGO experience. In addition, environmental (social) scores do not covary with an NGO’s experience in environmental (social) campaigns.
Result 4.
NGOs initially target AGM dates to gain visibility. They schedule subsequent campaigns a few months before AGM dates to influence the outcome of the votes.
Figure 4: More experienced NGOs campaign earlier
Note: This figure plots the estimated coefficients from (5). The dependent variable is the months between an NGO campaign and the targeted firm’s AGM date. Only firms with at least two campaigns are considered (1,136 observations). Standard errors are clustered at the NGO level.
6.2 NGO Success and Reputation
We investigate whether NGO reputation, as proxied by past campaigns on AGM dates, makes NGOs more effective at producing positive corporate change. We assess this channel with the following regression:
$\text{Success}_{i,n,y}$ $=\beta_{1}\,\text{Campaign on Same Cause}_{c(i,n,g),y}+\beta_{2}\,\text{Experience}_{g,y}$ (6)
$+\beta_{3}\,\text{Campaign on Same Cause}_{c(i,n,g),y}\cdot\text{Experience}_{g,y}$
$+\alpha_{n,y}+\alpha_{g,y}+\alpha_{i,y}+\alpha_{m(c(i,n,g))}+u_{i,n,y},$
where “$\text{Success}_{i,n,y}$” is defined as in (4)—namely, either a successful vote or a withdrawn proposal. On the right-hand side, “$\text{Campaign on Same Cause}_{i,n,y}$” is one if there has been a campaign by NGO $g$ against firm $n$ on cause $i$ in the ten months preceding the AGM and zero otherwise, and “$\text{Experience}_{g,y}$” indicates the number of campaigns NGO $g$ had launched in the past. We control for all confounding variations at the month, firm-by-year, NGO-by-year, and cause-by-year through fixed effects. Standard errors are clustered by firm.
Table 1: NGO experience at targeting AGMs and success at supporting current shareholder proposals
Dependent Variable: Proposal Passes (0/1)
Model: (1) (2) (3) (4)
Main Variables
Campaign on Same Topic 0.008 0.076 0.032 0.070
(0.064) (0.039) (0.056) (0.041)
Experience 0.018
Campaign on Same Topic $\times$ Experience 0.026
AGM Experience (0/1) -0.042 -0.041
(0.070) (0.069)
Campaign on Same Topic $\times$ AGM Experience (0/1) 0.169∗∗ 0.150∗∗
(0.074) (0.073)
Non-AGM Experience (0/1) 0.005
Campaign on Same Topic $\times$ Non-AGM Experience (0/1) 0.065
AGM Experience -0.018
Squared AGM Experience 0.002∗∗
Campaign on Same Topic $\times$ AGM Experience 0.074∗∗
Campaign on Same Topic $\times$ Squared AGM Experience -0.005∗∗∗
Fixed-effects
Firm x Year
Topic x Year
NGO x Year
Month
Fit statistics
Observations 38,613 38,613 38,613 38,613
R2 0.321 0.321 0.321 0.323
*–p$<0.1$; **–p$<0.05$; ***–p$<0.01$.
  • Note: This table reports the estimated coefficients from (6) in Column 1. In Column 1, Experience is capped at 5 to avoid outliers–as 85% of NGOs never reach that level of experience. In Column 2, we also consider the number of campaigns the NGO has launched in the past, which were not on the firm’s AGM day (Non-AGM Experience), and its interaction with Campaign on Same Topic. In Column 3 consider the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (AGM Experience), its square (Squared AGM Experience), and their interaction with Campaign on Same Topic. All regressions also include firm-by-year, NGO-by-year, topic-by-year, and month fixed effects. Clustered (firm) standard errors are reported in parentheses.
Column (1) of Table 1 reports the estimated $\{\beta_{1},\beta_{2},\beta_{3}\}$. Campaigning on the same cause as a shareholder proposal only weakly influences the outcome of the proposal, and the $\beta_{1}$ coefficient is not significantly different from zero. Reassuringly, the number of past campaigns affects a proposal’s success only if the NGO campaigns on a cause being discussed at the AGM ($\beta_{2}=0$ but $\beta_{3}>0$): such campaigns increase a proposal’s success by 8% on average.
Column (2) unpacks Experienceg,y by considering the timing of an NGO’s past campaigns. We replace “$\text{Experience}_{g,y}$” for the variable “$\text{AGM Experience}_{g,y}$,” which is 1 if NGO $g$ launched at least one campaign on the day of the AGM in past years and zero otherwise. The interaction coefficient is significant and economically important. While a campaign on the same cause as the shareholder proposal increases its success at the AGM by 8pp (26%), the effect increases by 14pp (45%) when that NGO has a history of targeting AGMs.
[…]
Finally, as in signaling games where there are decreasing returns to signal one’s type if the agent has already signaled the type in the past, we estimate decreasing returns from launching campaigns on AGM dates. Column (4) replaces “$\text{Experience}_{g,y}$” for the cumulative count of the campaigns when NGO $g$ targeted an AGM date (“$\text{AGM Experience}_{g,y}$”) and its squared value: the last two rows of Column (4) find that the coefficients of the interactions between those variables and “$\text{Campaign on Same Cause}_{i,n,g,y}$” imply a concave relationship between past AGM-targeting and proposals’ success. Therefore, the support of NGO’s to a shareholder proposal display diminishing returns. [^13] Note that the combined effect of AGM Experience and AGM Experience Squared never yields a negative probability of proposal success—even at the highest observed levels of AGM Experience—suggesting our reputation proxy does not exhibit negative returns.
This finding supports the results in Appendix Figure B4, discussed in Section 4, which demonstrate that AGM-targeting increases the probability that target receives negative media coverage only if the NGO has little campaign experience (i.e., having run less than four campaigns). Later campaigns do not generate additional awareness on AGM dates compared to other dates. In fact, “mature NGOs” seem to be better off campaigning one or two months earlier, indicating diminishing returns from campaigning on AGM dates, even in terms of visibility.
We conclude that,
Result 5.
Targeting AGMs helps an NGO driving support for shareholder proposals in future periods, but displays decreasing returns.
7 Discussion and Conclusion
By analyzing the timing of NGO campaigns, we demonstrate that this seemingly minor decision has substantial consequences for both targeted firms and the effectiveness of future initiatives. Furthermore, we demonstrate that NGOs’ tactics evolve over time as they seek to either indirectly pressure firms (“exposure”) or directly shape their operations (“influence”). Thus, systematically analyzing NGOs’ behavior can shed new light on how activism helps firms internalize externalities.
We first show that NGOs tend to launch their campaigns on the AGM date, as media coverage is higher during AGMs (Result 1). By doing so, NGOs elicit reactions from key stakeholders, such as shareholders and consumers (Result 2). The immediate goal of these campaigns is to influence a firm’s strategy by damaging its reputation and voicing stakeholder dissatisfaction.
However, campaigning before the AGM offers direct advantages. Specifically, when a shareholder proposal has already been filed at a firm, targeting that firm on the same issue increases the likelihood that the firm accepts the proposal outright—potentially leading to its withdrawal from the AGM ballot—or, at the very least, raises shareholder support for the proposal by an average of 8% (Result 3). Since shareholder proposals generally receive low support, early NGO campaigns can discipline managers (Levit and Malenko, 2011). Importantly, all these analyses control for NGO-by-year fixed effects and cause-by-year fixed effects, accounting for any time-varying confounding factors at the NGO level—such as experience or reputation—or at the cause level—such as specific news events—thus ensuring that globally renown NGOs (e.g., PETA) do not drive these results. Also, we provide evidence suggesting that NGO campaigns on AGMs do not target different firms compared to campaigns at other times, which helps mitigate selection issues (Appendix Table A2).
Therefore, NGOs face a trade-off, which they navigate by adapting their strategies over time. NGOs with limited activism histories tend to campaign on the AGM date, whereas more established NGOs prefer targeting firms just before the AGM to directly influence shareholder votes (Result 4). One interpretation is that as NGOs gain more experience with campaigning, they accumulate media exposure, reducing the pressure to campaign specifically on the AGM date.
Therefore, repeatedly targeting AGMs may yield diminishing returns (Result 5). First, an NGO may have already signaled its commitment to a cause, making subsequent signals weaker and reducing the effectiveness of AGM-date campaigning as a signaling device. This interpretation aligns with the view that AGM-date campaigns enhance an NGO’s “reputation.” For instance, following ”reputation-for-toughness” models, the NGO may accept an opportunity cost—delaying support for a shareholder proposal—to launch a campaign at a time when it imposes the greatest costs on a firm, thereby building a reputation that can be leveraged with other targets in the future. [^14] In these models, incumbent firms may incur short-term losses by setting predatory prices upon entry of a rival to deter future entrants (Kreps and Wilson, 1982; Milgrom and Roberts, 1982).
Second, stakeholders may come to expect more constructive approaches and proposals from established NGOs, leading NGOs to support earlier shareholder proposals or other unobserved tactics. Indeed, we find that mature NGOs—those with multiple prior campaigns in the sample—no longer inflict additional reputation losses by campaigning on the AGM date, whereas they did so in their first campaigns (compare the blue dots and red triangles in Appendix Figure B4). In fact, experienced NGOs generate the same or even more negative media coverage for a firm by launching campaigns one or two months before the AGM.
Other interpretations are unlikely to be relevant unless they covary with the number of AGM-date campaigns within a given year, NGO, firm, and cause. For instance, organizational aging (e.g., older NGOs becoming larger and more complex, with multiple ways to influence firms) is accounted for through fixed effects. These interpretations open avenues for future research, contributing to the broader discussion on civil society lobbying and activism (Bertrand et al., 2021, 2023).
In conclusion, NGOs influence corporate behavior, yet their incentives and strategies remain understudied. This paper analyzes NGO campaigns, leveraging the exogeneity of the corporate calendar to uncover their trade-offs. Similar approaches have identified stakeholder preferences, showing that AGMs are more than voting venues—they also serve as forums for firm-stakeholder negotiations, often unseen by researchers but inferred from underlying incentives (Fioretti et al., 2023).
We also acknowledge that NGOs not only campaign but also collaborate with corporations (Dahan et al., 2010). Future research could capitalize on similar exogenous variation to quantitatively study the incentives for collaboration.
Appendix A Omitted Tables
Table A1: Summary Statistics
Panel a: Variables varying by NGO
Mean 25% 50% 75% 95%
Number of campaigns by NGO 3.20 1 1 3 13
      – E campaigns 1.51 0 1 1 5
      – S campaigns 1.70 0 1 1 6
      – On AGM date 0.24 0 0 0 1
Campaign sentiment -1.72 -2 -2 -1 -1
Panel b: Variables varying by target firm
Mean 25% 50% 75% SD
Market Cap (USD bn) 175.66 21.71 75.05 209.47 263.66
Total Assets (USD bn) 98.81 21.80 59.27 138.68 124.36
Yearly Sales (USD bn) 37.26 6.85 17.35 42.41 150.58
Yearly Number of News 1,183 328 616 1,046 1,695.25
E score 5.66 4 5.58 7.3 2.22
S score 4.44 3.30 4.70 5.60 1.58
E campaign 0.47 0 0 1 0.50
S campaign 0.53 0 1 1 0.50
Panel c: Variables varying by target firm and NGO experience
High Experience Low Experience Test difference
Mean S.E. Mean S.E. p-value
Market Cap (decile) 9.23 0.10 9.07 0.04 0.13
Total Assets (decile) 9.26 0.08 9.23 0.04 0.72
Sales (decile) 9.31 0.09 9.14 0.04 0.07
Yearly Number of News (lag) 1,132.37 114.74 1,368.67 47.67 0.06
E score 5.64 0.12 5.72 0.06 0.58
S score 4.44 0.09 4.45 0.04 0.89
E campaign 0.46 0.03 0.5 0.01 0.15
S campaign 0.54 0.03 0.49 0.01 0.12
Panel d: OLS Regression of $y_{ny}=\beta_{0}+\beta_{1}\cdot\textit{NGO Experience}_{g(n)y}+\varepsilon_{ny}$
$\hat{\beta}_{1}$ (SE) w/o FEs $\hat{\beta}_{1}$ (SE) w/ FEs N
Dependent Variable ($y_{ny}$): (1) (2) (3) (4)
Market Cap Decile -0.002 (0.003) 0.003 (0.016) 1,338
Total Assets Decile 0.003 (0.004) -0.002 (0.012) 1,338
Sales Decile 0.001 (0.004) -0.002 (0.03) 1,337
Yearly Number of News 0.996 (12.4) -4.12 (16.5) 1,323
E score 0.004 (0.009) 0.041 (0.041) 1,338
S score -0.0006 (0.006) 0.019 (0.032) 1,338
E campaign 0.006∗∗∗ (0.002) 0.021∗∗ (0.010) 1,338
S campaign -0.006∗∗∗ (0.002) -0.018 (0.012) 1,338
  • Note: This table reports summary statistics on NGO campaigns (Panel a, NGO campaign data from Sigwatch and AGM dates from ISS), firms targeted by the campaigns (Panel b, accounting data from Compustat, ESG scores from MSCI, and news count based on RavenPack data), as well as an analysis of the relationship between NGO experience and targeted firms’ characteristics (Panels c and d). In Panel c, the Market Cap, Total Assets, and Sales deciles are based on the universe of US firms in Compustat on the quarter of the campaign. We use deciles instead of dollar values because highly ranked firms saw a large expansion in market cap and assets over the last years. Ranking measures like deciles avoid such a mechanical correlation with NGO Experience, which, being defined as the number of campaigns launched by NGO g in previous years, also grows over time. In Panel d, the fixed-effects are: NGO×$\times$×Year, Firm×$\times$×Year and Topic×$\times$×Year, except for the last two rows where Topic×$\times$×Year fixed effects are not included because of perfect collinearity with the dependent variables. Standard errors are clustered by firm.
Table A2: Campaign timing and characteristics of target firms
Characteristics of Target Firm
Market Cap Assets Sales Yearly Environmental Social Campaign
Dependent Variables: (decile) (decile) (decile) News Score Score Sentiment
Model: (1) (2) (3) (4) (5) (6) (7)
Variables
Campaign on AGM day (0/1) 0.469 0.495 0.374 83.2 0.469 -0.522 0.022
(0.628) (0.695) (0.550) (185.7) (0.919) (0.715) (0.383)
Fixed-effects
Industry-Year
NGO-Year
Topic-Year
Fit statistics
Observations 1,338 1,338 1,337 1,299 1,338 1,338 1,338
R2 0.93364 0.92669 0.90388 0.99232 0.95420 0.92167 0.84525
*–p$<0.1$; **–p$<0.05$; ***–p$<0.01$.
  • Note: This table reports the estimated coefficients from regressions of different target firm characteristics on a dummy indicating whether the campaign took place on the target firm’s AGM day. All regressions also include Industry-by-Year, NGO-by-Year, and Topic-by-Year fixed effects. The Market Cap, Assets, and Sales deciles are based on the entire universe of US firms in Compustat in the quarter of the campaign. Yearly News refers to the four-quarter moving average of news mentioning the target firm. Accounting data from Compustat. NGO campaign (including Sentiment) data from Sigwatch. News data from RavenPack. Environmental and Social scores are from MSCI. Standard errors are clustered by firm.
Appendix B Omitted Figures
Figure B1: Count of E&S shareholder proposals and their success
Note: This figure plots the number of E&S proposals at Russell 3,000 firms from 2006 to 2021 (solid line) and the number of successful E&S proposals (blue area) in the same period. A successful proposal is defined as a proposal that is either withdrawn before the AGM, or that goes to vote and passes. Data from ISS. Based on 5,937 shareholder proposals filed at 979 firms.
Figure B2: Persistence of the AGM month over time
Note: This figure plots the percentage of AGMs happening in each month of the year (y-axis), given the month of the previous year’s AGM (x-axis). A larger and darker circle indicates a greater proportion of meetings. Plot based on 38,471 meetings of 5,745 US companies, from 2010 to 2020.
Figure B3: Testing NGO campaign timing around the AGM date
Note: This figure plots the density test for the occurrence of NGO campaigns around firms’ AGM, following Cattaneo et al. (2020). The test is based on 521 NGO campaigns launched within 50 days of the firm’s closest AGM. The NGO data come from Sigwatch; the AGM dates are from ISS.
Figure B4: Media exposure of NGO campaigns around an AGM
(a) Main specification
(b) Controlling also for current E&S news
Note: The figure presents estimates from (1). The dependent variable is an indicator that a firm receives at least one negative E&S news in a week. Panel (b) also controls for the current count of E&S news to account for firm-week specific events that increase news coverage (not necessarily negative coverage). Standard errors are clustered at the firm level.
Figure B5: Examples of positive and negative E&S news
(a) Example of a positive E or S news
(b) Example of a negative E or S news
Notes: This figure displays examples of a positive (top) and negative (bottom) E&S news. News data from RavenPack.
Figure B6: Media sentiment of NGO campaigns around an AGM
Note:The figure presents estimates from (1). The dependent variable is the average sentiment of E&S news received by firm in a week. Standard errors are clustered at the firm level.
Figure B7: NGO campaigns and shareholder proposals
Note: This figure plots the estimated coefficients of the regression of a dummy indicating the occurrence of a shareholder proposal on month $t$ at firm $f$, relative to an NGO campaign targeting firm $f$ on the same cause, $k$ months ahead. The regression also includes month-by-year-by-industry-by-topic fixed effects. The regression is based on 1,227 NGO campaigns. Standard errors are clustered at the firm level.
Figure B8: Examples of positive and negative E&S news
(a) NGO Campaigns on AGM
(b) NGO Campaigns on Other Days
(c) Estimated Effect of Campaigning on AGM
Notes: Panels (a) and (b) plot the estimated coefficients of the regression of sales over (lagged) total assets, on a set of time dummies interacted with a dummy indicating the occurrence of an NGO campaign targeting firm $n$ from (3). Panel(c) plots the estimated $\hat{\beta}$ from
$\frac{\text{Sales}_{n,q}}{\text{Assets}_{n,q-1}}$ $=\sum^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}\cdot\text{NGO Campaign on AGM}_{n}\cdot\mathbf{1}{\{\text{distance}=\tau\}}\right)$
$+\sum^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\tilde{\beta}_{\tau}\cdot\mathbf{1}{\{\text{distance}=\tau\}}\right)+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$
which interacts the time dummy, the occurrence of a campaign, and a dummy indicating that the campaign happened on AGM day. The regressions also include firm and quarter-by-industry fixed effects. The reference period is -1 quarter, consistent with the other analyses. The number of observations is 10,188, 351,351, and 361,539, respectively. Standard errors are clustered at the firm level.
Figure B9: Campaign Timing and E&S Score Improvement
Note: This figure plots the estimated coefficients of the following regression:
$\Delta_{4}\text{E\&S Score}_{n,q}=\sum_{l=-6}^{5}\text{Campaign}_{n,q,l}+\textbf{X}_{n,q}\gamma^{\prime}+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$ (B1)
where $\Delta_{4}\text{E\&S Score}_{i,q}$ is the change in firm $n$’s average of its Environmental and Social scores between quarter $q$ and quarter $q+4$, and $\sum_{l=-6}^{6}\text{Campaign}_{n,q,l}$ is a set of dummy variables indicating the occurrence of an NGO campaign against firm $n$ in quarter $q$, being $l$ months away from the closest AGM. The regression also controls for firm size (logged total assets), current average of E and S scores, firm fixed effects ($\alpha_{n}$), and industry-by-quarter fixed effects ($\alpha_{s(n),q}$). ESG scores are from MSCI, NGO campaigns data from Sigwatch, firm characteristics from Compustat, and AGM dates from ISS. The regression has 47,082 observations. Standard errors are clustered at the firm level.
Unlike an event study, equation (B1) is a distributed leads-and-lags regression. We omit a reference period, so the coefficients are interpreted relative to firms that did not experience a campaign. This choice allows us to directly compare the effects of early campaigns with those launched on the AGM date relative to firms without campaigns. Results remain similar if we instead set month –4 as the reference period, as in the other analyses.
Figure B10: NGO experience and the timing of NGO campaigns – Alternative set of Fixed Effects
Note: This figure plots the estimated coefficients from the regressions of the distance between NGO campaigns and the targeted firm’s closest AGM (in months) on dummies indicating the experience of the NGO, measured by the number of campaigns launched before. The regression includes replaces NGO-by-year fixed effects with separate fixed effects for each NGO and each year. Hence the coefficients $\{\hat{\beta}_{\tau}\}_{\tau}$ in (5) is also identified from variation in NGO experience over years. Standard errors are clustered at the NGO level.
Figure B11: Campaign timing for NGOs whose first campaign is after the AGM
Note: This figure plots the estimated coefficients from the regressions (5) of the distance between NGO campaigns and the targeted firm’s closest AGM (in months) on dummies indicating the experience of the NGO, measured by the number of campaigns launched before. The panel focuses on NGOs whose first campaign was within the two weeks following the AGM. Standard errors are clustered at the NGO level.
Figure B12: NGO experience and target firm’s size
Note: This figure plots the estimated coefficients from regressing the target firm’s market cap (in deciles to control for the fact that the tails of the market cap distribution have thickened in recent years) on dummies indicating the experience of the NGO, measured by the number of campaigns it had launched in previous year as in (5). Standard errors are clustered at the firm level. The number of observations is 1,338.
Figure B13: NGO experience and target firm’s E&S characteristics
(a) Target firm’s E score for Environmental campaigns
(b) Target firm’s S score for Social campaigns
Note: This figure plots the estimated coefficients from regressing either a target firm’s environmental score (top panel) or its social score (bottom panel) on dummies indicating the experience of the NGO, measured by either the number of environmental (top panel) or social (bottom panel) campaigns it had launched in previous year as in (5). Standard errors are clustered at the firm level. The number of observations is 636 and 698, respectively.
Appendix C Matching Proposals And NGO Campaigns
This section outlines the process of matching each shareholder proposal in the ISS dataset with a relevant NGO campaign cause.
Sigwatch assigns a“Generic Topic” to each campaign in their dataset. To streamline the analysis, we combine certain causes when they are difficult to distinguish or when one category has few campaigns and can reasonably fit within a broader category. Specifically, we merge “Labor Rights” with “Human & Labor Rights,” “Children” with ”Human & Labor Rights,” “CSR / ESG” with “CSR,” and “Biodiversity” and “Wildlife” with ”Environment.” For campaigns under the “Safety” cause, which includes only 12 entries, we examine each campaign’s specific purpose and classify it as either “Product Safety” or “Food Safety.”
Using these consolidated “Generic Cause Names”, we then categorize each shareholder proposal based on its short description. When a shareholder proposal is submitted, it includes a brief, one-sentence description, which is provided in the ISS dataset and serves as the foundation for cause classification. We start by assigning proposals to causes based on keywords in their descriptions and then manually classify any remaining proposals.
Table C1: Proposal Causes
NGO Campaign Category Example of Shareholder Proposals
Topic Classification (E or S)
Climate Change E Encourage and Support Carbon Tax to Address Climate Change
Environment E Stop financing mountain top removal coal mining
Pollution E Report on Efforts to Reduce Pollution From Products and Operations
Rainforest E Report on impact on rainforest
Sustainability E Establish board committee on sustainability
Waste Generation E Do not increase radioactive waste production
Water E Report on supply chain water impacts
Animal Rights & Welfare S Review suppliers’ animal slaughter methods
CSR S Disclose Charitable Contributions
Food Safety S Label Genetically Modified Ingredients
Health S Reduce Tobacco Harm to Health
Human & Labor Rights S End Sri Lanka Trade Partnerships Until Human Rights Improve
Indigenous People S Adopt Global Policy Regarding the Rights of Indigenous People
Intellectual Property Rights S Implement a Process to Fully Vet Whether a
Supervisor is a Bona Fide Inventor Before a Patent is Filed
Politics S Affirm Political Nonpartisanship
Product Safety S Report on Nanomaterial Product Safety
Worker Safety S Report on Reducing Risks of Accidents
  • Note: This Table reports the 17 topics assigned to each shareholder proposal in the sample. Shareholder proposals data from ISS. The topics are from Sigwatch’s categorization of campaigns.