### v6

# NGO Activism:

Abstract
We examine how NGOs time campaigns to influence corporate decisions at Annual General Meetings (AGMs). Using data on 2,500 U.S. campaigns, we find that NGOs frequently launch campaigns on the targeted firm’s AGM date, which boosts media exposure and triggers future shareholder proposals but does not sway current AGM votes or affect sales. As NGOs gain experience and reputational capital, they shift to earlier pre-AGM campaigns that increase shareholder support at the AGM and lead to improvements in corporate behavior. These patterns reveal a strategic trade-off between maximizing exposure and achieving direct influence.
JEL classifications: L21, L31
Keywords: Non-Governmental Organizations, NGO campaigns, social and environmental responsibility, ESG, firm dynamics, non-profit objectives
1 Introduction
Non-governmental organizations (NGOs) are increasingly prominent actors in shaping firms’ environmental and social policies (Baron, 2001; Egorov and Harstad, 2017; Yaziji and Doh, 2009). Yet, their strategic choices remain poorly understood. Much of the literature treats NGOs as either principled advocates motivated by social good (Besley and Ghatak, 2005; Bénabou and Tirole, 2006; Daubanes and Rochet, 2019) or as dysfunctional entities prone to corporate capture and strategic manipulation (Grossman and Helpman, 1994; Bertrand et al., 2020; Akey et al., 2021). We argue and demonstrate empirically that NGOs adapt their strategies in response to incentives, shaping which stakeholders they engage and ultimately affecting the social and environmental performance of targeted firms.
To study how NGOs respond to incentives and their implications for firms, we focus on a setting where firms are especially exposed to outside pressure: the Annual General Meeting (AGM). AGMs are predictable focal points at which executives present results, shareholders vote on proposals, and media attention peaks (Fioretti et al., 2025). Figure 1 shows that NGO campaigns cluster heavily around AGM weeks: incidence is four to six times higher than in non-AGM weeks (Panel a), and media coverage spikes simultaneously (Panel b). These striking patterns raise a central question: do NGOs target AGMs mainly to maximize exposure, or do they also use campaigns to exert influence over corporate outcomes?
Figure 1: Timing of NGO campaigns and media attention around AGMs
(a) Campaigns peak during AGMs
(b) Media coverage rises during AGMs
Note: Panel (a) displays the number of NGO campaigns launched relative to the date of the Annual General Meeting of the targeted firm (week 0), based on 1,544 NGO campaigns. Panel (b) compares the weekly average count of news during the AGM week (gray bars) and outside of the AGM week (white bars), based on weekly observations for 5,211 US firms. The NGO data come from Sigwatch; the news data are from Ravenpack.
Examining campaign timing reveals that less experienced NGOs concentrate campaigns on the AGM date, where visibility is maximized but most shareholder votes are already casted (Gent et al., 2015; Bebchuk et al., 2020). These campaigns cannot change immediate voting outcomes but are found to mobilize other stakeholders, such as future shareholder proponents. For instance, as the NGO Public Eye put it in its 2025 campaign timed on the AGM date of Inditex: “As shareholders, you have the power to put Inditex on a sustainable course” (Clean Clothes Campaign, 2025). Indeed, we find little evidence that other stakeholders are moved by these campaigns. For example, we find no evidence that sales are differently affected based on campaign timing relative to the AGM date, indicating that consumers may not be the main audience of NGOs.
While they do not sway votes or sales, AGM-timed campaigns significantly increase the visibility of NGOs themselves. Using Google search data, we find that such campaigns generate sharp increases in search intensity for the campaigning NGO, particularly when it is relatively inexperienced. These visibility gains persist for several months and are substantially larger than those observed for campaigns launched at other times. This suggests that AGM-day actions, while limited in shifting immediate corporate outcomes, are instrumental in boosting public recognition of NGOs.
As NGOs accumulate visibility through AGM-date campaigns, their strategic calculus changes. Because most shareholder votes are cast well before the meeting, campaigns on the day itself rarely sway immediate outcomes. By contrast, campaigns launched in the months leading up to the AGM substantially raise the likelihood that related shareholder proposals succeed. Intervening one to three months before the meeting boosts average shareholder support by 8%, while even earlier campaigns often prompt firms to settle before proposals reach a vote, leading to measurable improvements in environmental and social performance.
Consistent with this shift from exposure to influence, we find that NGOs progressively move their campaigns forward in time as they gain experience and reputational capital. First-time campaigns are concentrated on the AGM date, where visibility is highest, while subsequent efforts are launched progressively earlier, when they can still shape shareholder decisions. This pattern is especially pronounced among NGOs with stronger reputations, proxied by higher search visibility, which campaign several months ahead of the AGM rather than on the day itself.
Repeated targeting of AGMs also plays a distinct role. By returning to these focal moments, NGOs build credibility and reputational standing, which in turn increases the effectiveness of later campaigns. Yet such returns are not indefinite: once credibility is established, additional AGM-day campaigns yield diminishing benefits. Taken together, this lifecycle of tactics—initially focused on visibility, later on direct influence—shows how NGOs adaptively balance exposure, reputation, and pressure on firms, with implications for firms’ social and environmental impact.
In this sense, AGM-day campaigns act as costly signals of commitment that help NGOs build credibility (Kreps and Wilson, 1982; Milgrom and Roberts, 1982). As NGOs accumulate reputational capital (Bar-Isaac et al., 2008; Baur and Palazzo, 2011; Mitchell and Stroup, 2017), their strategy shifts: campaigns are launched months before the AGM, when they can still sway shareholder support and discipline managers.
This dynamic lifecycle highlights NGOs’ role as coordinators that align stakeholder actions and channel them into constructive corporate change, as their strategies shift over time to balance visibility and influence (Briscoe and Gupta, 2016), with campaign tactics evolving dynamically across the life cycle to shape both corporate responses and stakeholder engagement (Cabral and Hortacsu, 2010).
By documenting this strategic adaptation, we provide new conditions under which firms internalize their externalities (Fioretti, 2022). More broadly, our work offers empirical support for theories asserting that stakeholders contribute resources only when they expect to capture part of the value created, underscoring the interplay between stakeholder involvement, reputation, and corporate change (Porter and Kramer, 2011; Kaplan, 2019; McGahan and Pongeluppe, 2023).
Prior studies highlight the role of shareholders influencing firms via voice through voting or exit via divestments (e.g., Cuñat et al., 2016; Broccardo et al., 2022; Bonnefon et al., 2022; Hart et al., 2022; Saint-Jean, 2024), and of consumers through willingness to pay (Gneezy et al., 2010). We instead show that NGOs act as coordination mechanisms, enabling diverse stakeholders to collectively discipline firms (Koenig and Poncet, 2019; Chuah et al., 2024; Mazet-Sonilhac and Mésonnier, 2023; McGahan, 2023). In doing so, NGOs expand the set of channels through which civil society shapes corporate conduct.
Finally, our methodology relates to the literature on the strategic timing of actions to exploit fluctuations in media attention. Prior work shows that politicians and governments time executive orders, military operations, and even disaster relief to coincide with or avoid competing news (Eisensee and Strömberg, 2007; Durante and Zhuravskaya, 2018; Djourelova and Durante, 2022), and that social movements strategically leverage focal dates to amplify visibility (Madestam et al., 2013). We show that NGOs adopt similar tactics in the corporate domain.
The paper is organized as follows. Section 2 outlines the data, and Section 3 provides background on AGM dates. Sections 4 and 5 examine NGO campaigns as tools to gain media exposure or to influence firms through key stakeholders, respectively. Section 6 brings these channels together by analyzing how NGO reputation enhances the success of shareholder proposals. Section 7 discusses the results and concludes.
2 Data
Our dataset integrates four sources: NGO campaigns, online attention, shareholder proposals, and firm-level news.
First, we use data from Sigwatch, a consulting firm tracking global NGO campaigns against corporations (Hatte and Koenig, 2020). The data record the campaign’s public release date, cause (e.g., “GMOs in food”), the targeted company and its prominence, and sentiment (ranging from –2 to +2, typically negative). We focus on single-firm, negative-sentiment campaigns where the firm appears in the headline. [^1] Since our analysis links campaign timing with a firm’s AGM date, we exclude multi-firm campaigns as their AGM dates may differ. Single-firm campaigns represent 93% of the original sample. The final dataset spans 2,512 campaigns launched by 731 NGOs against 523 firms.
Second, we collect monthly Google Trends scores for all NGOs in our sample from 2010 to 2022. Scores are normalized on a 0–100 scale, either (i) relative to the NGO’s own peak searches (within-NGO) or (ii) relative to a benchmark NGO in the same month (between-NGOs). We use Greenpeace—the most searched NGO—as the benchmark for comparability across organizations.
Third, we use data from Institutional Shareholder Services (ISS) on proposals at the Annual General Meetings (AGMs) of the 3,000 largest US-listed firms. We select 4,483 environmental and social (E&S) proposals filed between 2010 and 2021 (837 firms in total). About one-third of these proposals are successful, either withdrawn before the AGM (after firm commitments) or passed in a vote. Both the number of E&S proposals and their success rate increased steadily over the period.
Lastly, we obtain firm-level news from RavenPack, which transforms unstructured news into structured data with sentiment and event tagging. We count news items related to US-listed firms from 2010 to 2020, retaining only items with a Relevance Score of at least 75%. The final sample includes 5,212 firms.
We link Sigwatch, ISS, and RavenPack data using standard identifiers (CUSIP, ISIN, year). Appendix Table A1 reports descriptive statistics. Panel (a) shows that, on average, NGOs launch 3.2 campaigns during the sample, 24% of which are on the target AGM date, though over half of NGOs appear only once. Most campaigns focus on social rather than environmental issues, but the distribution is highly skewed: a small set of NGOs accounts for a disproportionate share of activity, while the median NGO runs only a single campaign. Panel (b) documents substantial variation across targeted firms in size and news coverage. Targets range from very large firms with extensive media presence to smaller firms with limited coverage, suggesting that NGOs are not exclusively focused on high-visibility corporations. 53% of the campaigns are on social causes, with the remaining being environmental. We return to the remaining panels of Table A1 in Section 4, after first providing background on AGM dates.
3 Institutional Background on AGM Dates
Our empirical approach examines the timing of NGO campaigns relative to AGMs, leveraging the exogeneity of AGM dates, which remain stable within firms across years due to SEC regulations. Rule 14a-8 mandates shareholder proposals be filed 120 days before the proxy statement, anchoring AGM dates across years (Glac, 2014). Appendix Figure B1 confirms this stability, with exceptions largely due to pandemic-related adjustments allowed by the SEC (2020) to facilitate shareholder attendance at AGMs. [^2] Results remain robust excluding the covid period (2020–2021).
While external events may shape NGOs’ incentives to campaign, they do not alter the timing of AGMs. This stability makes AGMs a useful benchmark: NGOs can choose to react immediately to a negative event concerning a firm, or alternatively wait and align their campaign with its AGM, when the firm is most exposed to scrutiny, as shown in Panel (b) of Figure 1, which shows that news coverage spikes in AGM weeks. Consistent with this view, we show in the Appendix that AGM-timed campaigns generate more negative news coverage than those launched at other times (Appendix Figure B2). Taken together, these patterns highlight why AGMs provide a natural focal point to study the strategic trade-off between maximizing visibility and exerting direct influence.
4 Timing Campaigns for Visibility
Having shown that NGO campaigns cluster around AGM dates and that they increase a target’s negative media coverage, we next ask whether such timing enhances NGOs’ visibility. We use Google search activity as a proxy for NGO visibility and estimate event-study regressions comparing campaigns launched on the AGM date with those launched at other times:
$\text{Google Trends Score}_{g,m,y}$ $=\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}^{AGM}\cdot\text{AGM Campaign}_{c}\cdot\mathbf{1}_{\{M_{c,m,y}=\tau\}}\right)$ (1)
$\quad+\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}^{Non-AGM}\cdot\mathbf{1}_{\{M_{c,m,y}=\tau\}}\right)$
$\quad+\alpha_{c,y}+\alpha_{m,y}+u_{g,m,y}.$
Because NGOs typically campaign multiple times, we follow Baker et al. (2022a) and stack all campaign events. Each event is denoted $c(i,n,g,y,m)$, where campaign $c(\cdot)$ is identified by cause $i$ (e.g., pollution), NGO $g$ (e.g., Greenpeace), target firm $n$ (e.g., Shell), calendar year $y$, and month $m$. We then analyze variation in the Google Trends score for NGO $g$ within a five‐month window around the start of $g$’s campaign; this window may spill into years y1y-1 or $y+1$ if the campaign occurs early or late in year $y$.
The main explanatory variables are event-time dummies, $\mathbf{1}_{\{M_{c,m,y}=\tau\}}$, equal to one if month $m$ is exactly $\tau$ months from campaign $c$, with $\tau\in[-5,5]$. The omitted category is the month immediately preceding the campaign ($\tau=-1$). These indicators are interacted with $\text{AGM Campaign}_{c}$, a dummy equal to one if campaign $c$ was launched on the target firm’s AGM date. Thus $\beta^{AGM}_{\tau}$ estimates the increase in visibility for NGO campaigning on an AGM date, over the direct effect of campaigning at any date, $\beta^{Non-AGM}$.
The remaining terms include fixed effects at the NGO–firm–year level ($\alpha_{c,y}$), which absorb all variation across NGO-firm pairs over years, as well as month-by-year fixed effects ($\alpha_{m,y}$). Our main coefficients of interest ($\beta^{AGM}_{\tau}$) are therefore identified from within-pair variation across months. Standard errors are clustered at the NGO level.
Under the hypothesis that AGM campaigns affect the visibility of experienced and inexperienced NGOs differently, we estimate (1) separately for an NGO’s first few campaigns and for later ones. Panel (a) of Figure 2 reports the estimated coefficients of $\beta^{AGM}_{\tau}$, splitting the sample into an NGO’s first three campaigns (blue dots) and its subsequent campaigns (red triangles). The results indicate that AGM campaigns generate a sharp and statistically significant increase in NGO visibility, raising Google Trends scores by almost 50% in the following three months. For later campaigns, the effect is weaker—about 40% in the first month and around 10% in the next two months—and not statistically significant in any period, consistent with diminishing visibility returns as NGOs become more established. These patterns suggest that inexperienced NGOs rely more heavily on AGM timing to build name recognition, whereas established NGOs gain less from repeating the same tactic.
Figure 2: Increased NGO visibility around an AGM campaign
(a) The first AGM campaigns bring visibility
(b) No evidence of Selection
Note: Panel (a) plots the estimated $\beta^{AGM}$ coefficients from (1). The dependent variable is the monthly Google Trends score for the NGO. Blue dots are estimated by restricting the sample to an NGO’s first three campaigns, while red triangles consider later campaigns. Standard errors are clustered at the campaign level. Panel (b) reports the estimated coefficients of the regression of an indicator that an NGO campaign was launched on the target firm’s AGM date on the displayed (standardized) variables, and year fixed effects. Error bars shows 90% (shaded) and 95% confidence intervals using standard errors that are clustered by firm.
We show that these results are robust to addressing three main concerns. First, the stacked event-study framework implicitly treats campaign timing as simultaneous across NGOs, whereas in reality campaigns are staggered across organizations and calendar months. This misalignment can distort estimates when treatment effects are heterogeneous across groups or over time. Second, some NGOs launch multiple campaigns within the same year, which may confound the analysis if two campaigns fall within five months of each other. To address both issues, Appendix C replicates the analysis using the staggered difference-in-differences estimator of Sun and Abraham (2021), which explicitly accounts for variation in treatment timing, as well as a distributed-lag specification that flexibly controls for past campaigns within a year. Both approaches yield similar results, with much larger gains for “unexperienced” NGOs.
Third, although we find no evidence of pre-trends, NGOs might still select targets strategically to raise visibility—for example, by focusing on firms that are more vulnerable to pressure. Panel (b) of Figure 2 addresses this concern by regressing an indicator for AGM-timed campaigns on firms’ financial and ESG characteristics and the cause of the campaign, with year fixed effects and firm-clustered standard errors. [^3] The independent variables include the target firm’s market capitalization (Market Cap), yearly sales (Sales), total assets (Total Assets), number of news over the past 12 months (News count), the firm’s Environmental and Social scores (Env. score, Soc. score), how negative the campaign is towards the firm (Campaign sentiment), and an indicator for the campaign cause being related to an environmental topic (Env. campaign). Note that no coefficient is significantly different from zero. We find that a one standard deviation increase in any of these variables moves the likelihood of an AGM campaign by less than 5% of a standard deviation, indicating little selection on observables.
Further evidence on selection based on NGO experience comes from Table A1. Panel (c) compares firms targeted by NGOs with high versus low campaign experience and finds no significant differences across the same variables studied in Panel (b) of Figure 2. Panel (d) corroborates this by regressing firm characteristics on NGO experience (i.e., number of past campaigns), showing no correlation once fixed effects are included. The only systematic trend is that NGOs increasingly focus on environmental rather than social issues in later years, reflecting their rising importance.
Thus, the evidence presented so far points to:
Result 1.
Gains in visibility are greater for unexperienced NGOs campaigning on the target’s AGM day than for experienced NGOs.
For clarity, each of the next sections highlights its main findings in a similar manner.
4.1 Stakeholders’ Reactions
We next ask how campaign timing matters for stakeholders and firms’ social and environmental outcomes. We focus on shareholders and consumers, with supporting exhibits in the Online Appendix.
4.1.1 Shareholders
Shareholders can be seen as primary proponent of firm’s social values (e.g., Naaraayanan et al., 2021; Baker et al., 2022b). Shareholders may respond more to causes campaigned during AGMs, when media coverage is highest. We define as “related” a shareholder proposal on the same cause of the campaign in the previous year. Appendix D describes the matching procedure.
For each campaign in our sample, we examine the probability that the next AGM includes a related shareholder proposal, depending on the campaign’s timing. To this end, we recalendarize time so that month 0 corresponds to the AGM month for each firm. As discussed in Section 3, AGM months are stable over time; thus, this recentering does not alter the distribution of months across years or risk double counting. [^4] All subsequent analyses use the same recalendarization. Months –5 and +5 are defined as 45-day intervals to ensure the recalendarized year covers the entire calendar. We estimate the following regression:
$\text{Proposal on Same Cause}_{i,n,y}$ $=\sumop\displaylimits^{\tau=-4}_{\begin{subarray}{c}\tau=-15\\ \tau\neq-4\end{subarray}}\left(\beta_{\tau}\cdot\mathbf{1}_{\{\text{Distance to AGM${}_{c(i,n,g,y,m)}$}=-\tau\}}\right)$ (2)
$+\alpha_{i,s(n),g,m,y}+\gamma_{n}+u_{i,n,y}$
For each campaign $c(i,n,g,y,m)$ launched by NGO $g$ against firm $n$ on cause $i$ in year $y$ and month $m$, we analyze the likelihood that the next AGM of firm $n$ includes a shareholder proposal on the same cause. Formally, $\text{Proposal on Same Cause}_{i,n,y}$ equals one if such a proposal is observed within 4 to 15 months after campaign $c(\cdot)$, and zero otherwise, with timing measured by the indicator $\mathbf{1}_{\{\text{Distance to AGM}_{c(i,n,g,y,m)}=-\tau\}}$. Since SEC 14a8(e)(2) rule prevents shareholders from including proposals in the proxy statement within 120 days of its release, we cap the analysis at four months before the previous year’s AGM and set the reference month to $\tau=-4$ because of SEC rules.
Although firm-by-time fixed effects are not possible in this setting as they would absorb the variation used to estimate $\beta_{\tau}$, this analysis purges all the variation from cause-by-industry-by-NGO-by-month-by-year through $\alpha_{i,s(n),g,m,y}$ and firm invariant characteristics through $\gamma_{n}$, which we view as the main confounding factor since issue salience varies systematically across firms, industries, causes, and NGOs over time. We cluster standard errors by firm.
Figure B3 shows that only campaigns launched twelve months before the next AGM increase the likelihood of a shareholder proposal on the same issue. The effect is substantial—an increase of 12 percentage points, or 44% relative to the baseline. This timing suggests that campaigns around an AGM may spur proposals at the next one, possibly due to heightened media attention.
Importantly, we can rule out the possibility that NGOs partnering with shareholders drive this outcome. If that were the case, submitting a shareholder proposal followed by an early campaign would allow the partners to immediately gain broad support on the proposal, without having to wait a year to advocate for change at the next AGM.
4.1.2 Consumers
Drawing from quarterly data from Compustat, we investigate the responses of consumers to the timing of NGO campaigns by studying the relation between the target’s “Sales to Asset ratio” and the distance between the campaign and the target’s AGM through
$\frac{\text{Sales}_{n,q}}{\text{Assets}_{n,q-1}}$ $=\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}\cdot\text{NGO Campaign}_{n}\cdot\mathbf{1}_{\{\text{distance}=\tau\}}\right)+\alpha_{n,y}+\alpha_{s(n),q}+u_{n,q},$ (3)
where $q$ indicates a quarter and $\alpha_{n,y}$ and $\alpha_{s(n),q}$ are fixed effects at either the firm-by-year level or the industry-by-quarter-by-year level, respectively. The dependent value normalizes $n$’s sales by its asset to account for time variation in relative firm sizes, which is large in our ten-year sample. Standard errors are clustered at the firm level.
Appendix Figure B4 shows the estimated $\hat{\beta}_{\tau}$} coefficients separately for NGO campaigns (i) planned on an AGM date (Panel a) and (ii) not planned on an AGM date (Panel b). We find no consumer response in either case, indicating that shareholders may be the primary audience for NGO campaigns. [^5] Unreported results show also no difference between experienced and unexperienced NGOs.
4.2 Does AGM-Targeting Create Positive Change?
If public shaming effectively influenced firm behavior, the AGM month should deliver the largest improvements in E&S scores, given that it triggers heightened negative news coverage of the target firm (Appendix Figure B2). However, Appendix Figure B5 shows that only campaigns launched 3 to 4 months before the AGM lead to subsequent increases in firms’ E&S scores—rising by around 3% over the following year. By contrast, campaigns launched at other times, including during the AGM month, have no economically or statistically significant effect.
Although ESG scores are an imperfect measure of a firm’s social impact (e.g., Kotsantonis and Serafeim, 2019; Allcott et al., 2022), this finding suggests that NGOs’ preference for targeting AGM dates may be driven by factors beyond their direct influence on firms. Therefore, the results in this section suggest that:
Result 2.
NGO campaigns on AGM dates spur shareholders to place related proposals at future AGMs of the targeted firm, but they produce no contemporaneous improvements in the firm’s E&S scores.
5 Timing Campaigns to Influence AGM Outcomes
Early campaigns could benefit an NGO aiming to influence current AGM outcomes. Since most votes are cast by proxy before the AGM, targeting the AGM date may limit an NGO’s ability to sway a significant share of the votes. For instance, over 85% of vote-eligible shares for Microsoft’s 2023 AGM were cast by proxy before the meeting, indicating that campaigning around AGM dates may have little impact on shareholders’ votes (Microsoft, 2023).
We define a proposal as successful if either it is withdrawn by the shareholder before the AGM or it secures at least $50\%+1$ vote at the AGM. We first look at the influence of NGO campaigns on the success of related shareholder proposals using the following specification on the whole sample of shareholder proposals:
$\text{Success}_{i,n,m,y}=\sumop\displaylimits^{0}_{\begin{subarray}{c}\tau=-6\end{subarray}}\beta_{\tau}\cdot\text{Campaign}_{i,n,m+\tau,y}+\alpha_{n,y}+\alpha_{i,y}+\alpha_{m}+u_{i,n,m,y},$ (4)
where $\text{Success}_{i,n,m,y}$” is an indicator for the success of proposal $i$ at firm $n$ in month $m$ in year $t$, and $\text{Campaign}_{i,n,m+\tau,t}$” is an indicator for the occurrence of an NGO campaign on the same cause as proposal $i$, $\tau$ months before the AGM. The reference period covers the months that are more than six months before the AGM of year yy—i.e., $m\in[-11,-7]$, as campaigns launched during that period are more likely to be in reaction to the previous year’s AGM results. We account for firm-by-year, and cause-by-year fixed effects, as well as seasonality with month fixed effects, thereby exploiting only variation across campaigns and causes within firm-year pairs. The standard errors are clustered by firm.
Figure 3: NGOs’ campaigns before the AGM help E&S shareholder proposals succeed
(a) Proposal accepted or withdrawn
(b) Vote share in favor of proposal
Note: The figure plots the estimated coefficients from (4). The dependent variables are an indicator variable that is one if a shareholder proposal is withdrawn or successful if voted (911 proposals) in Panel (a), and the percentage of votes in favor of a proposal (2,078 proposals) in Panel (b). Standard errors are clustered at the firm level.
Panel (a) Figure 3 plots the estimated ${\beta}_{\tau}$. We find that NGO campaigns happening 6 months and 1 month before the AGM positively influence the outcome of related shareholder proposals. They increase, respectively, the probability of success by 2 percentage points and 3 percentage points more than a campaign launched more than six months earlier, corresponding to a relative increase between 6 and 8%.
To understand why campaigns launched six months before the AGM are also somewhat successful, we re-estimate (4) on the subset of proposals that went to a vote, using the percentage of votes in favor as the dependent variable. [^6] Since this analysis focuses on voted proposals, we also control for the voting recommendations of management and proxy advisors such as ISS (“For,” “Against,” or “Abstain”). Panel (a) of Figure 3 shows that only NGO campaigns launched one or three months before the AGM increase support for related shareholder proposals, with an increase of nearly 2 percentage points. Given the generally low baseline support for these proposals (24%), this corresponds to an 8% increase in shareholder support when an NGO campaign addresses the same cause, on average.
In contrast to Panel (a), Panel (b) detects no significant increase in support for campaigns initiated as early as six (or five) months before the AGM. This suggests that the difference between the two figures arises from shareholder proposals being withdrawn after reaching an agreement with the firm. As a result of these early campaigns, the target firm may take actions to resolve the issue highlighted in both the campaign and the shareholder proposal, leading to an improvement in its E&S score, which supports the finding in Section 4.2 that only early campaigns result in better E&S scores.
Note that the subset of shareholder proposals going to a vote is not random; an argument raised, for instance, by Flammer (2015). Given the generally low support of shareholder proposals going to vote, managers may let proposals that are the least likely to succeed go to a vote. However, those votes can still be helpful in disciplining firms as large shareholder support to a cause may be hard to conceal for managers (Levit and Malenko, 2011). This reinforces the importance of analyzing all E&S proposals, including those withdrawn as in Panel (a).
The evidence in this section points to:
Result 3.
Pre-AGM campaigns raise shareholder support and proposal success, while AGM-day campaigns have little effect.
6 Exposure vs. Influence
Sections 4 and 5 outlined a tradeoff: campaigns launched close to an AGM maximize exposure, while those launched earlier maximize influence. In this section, we examine how NGOs navigate this tradeoff as their visibility increases. We first study whether greater experience or reputation leads NGOs to shift campaign timing from the AGM date to earlier months. We then ask whether such reputation translates into greater effectiveness at influencing shareholder proposals.
6.1 NGO Experience and Campaign Timing
We first examine whether NGOs, as they gain experience, shift from targeting AGMs to launching earlier campaigns. Specifically, we estimate
$\text{Distance to AGM}_{c(i,n,g,y,m)}$ $=\sumop\displaylimits_{\tau=0.1}^{0.9}\beta_{\tau}\cdot\mathbf{1}_{\left\{\text{Campaign \#}_{g,y-1}=\tau\right\}}$ (5)
$\quad+\alpha_{t(\cdot),y}+\alpha_{g,y}+\alpha_{n,y}+u_{c(i,n,g,y,m)}.$
where the dependent variable measures the number of months between campaign $c(\cdot)$ and the targeted firm’s AGM, and the key explanatory variable is the sequence number of campaigns by NGO $g$ by year $y$. $\alpha_{t(\cdot),y}$, $\alpha_{g,y}$, and $\alpha_{n,y}$ are topic-by-year, NGO-by-year, and firm-by-year fixed effects, respectively. Identification thus comes from NGOs launching multiple campaigns in the same year. [^7] Appendix Figure B6 shows similar results when controlling only for firm and NGO fixed effects, so that the estimation of $\hat{\beta}_{\tau}$ come from variation across both campaigns and years.
Panel (a) of Figure 4 shows that first campaigns tend to occur right at the AGM, while subsequent ones are launched progressively earlier. This pattern is consistent with NGOs initially using AGMs to gain visibility and later exploiting their visibility to influence voting outcomes.
These patterns are not driven by alternative explanations. First, they are unlikely to reflect NGOs learning that campaigning on the AGM date does not affect voting, since most shareholders have already voted by proxy. Indeed, the same downward trend in campaign timing holds when focusing on NGOs whose first campaign occurred after the AGM (Appendix Figure B7). Second, they are not driven by selection of more “influenceable” targets: neither target size (market capitalization) nor ESG scores covary with NGO experience (Appendix Figures B8 and B9).
6.2 Reputation and Campaign Timing
We next use Google Trends scores to proxy for an NGO’s reputation. We compute the average search score for NGO $g$ in year $y$ relative to Greenpeace in year y1y-1, discretized into deciles and regress:
$\text{Distance to AGM}_{c(i,n,g,y,m)}$ $=\sumop\displaylimits_{\tau=0.1}^{0.9}\left(\beta_{\tau}\cdot\mathbf{1}_{\left\{\text{Relative Google Score}_{g,y-1}=\tau\!\right\}}\right)$ (6)
$\quad+\alpha_{t(\cdot)}+\alpha_{y}+u_{c(i,n,g,y,m)},$
which is similar to model (6) with $\text{Relative Google Score}_{g,y-1}$ as a measure of NGO visibility. $\alpha_{t(\cdot)}$ and $\alpha_{y}$ are topic and year fixed effects. Panel (b) of Figure 4 shows that only the most visible NGOs—those in the top deciles of relative searches—campaign several months ahead of AGMs, while less visible NGOs continue to concentrate on AGM dates.
Taken together, these results indicate a systematic shift: NGOs start by targeting AGMs to maximize exposure, and as their reputation or experience grows, they move campaigns forward in time to increase their chances of influencing firm decisions.
Result 4.
NGOs initially target AGM dates to gain visibility. As their experience or reputation increases, they advance campaigns by several months to influence the outcome of shareholder proposals.
Figure 4: NGO reputation and campaign timing
(a) More experienced NGOs campaign earlier
(b) More visible NGOs campaign earlier
Note: This figure plots the estimated coefficients from (5). The dependent variable is the months between an NGO campaign and the targeted firm’s AGM date. Standard errors are clustered at the NGO level.
6.3 NGO Reputation and AGM Influence
We now ask whether NGO experience and reputation, as measured in the previous sections, translates into greater influence over shareholder proposals. To test this, we estimate
$\text{Success}_{i,n,y}$ $=\beta_{1}\,\text{Campaign on Same Cause}_{c(i,n,g,y,m)}+\beta_{2}\,\text{Experience}_{g,y-1}$ (7)
$+\beta_{3}\,\text{Campaign on Same Cause}_{c(i,n,g,y,m)}\cdot\text{Experience}_{g,y-1}$
$+\alpha_{n,y}+\alpha_{g,y}+\alpha_{i,y}+\alpha_{m(c(i,n,g))}+u_{i,n,y},$
where $\text{Success}_{i,n,y}$” indicates whether a proposal is withdrawn or passes a vote, as in equation (4). The key variable $\text{Campaign on Same Cause}_{i,n,y}$” equals one if NGO $g$ campaigned against firm $n$ on cause $i$ in the ten months preceding the AGM. $\text{Experience}_{g,y-1}$” counts NGO $g$’s past campaigns. We absorb firm-by-year, NGO-by-year, cause-by-year, and month variation through fixed effects, and cluster standard errors at the firm level.
Table 1: NGO experience at targeting AGMs and success at supporting current shareholder proposals
Dependent Variable: Proposal Passes (0/1)
Model: (1) (2) (3) (4)
Campaign on Same Topic -0.0003 -0.013 -0.005 -0.001
(0.026) (0.028) (0.025) (0.030)
Experience -0.034
Campaign on Same Topic $\times$ Experience 0.078∗∗∗
Experience square 0.008∗∗
Campaign on Same Topic $\times$ Experience square -0.015∗∗
Non-AGM Experience 0.0008
Campaign on Same Topic $\times$ Non-AGM Experience -0.011
AGM Experience 0.005 -0.034
(0.020) (0.023)
Campaign on Same Topic $\times$ AGM Experience 0.048 0.091∗∗∗
(0.026) (0.032)
AGM Experience square 0.010∗∗
Campaign on Same Topic $\times$ AGM Experience square -0.013∗∗
Campaign on Same Topic $\times$ Google Search Popularity 0.071
Campaign on Same Topic $\times$ Google Search Popularity square -0.021
Fixed-effects
CUSIP x Year
NGO x Year
Topic x Month-Year
Observations 36,543 35,778 36,543 33,142
R2 0.61894 0.61698 0.61995 0.62078
*–p<0.1<0.1; **–p<0.05<0.05; ***–p<0.01<0.01.
  • Note: This table reports the estimated coefficients from (7) in Column 1. In Column 2, we also consider the number of campaigns the NGO has launched in the past which were not on the firm’s AGM day (Non-AGM Experience) and the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (AGM Experience), and their interaction with Campaign on Same Topic. In Column 3, we focus on the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (AGM Experience), its square, and their interaction with Campaign on Same Topic. In Column 4, we replace Experience from model 7 by NGO $g$’s average Google Trends Score relative to GreenPeace the year before the campaign. All regressions also include firm-by-year, NGO-by-year, topic-by-month-by-year fixed effects. Clustered (firm) standard errors are reported in parentheses.
Column (1) of Table 1 shows that campaigning on the same cause as a shareholder proposal has little effect on its own ($\beta_{1}\approx 0$). What matters is the interaction: proposals are more likely to succeed when the campaigning NGO has accumulated experience ($\beta_{3}>0$). A one-standard-deviation increase in past campaigns on the same cause raises proposal success by about 8% on average.
Column (2) distinguishes between campaigns held on AGM dates and those held at other times. Only the former contribute to proposal success. In fact, one-standard-deviation more AGM campaigns increases the probability of success by roughly 4.5 percentage points (about 15%), while non-AGM campaigns have no measurable effect.
Column (3) tests for diminishing returns by including squared terms. The results confirm a concave pattern: additional AGM campaigns continue to improve proposal success, but at a decreasing rate. Importantly, the combined effect remains non-negative even for the most active NGOs. [^8] The marginal effect of AGM Experience and its square never turns negative, so our proxy for reputation does not imply adverse returns.
We corroborate these findings with an alternative proxy: NGO visibility measured by Google Trends. Specifically, Column (4) replaces experience by Google Search Popularityg,y, which we define as the average search score of NGO $g$ relative to Greenpeace in year y1y-1.999The table drops the direct effect of this variable on the dependent variable, as it is absorbed by NGO-by-Year fixed effects. The interaction between this measure and campaigning on the same cause is positive, with a concave shape. Thus, both proxies—past AGM campaigns and Google visibility—suggest that reputation enhances the effectiveness of NGO campaigns, albeit with diminishing marginal returns.
These results also resonate with the media evidence in Section 4: AGM targeting increases negative media coverage only for relatively inexperienced NGOs. More seasoned NGOs gain little additional visibility from campaigning exactly on AGM dates and may be more effective when campaigning slightly earlier. These results support our view that:
Result 5.
Targeting AGMs builds NGO reputation and increases their ability to raise the supported of related shareholder proposals, but with diminishing returns.
7 Discussion and Conclusion
Examining NGO campaign timing reveals that not just the decision to intervene, but its timing, can materially affect firm behavior and shape the course of future activism. Over time, NGOs shift from visibility-driven exposure tactics to influence current strategies, highlighting how the temporal dimension of their strategy channels pressure that leads firms to internalize social and environmental externalities.
We first show that NGOs tend to launch their campaigns on the AGM date, likely motivated by the higher media coverage surrounding these events (Result 1). By doing so, NGOs elicit reactions from key stakeholders, such as shareholders. However, the immediate purpose of these campaigns is primarily to gain visibility by voicing stakeholder concerns; they have no direct effect on firms’ current E&S scores, which improve only when campaigns are launched ahead of the AGM (Result 2).
Hence, campaigning before the AGM offers direct advantages. When an NGO targets a firm on an issue for which a shareholder proposal has already been filed, it boosts the proposal’s chance of success—often leading to its withdrawal from the ballot—and raises average shareholder support by 8% (Result 3). Given the typically low baseline support for such proposals, early NGO interventions can effectively discipline managers who must contend with heightened shareholder pressure (Levit and Malenko, 2011). Importantly, all analyses include NGO-by-year and cause-by-year fixed effects, thereby accounting for time-varying confounders at both the NGO level—such as experience or reputation—and the cause level—such as specific news events. This ensures that globally renowned NGOs (e.g., PETA) do not drive the results. Furthermore, the analyses in Sections 4 and 6 show little evidence of selection on observables, alleviating concerns that NGOs disproportionately target firms that are easier to influence or more likely to generate visibility.
Our results suggest that NGOs face a trade-off, which they navigate by adapting their strategies over time. NGOs with limited activism histories (or low visibility) tend to campaign on the AGM date, whereas more established NGOs prefer targeting firms just before the AGM to directly influence shareholder votes (Result 4). One interpretation is that as NGOs gain more experience with campaigning, they accumulate media exposure, reducing the pressure to campaign specifically on the AGM date.
Therefore, repeatedly targeting AGMs may yield diminishing returns (Result 5). First, an NGO may have already signaled its commitment to a cause, making subsequent signals weaker and reducing the effectiveness of AGM-date campaigning as a signaling device. This interpretation aligns with the view that AGM-date campaigns enhance an NGO’s “reputation.” For instance, following “reputation-for-toughness” models, the NGO may accept an opportunity cost—delaying support for a shareholder proposal—to launch a campaign at a time when it imposes the greatest costs on a firm, thereby building a reputation that can be leveraged with other targets in the future. [^10] In these models, incumbent firms may incur short-term losses by setting predatory prices upon entry of a rival to deter future entrants (Kreps and Wilson, 1982; Milgrom and Roberts, 1982).
Second, stakeholders may come to expect more constructive approaches from established NGOs, such as supporting related shareholder proposals. Consistent with this hypothesis, AGM-date campaigns by mature NGOs—those with multiple prior campaigns—do not generate more negative media coverage than campaigns at other times. By contrast, the very same NGOs do so in their initial campaigns (compare the blue dots and red triangles in Appendix Figure B10). Hence, we observe not only a shift in strategy as NGOs mature (Figure 4) but also a change in their capacity to inflict reputational harm on firms.
Other interpretations are unlikely to be relevant unless they covary with the number of AGM-date campaigns within a given year, NGO, firm, and cause. For instance, organizational aging (e.g., older NGOs becoming larger and more complex) is accounted for through fixed effects. These interpretations open avenues for future research, contributing to the broader discussion on civil society lobbying and activism (Bertrand et al., 2021, 2023).
In conclusion, NGOs influence corporate behavior, yet their incentives and strategies remain understudied. We leverage the exogeneity of corporate events to uncover a trade-off between visibility and influence that NGOs face when deciding when to campaign, and show that how this trade-off is resolved determines which stakeholders are mobilized (e.g., shareholders today versus shareholders tomorrow). Hence, NGOs act as a coordination device, with distinct implications for the social and environmental impact of targeted firms.
We also acknowledge that NGOs not only campaign but also collaborate with corporations (Dahan et al., 2010). Future research could capitalize on similar exogenous variation to quantitatively study the incentives for collaboration.
[…]
Appendix A Omitted Tables
Table A1: Summary Statistics
Panel a: Variables varying by NGO
Mean 25% 50% 75% 95%
Number of campaigns by NGO 3.20 1 1 3 13
     – Campaigns on Env. topic 1.51 0 1 1 5
     – Campaigns on Soc. topic 1.70 0 1 1 6
     – On AGM date 0.24 0 0 0 1
Campaign sentiment -1.72 -2 -2 -1 -1
Panel b: Variables varying by target firm
Mean 25% 50% 75% SD
Market Cap (USD bn) 175.66 21.71 75.05 209.47 263.66
Total Assets (USD bn) 98.81 21.80 59.27 138.68 124.36
Yearly Sales (USD bn) 37.26 6.85 17.35 42.41 150.58
Yearly Number of News 1,183 328 616 1,046 1,695.25
Environmental score 5.66 4 5.58 7.3 2.22
Social score 4.44 3.30 4.70 5.60 1.58
Campaign on Env. topic 0.47 0 0 1 0.50
Campaign on Soc. topic 0.53 0 1 1 0.50
Panel c: Variables varying by target firm and NGO experience
NGO w/ AGM Experience NGO w/o AGM Experience Test difference
Mean S.E. Mean S.E. p-value
Market Cap (decile) 9.23 0.10 9.07 0.04 0.13
Total Assets (decile) 9.26 0.08 9.23 0.04 0.72
Sales (decile) 9.31 0.09 9.14 0.04 0.07
Yearly Number of News (lag) 1,132.37 114.74 1,368.67 47.67 0.06
Environmental score 5.64 0.12 5.72 0.06 0.58
Social score 4.44 0.09 4.45 0.04 0.89
Campaign on Env. topic 0.46 0.03 0.5 0.01 0.15
Campaign on Soc. topic 0.54 0.03 0.49 0.01 0.12
Panel d: OLS Regression of $z_{ny}=\beta_{0}+\beta_{1}\cdot\textit{NGO Experience}_{g(n)y}+\varepsilon_{ny}$
$\hat{\beta}_{1}$ (SE) w/o FEs $\hat{\beta}_{1}$ (SE) w/ FEs N
Dependent Variable ($z_{ny}$): (1) (2) (3) (4)
Market Cap Decile -0.002 (0.003) 0.003 (0.016) 1,338
Total Assets Decile 0.003 (0.004) -0.002 (0.012) 1,338
Sales Decile 0.001 (0.004) -0.002 (0.03) 1,337
Yearly Number of News 0.996 (12.4) -4.12 (16.5) 1,323
Environmental score 0.004 (0.009) 0.041 (0.041) 1,338
Social score -0.0006 (0.006) 0.019 (0.032) 1,338
Campaign on Env. topic 0.006∗∗∗ (0.002) 0.021∗∗ (0.010) 1,338
Campaign on Soc. topic -0.006∗∗∗ (0.002) -0.018 (0.012) 1,338
[…]
Appendix B Omitted Figures
Figure B1: Persistence of the AGM month over time
Note: This figure plots the percentage of AGMs happening in each month of the year (y-axis), given the month of the previous year’s AGM (x-axis). A larger and darker circle indicates a greater proportion of meetings. Plot based on 38,471 meetings of 5,745 US companies, from 2010 to 2020.
Figure B2: Firms’ negative media coverage increases if NGOs campaign during AGMs
Note: The figure plots the coefficients $\beta_{\tau}$ from the following event study regression where we re-calendarize time around an AGM date so that the AGM month is month 0 in each year for each firm:
$News_{n,w,y}$ $=\sumop\displaylimits_{\begin{subarray}{c}\tau=-5\\ \tau\neq-4\end{subarray}}^{5}\!\left(\beta_{\tau}\cdot 1\!\left\{M_{n,w,y}=\tau\!\right\}\cdot NGO\,Campaign_{n,w,y}\!\right)$
$\qquad+X_{n,w,y}\gamma^{\prime}+\alpha_{n,y}+\alpha_{w,y,s(n)}+u_{n,w,y},$
where $News_{n,w,y}$ is an indicator for whether firm $n$ in week $w$ of year $y$ experiences at least one negative E&S news event. The set of dummies $1\{M_{n,w,y}=\tau\}$ indicate the distance in months from the closest AGM. The function $M_{n,w,y}$ assigns week $w$ to each month; the indicator is 1 if the week falls in month $\tau$, where a month is made up of 4 weeks, and zero otherwise. We interact these event-study indicators, with the variable $NGO\,Campaign_{n,w,y}$, an indicator variable equal to 1 if firm n experienced an NGO campaign in week $w$ of year $y$, and 0 otherwise. The vector $X_{n,w,y}$ includes the direct effect of these variables. The reference month is $\tau=-4$ following SEC rules on shareholder proposals. Note that, because of the stability of AGM dates across years within firms (Appendix Figure B1), the distribution of months fixed effects is unaffected by the re-calendarization, so that months are not double counted across years. Standard errors are clustered by firm.
Figure B3: NGO campaigns and shareholder proposals
Note: This figure plots the estimated coefficients of the regression of a dummy indicating the occurrence of a shareholder proposal on month $t$ at firm $f$, relative to an NGO campaign targeting firm $f$ on the same cause, $k$ months ahead. The regression also includes campaign and month-by-year-by-industry-by-topic fixed effects. The regression is based on 1,227 NGO campaigns. Standard errors are clustered at the firm level.
Figure B4: NGO campaign timing and consumer backlash
(a) NGO Campaigns on AGM dates
(b) NGO Campaigns on Other Days
Notes: The figure plots the estimated ${\beta}_{\tau}$ from (3). Panel (a) focuses on NGO campaigns on AGM dates, while Panel (b) focuses on campaigns on dates other than AGM dates. The reference period is -1 quarter, consistent with the other analyses. The number of observations is 239,818 in Panel (a) and 239,925 in Panel (b). Standard errors are clustered at the firm level.
Figure B5: Campaign Timing and E&S Score Improvement
Note: This figure plots the estimated coefficients of the following regression:
$_{5}\text{E\&S Score}_{n,q}=\sumop\displaylimits_{l=-6}^{5}\text{Campaign}_{n,q,l}+X_{n,q}\gamma^{\prime}+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$ (B1)
where $_{5}\text{E\&S Score}_{i,q}$ is the change in firm $n$’s average of its Environmental and Social scores between quarter q1q-1 and quarter $q+4$, and $\sumop\displaylimits_{l=-6}^{6}\text{Campaign}_{n,q,l}$ is a set of dummy variables indicating the occurrence of an NGO campaign against firm $n$ in quarter $q$, being $l$ months away from the closest AGM. The regression also controls ($X_{n,q}$) for firm size (logged total assets), lagged average of E and S scores, firm fixed effects ($\alpha_{n}$), and industry-by-quarter fixed effects ($\alpha_{s(n),q}$). ESG scores are from MSCI, NGO campaigns data from Sigwatch, firm characteristics from Compustat, and AGM dates from ISS. The regression has 47,082 observations. Standard errors are clustered at the firm level.
Unlike an event study, equation (B1) is a distributed leads-and-lags regression. We omit a reference period, so the coefficients are interpreted relative to firms that did not experience a campaign. This choice allows us to directly compare the effects of early campaigns with those launched on the AGM date relative to firms without campaigns. Results remain similar if we instead set month –4 as the reference period, as in the other analyses.
Figure B6: NGO experience and the timing of NGO campaigns – Alternative set of Fixed Effects
Note: This figure plots the estimated coefficients from the regressions of the distance between NGO campaigns and the targeted firm’s closest AGM (in months) on dummies indicating the experience of the NGO, measured by the number of campaigns launched before. The regression includes replaces NGO-by-year fixed effects with separate fixed effects for each NGO and each year. Hence the coefficients $\{\hat{\beta}_{\tau}\}_{\tau}$ in (5) is also identified from variation in NGO experience over years. Standard errors are clustered at the NGO level.
Figure B7: Campaign timing for NGOs whose first campaign is after the AGM
Note: This figure plots the estimated coefficients from the regressions (5) of the distance between NGO campaigns and the targeted firm’s closest AGM (in months) on dummies indicating the experience of the NGO, measured by the number of campaigns launched before. The panel focuses on NGOs whose first campaign was within the two weeks following the AGM. Standard errors are clustered at the NGO level.
Figure B8: NGO experience and target firm’s size
Note: This figure plots the estimated coefficients from regressing the target firm’s market cap (in deciles to control for the fact that the tails of the market cap distribution have thickened in recent years) on dummies indicating the experience of the NGO, measured by the number of campaigns it had launched in previous year as in (5). Standard errors are clustered at the firm level. The number of observations is 1,338.
Figure B9: NGO experience and target firm’s E&S characteristics
(a) Target firm’s E score for Environmental campaigns
(b) Target firm’s S score for Social campaigns
Note: This figure plots the estimated coefficients from regressing either a target firm’s environmental score (top panel) or its social score (bottom panel) on dummies indicating the experience of the NGO, measured by either the number of environmental (top panel) or social (bottom panel) campaigns it had launched in previous year as in (5). Standard errors are clustered at the firm level. The number of observations is 636 and 698, respectively.
Figure B10: Media coverage of NGO-target firms around their AGM
Note: To test if media exposure of NGO campaigns is highest during the AGM week, we aggregate negative E&S news by firm ($n$) and week-by-year ($w,y$) and perform the event study:
$\text{News}_{n,w,y}$ $=\sumop\displaylimits_{\begin{subarray}{c}\tau=-5\\ \tau\neq-4\end{subarray}}^{5}\left(\beta_{\tau}\cdot\mathbf{1}_{\{M_{n,w,y}=\tau\}}\cdot\text{NGO Campaign}_{n,w,y}\right)$
$\qquad+X_{n,w,y}\gamma^{\prime}+\alpha_{n,y}+\alpha_{w,y,s(n)}+u_{n,w,y},$
[…]
Appendix C Robustness of Google Trends Results
This section discusses the robustness of the results presented in Figure 2 to alternative specifications. In particular, we replicate the analysis using (i) a staggered difference-in-differences estimator (Sun and Abraham, 2021), which accounts for heterogeneous campaign timing across NGOs, and (ii) a distributed lags specification, which controls for the possibility of overlapping campaigns within the same year.
C.1 Staggered Difference-in-Differences Specification
We first replicate Figure 2, applying the staggered difference-in-differences approach to account for heterogeneous campaign timing across NGOs (Sun and Abraham, 2021).
Estimated coefficients $\beta_{\tau}^{AGM}$ are reported in Figure C1. The results are consistent with our baseline findings: AGM-timed campaigns (blue dots) generate sizeable visibility gains for early campaigns, raising Google search intensity by about 58%. By contrast, experienced NGOs see a more modest increase (33%) in search intensity, and only in the month of the campaign.
Figure C1: Staggered difference-in-differences specification
Note: The Figure plots the estimated $\beta^{AGM}$ coefficients from (1) using a staggered difference-in-differences estimator (Sun and Abraham, 2021). The dependent variable is the monthly Google Trends score for the NGO. Blue dots are estimated by restricting the sample to an NGO’s first three campaigns, while red triangles consider later campaigns. Standard errors are clustered at the campaign level.
C.2 Distributed Lags Specification
In the baseline event-study, if an NGO launches several campaigns in close succession, their effects may overlap in calendar time. For example, if one campaign takes place at $t=0$ and another at $t=2$, the months around $t=2$ capture the lingering impact of both events. This overlap can bias the estimated effects in the previous analyses. To address this, we estimate a distributed lags specification that explicitly includes leads and lags of both AGM and non-AGM campaigns:
$\text{Google Trends Score}_{g,m,y}$ $=\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-5\end{subarray}}\left(\beta_{\tau}^{AGM}\cdot\text{AGM Campaign}_{g,m+\tau}\right)$ (C1)
$\quad+\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-5\end{subarray}}\left(\beta_{\tau}^{Non-AGM}\cdot\text{Non-AGM Campaign}_{g,m+\tau}\right)$
$\quad+\alpha_{m,y}+\alpha_{g,y}+u_{g,m,y},$
By doing so, the regression attributes observed variation in search activity in each month to the full sequence of campaigns occurring in that window, rather than assigning the entire effect to whichever campaign happens to be closest in event time. Put differently, the distributed lags approach disentangles the impact of overlapping campaigns by allowing each campaign to contribute separately to visibility over several months. This ensures that the estimated $\beta^{AGM}_{\tau}$ captures the marginal visibility effect of a single campaign, net of any other campaigns that may have occurred nearby.
Figure C2: Distributed lags specification
(a) First three campaigns
(b) Later campaigns
Note: The two plots present the estimated $\beta^{AGM}$ with blue circles and $\beta^{Non-AGM}$ with red triangles coefficients from (C1) on two different samples. The first panel subsets the data to each NGO’s first three campaigns, while the second panel examines later campaings. The dependent variable is the monthly Google Trends score for the NGO. Standard errors are clustered at the campaign level.
Figure C2 reports the estimated coefficients. When restricting attention to an NGO’s first three campaigns, the results confirm that AGM-timed campaigns lead to significant short-run increases in visibility (blue dots), while later campaigns generate little or no effect. Whether early or late, non-AGM campaings have a limited effect on an NGO’s visibility.
Overall, these robustness checks confirm that our main results are not driven by campaign timing heterogeneity across NGOs or by overlapping campaigns within a year.
Appendix D Matching Proposals And NGO Campaigns
This section outlines the process of matching each shareholder proposal in the ISS dataset with a relevant NGO campaign cause.
Sigwatch assigns a“Generic Topic” to each campaign in their dataset. To streamline the analysis, we combine certain causes when they are difficult to distinguish or when one category has few campaigns and can reasonably fit within a broader category. Specifically, we merge “Labor Rights” with “Human & Labor Rights,” “Children” with ”Human & Labor Rights,” “CSR / ESG” with “CSR,” and “Biodiversity” and “Wildlife” with ”Environment.” For campaigns under the “Safety” cause, which includes only 12 entries, we examine each campaign’s specific purpose and classify it as either “Product Safety” or “Food Safety.”
Using these consolidated “Generic Cause Names”, we then categorize each shareholder proposal based on its short description. When a shareholder proposal is submitted, it includes a brief, one-sentence description, which is provided in the ISS dataset and serves as the foundation for cause classification. We start by assigning proposals to causes based on keywords in their descriptions and then manually classify any remaining proposals.
Table D1: Proposal Causes
NGO Campaign Category Example of Shareholder Proposals
Topic Classification (E or S)
Climate Change E Encourage and Support Carbon Tax to Address Climate Change
Environment E Stop financing mountain top removal coal mining
Pollution E Report on Efforts to Reduce Pollution From Products and Operations
Rainforest E Report on impact on rainforest
Sustainability E Establish board committee on sustainability
Waste Generation E Do not increase radioactive waste production
Water E Report on supply chain water impacts
Animal Rights & Welfare S Review suppliers’ animal slaughter methods
CSR S Disclose Charitable Contributions
Food Safety S Label Genetically Modified Ingredients
Health S Reduce Tobacco Harm to Health
Human & Labor Rights S End Sri Lanka Trade Partnerships Until Human Rights Improve
Indigenous People S Adopt Global Policy Regarding the Rights of Indigenous People
Intellectual Property Rights S Implement a Process to Fully Vet Whether a
Supervisor is a Bona Fide Inventor Before a Patent is Filed
Politics S Affirm Political Nonpartisanship
Product Safety S Report on Nanomaterial Product Safety
Worker Safety S Report on Reducing Risks of Accidents
  • Note: This Table reports the 17 topics assigned to each shareholder proposal in the sample. Shareholder proposals data from ISS. The topics are from Sigwatch’s categorization of campaigns.