### v8

# NGO Activism:

Abstract
This paper studies how the timing of NGO activism shapes its effectiveness in influencing corporate behavior. Using data on 2,500 campaigns targeting U.S. firms, we show that campaigns timed at annual general meetings (AGMs) generate large visibility gains but little contemporaneous influence, while campaigns launched before the AGM significantly increase shareholder proposal success and improve firms’ environmental and social performance. We develop a dynamic model in which NGOs trade off awareness building and credibility formation, generating a lifecycle in activism from visibility-seeking to influence-oriented engagement. Therefore, NGOs’ objectives evolve endogenously to coordinate stakeholder pressure and shape corporate behavior.
JEL classifications: L21, L31
Keywords: Non-Governmental Organizations, NGO campaigns, social and environmental responsibility, ESG, firm dynamics, non-profit objectives
1 Introduction
Non-governmental organizations (NGOs) have become influential actors in shaping firms’ environmental and social policies (e.g., baron2001private; egorov2017private; yaziji2009ngos). Yet, we know remarkably little about their strategic behavior. NGOs are highly opaque: they disclose almost nothing about internal decisions, incentives, or tactics, making most of their actions effectively unobservable (maxwell2012economic). As a result, the literature largely portrays NGOs either as principled advocates driven by social missions (e.g., besley2005competition; benabou2006incentives; daubanes2019rise) or as organizations vulnerable to strategic misuse and capture (e.g., grossman1994protection; bertrand2020tax; akey2021hacking). We argue instead that NGOs respond systematically to incentives generated by stakeholder attention and expectations, shaping NGOs’ strategic choices, with consequences for corporate behavior and social and environmental outcomes.
Figure 1: Timing of NGO campaigns and media attention around AGMs
(a) Campaigns peak during AGMs
(b) Media coverage rises during AGMs
Note: Panel (a) displays the number of NGO campaigns launched relative to the date of the Annual General Meeting of the targeted firm (week 0), based on 1,544 NGO campaigns. Panel (b) compares the weekly average count of news during the AGM week (gray bars) and outside of the AGM week (white bars), based on weekly observations for 5,211 US firms. The NGO data come from Sigwatch; the news data are from Ravenpack.
A rare opportunity to observe these strategic responses arises from campaign timing. NGO campaigns must be public to be effective, and their timing provides one of the few visible choices NGOs make. We focus on how NGOs align campaigns with a moment when firms are uniquely exposed to outside scrutiny: the Annual General Meeting (AGM). AGMs are predictable focal events at which executives report results, shareholders vote, and media attention peaks (fioretti2024shared). Panel (a) of Figure 1 shows that NGO campaigns are five to six times more likely to occur in AGM weeks than at other times, and Panel (b) shows that media coverage of targeted firms peaks at the same moment. These patterns raise a central question: do NGOs target AGMs primarily to maximize exposure, or do they use campaign timing to exert influence on firms? Is there a tradeoff between exposure and influence? More broadly, what incentives shape NGO strategy, and how do their choices affect the firms they target?
Leveraging the AGM as a predictable focal point, we address these questions and uncover a lifecycle in NGO activism. Younger NGOs concentrate on AGM-day campaigns that build visibility, while more established NGOs shift to pre-AGM campaigns that influence shareholder decisions and firms’ environmental and social performance. A simple dynamic model, in which NGOs trade off awareness building against credibility formation, rationalizes this timing pattern. Taken together, our results show that NGO behavior is governed by incentives that determine which stakeholders are mobilized and when, and thus whether campaigns generate symbolic exposure or substantive corporate change. In particular, because most shareholder votes are cast before the AGM, campaigns timed on the meeting date primarily affect visibility, whereas campaigns launched earlier can influence voting outcomes – linking NGO effectiveness to institutional features such as AGM and proxy rules and disclosure timing.
Imagine an NGO uncovers troubling information about a firm and seeks to change its behavior. Because Annual General Meetings (AGMs) occur at predictable times, the NGO can choose whether to campaign before the AGM—when shareholder votes and negotiations are still open—or at the AGM itself, when public and media attention peak but most proxy votes have already been cast. If NGOs simply campaigned as soon as concerns arose, we would not expect the sharp clustering of campaigns around AGMs shown in Figure 1. The pronounced spike instead suggests that campaign timing is a strategic choice.
To study this choice, we assemble a novel dataset of roughly 2,500 campaigns by international NGOs targeting U.S. firms, linked to media coverage, Google search activity, NGO financial statements, shareholder proposals, and environmental and social (E&S) outcomes. SEC rules tie proxy filings and AGM dates to the fiscal year, so each firm’s AGM recurs within a narrow window each year, creating a predictable surge in visibility alongside a fixed deadline for shareholder voting. This institutional structure generates sharp variation in when campaigns occur relative to when shareholder decisions are made, which we exploit to distinguish exposure-oriented from influence-oriented activism and their drivers.
Our first set of results shows that campaign timing is tightly linked to how NGOs affect firms. In their first campaigns, NGOs tend to target the AGM date, generating large visibility shocks – raising media coverage, online search activity, donations, and the likelihood that related shareholder proposals are filed at the next AGM – but having little effect on contemporaneous proposal success or firms’ E&S performance. This pattern suggests that shaming campaigns alone may have limited influence on corporate decisions.
As NGOs gain experience and reputation, campaigns shift into the months before the AGM, when proxy voting and negotiations are still open. These earlier actions significantly increase the success and vote share of E&S proposals and are followed by improvements in firms’ E&S scores, consistent with genuine changes in corporate policy rather than pure reputational damage.
These patterns are not explained by NGOs gradually moving from easier to harder targets: observable characteristics of target firms do not change systematically as NGOs accumulate campaigns. To address selection on unobservables, we exploit quasi-exogenous variation from the publication date of the Fashion Transparency Index, which is fixed by the Rana Plaza anniversary and thus pins down both when campaigns occur and which firms are targeted. Using the distance between index publication and firms’ AGM dates as a source of exogeneous timing variation, we find similar effects both in signs and magnitudes to those estimated in the whole dataset.
We then show that NGOs’ use of AGM visibility versus pre-AGM influence evolves systematically over their organizational life cycle. NGOs with little prior AGM exposure or public visibility concentrate campaigns on the AGM date, where media attention peaks and visibility gains are largest. As NGOs accumulate visibility – proxied by past AGM campaigns, search intensity, and organizational size – they progressively shift campaigns earlier in the AGM cycle, when shareholder votes have not yet been cast and managerial concessions are more likely.
NGO income statement data further indicate that early campaigns are costlier and less visibility-oriented than AGM-day campaigns. For younger NGOs, AGM-day actions generate sizable short-run increases in donations and returns on assets, consistent with shaming campaigns being relatively inexpensive and effective visibility tools. For more established NGOs, the marginal fundraising benefit of AGM campaigns is small, and non-AGM campaigns do not produce comparable financial gains for any group.
The empirical patterns point to a sharp distinction between campaigns launched before the AGM and those timed on the meeting date. Campaigns before the AGM are associated with greater influence on shareholder votes and firm outcomes, while AGM-day campaigns primarily generate visibility. We interpret these findings through a dynamic model in which an NGO, shareholders, and the firm interact repeatedly over time, and in which campaign timing is the NGO’s central strategic choice.
In each period, the public state is summarized by two variables. The first is an awareness stock, which captures how salient the issue is among dispersed shareholders and the broader public. Awareness increases with highly visible actions—most notably campaigns timed on the AGM—and decays over time in the absence of renewed attention. The second is a credibility belief, which reflects stakeholders’ assessment of the NGO’s resolve and effectiveness in sustaining pressure on firms. This belief evolves endogenously as shareholders and firms observe the NGO’s actions and their outcomes.
Given the current state, the NGO chooses whether to campaign before shareholder votes are cast or at the AGM. AGM-day campaigns maximize visibility and generate warm-glow benefits for the NGO, but they occur after most proxy votes are locked in and therefore have little contemporaneous influence on firm decisions. Early campaigns, by contrast, are less visible and forego some warm-glow benefits, but they can affect shareholder voting and negotiations with the firm. Importantly, early campaigns are also more informative: they require sustained engagement and, when successful, reveal persistence and competence, strengthening the NGO’s credibility. AGM-day campaigns are comparatively cheap and visible, but once an NGO is established they convey little new information about its type.
Shareholders observe campaign timing and, when relevant, campaign outcomes. Based on the current levels of awareness and credibility, they decide whether to file a proposal and how much support to give it. Higher awareness increases the likelihood that shareholders notice and coordinate around an issue, while higher credibility raises the expected effectiveness of activism and the probability of success. The firm, anticipating these responses, adjusts its behavior accordingly.
This dynamic interaction yields a simple equilibrium structure. When credibility and awareness are low, the NGO optimally campaigns at the AGM: visibility and warm-glow benefits dominate, and early engagement is unlikely to be persuasive. As awareness accumulates and credibility improves, the NGO crosses a threshold at which the influence and reputational gains from early campaigning outweigh the loss in visibility. At that point, it switches to pre-AGM campaigns to exert direct pressure on shareholder decisions and firm policies. This cutoff behavior reflects evolving stakeholder expectations: actions that initially signal resolve later appear routine, leading established NGOs to rely less on public shaming and more on early, substantive engagement (baur2011moral; mitchell2017reputations).
A purely visibility-driven model would predict persistent AGM-day campaigning. Matching the observed shift from late, high-visibility actions to earlier, influence-oriented engagement instead requires a dynamic reputation-building mechanism in the spirit of intertemporal signaling models (kreps1982reputation; milgrom1982limit; cabral2010dynamics), combined with a deterministic awareness stock that separates visibility-driven exposure from credibility-driven influence. This framework provides a unified explanation for why NGOs initially seek attention and later convert reputation into corporate impact.
Our analysis yields three main contributions. First, we offer a new quantitative framework to examine NGO strategy and to infer how NGOs respond to incentives, which stakeholders they prioritize at different stages, and how their objectives evolve. Classic corporate-finance approaches tie organizational objectives to ownership structure (e.g., grossman1977value; grossman1979theory), while stakeholder theory emphasizes salience based on power, legitimacy, and urgency (freeman1984strategic; mitchell1997toward). Our results show that NGO objectives shift with how donors and investors interpret NGOs’ actions and with credibility concerns about timing. In this sense, we provide a quantitative setting to study how organizational objectives emerge from interactions with multiple stakeholders rather than being fixed by a single principal—for example, an entrepreneur defining the firm’s goals (e.g., scott2002love; romer2006firms; duarte2025competitive; fioretti2024shared). [^1] Our approach also relates to work on the strategic timing of actions around news cycles and focal dates (eisensee2007news; madestam2013do; durante2018attack; djourelova2022media), but shifts the focus to NGOs in the corporate domain and to the management of reputational capital over time.
Second, we contribute to the corporate-finance and stakeholder-governance literatures by identifying mechanisms through which firms partially internalize social and environmental externalities (fioretti2022caring; allcott2022economic). Our setting complements work on shareholder activism and gadflies (e.g., brav2008hedge; gantchev2021costs; malenko2024information): NGOs do not vote but act as external coordinators that put issues on the ballot and help align dispersed investors’ behavior, shaping both the filing and success of shareholder proposals (see also chuah2024shareholder; mazet2023some; mcgahan2023new).
Third, our approach has implications for policy. For boards, early, credibility‐backed campaigns are informative signals that warrant engagement rather than dismissal (dimson2015active; kolbel2017media). For investors and stewardship codes, credible NGOs supply information that helps coordinate dispersed shareholders on E&S issues, complementing traditional voice and exit (oehmke2020theory; gans2021exit; broccardo2020exit; saint2024exit; bonnefon2025moral). For regulators, rules on AGM timing, proposal access, and disclosure shape the returns to NGO campaigns and thus the scope for private activism to substitute for, or complement, formal regulation.
The paper proceeds as follows. Section 2 presents the model. Section 3 describes the data, and Section 4 outlines the empirical strategy. Sections 5 and 6 examine NGO campaigns as tools for gaining media exposure and influencing firms through key stakeholders. Section 7 brings these channels together by analyzing how NGO reputation enhances the success of shareholder proposals, and Section 8 provides evidence from the quasi-experimental case study of the Fashion Transparency Index. Section 9 discusses implications for policy and future research, and Section 10 concludes.
2 Model
This section develops a model in which the timing of an NGO campaign conveys information about the NGO’s credibility and shifts the salience of the underlying issue among shareholders. Two forces interact: warm glow, which makes high‑visibility AGM campaigns intrinsically valuable, and reputation, which evolves by Bayesian learning as timing and proposal outcomes are observed.
2.1 Environment and Timing
Time is discrete, $t=0,1,2,\ldots$. The public state at the start of period $t$ is $(A_{t},\pi_{t})$, where $A_{t}$ denotes the stock of awareness, or issue salience among dispersed shareholders on the topic championed by the NGO, and $\pi_{t}\in[0,1]$ is the belief that the NGO is the credible type ($\theta=C$) rather than the weak one ($\theta=W$). A credible NGO can coordinate investors, sustain pressure, and translate outreach into voting support, whereas a weak NGO lacks this capacity and therefore exerts little direct influence on the target firm.
Timeline.
Given $(A_{t},\pi_{t})$, this is the order of play in each period $t$:
  1. 1.
    Filing decision. A (representative) shareholder decides whether to file a proposal for the AGM that takes place at the end of period $t$. A proposal is filed if the perceived probability of success covers a fixed cost $f$.
  2. 2.
    Campaign timing decision. The NGO chooses $d_{t}\in\{\mathrm{early},\mathrm{AGM}\}$. Early campaigns occur before proxy votes are locked; AGM campaigns occur when most votes are already cast.
  3. 3.
    Outcome. If a proposal was filed, the proposal is voted and an outcome $y_{t}\in\{0,1\}$ is realized at the end of period $t$.
  4. 4.
    Updating. Observers see $d_{t}$ and, if a proposal was on the ballot, $y_{t}$. Awareness and beliefs update to $(A_{t+1},\pi_{t+1})$.
Awareness law of motion.
The awareness stock $A_{t}$ summarizes issue salience among dispersed shareholders and evolves according to
$A_{t+1}=\rho\cdot A_{t}+(1-\rho)\cdot v(d_{t}),\qquad A_{0}=0,\quad\rho\in(0,1),\quad\text{ and }\quad v(\text{AGM})>v(\text{early}),$ (1)
so that awareness is a convex combination between the visibility of the chosen timing, $v(d_{t})$, and past awareness. The function $v(d_{t})$ captures the visibility of the issue–firm link among shareholders: the condition $v(\text{AGM})>v(\text{early})$ implies that awareness increases when the NGO campaign aligns with the AGM, when firms are most exposed to media and investor scrutiny, while the others are regularity conditions.
NGO preferences.
Warm glow from visibility and contemporaneous influence on shareholder outcomes jointly determine one-period payoffs. The per-period utility is
$u(d_{t};\pi_{t},A_{t})=\alpha\cdot v(d_{t})+B\cdot{}_{d_{t}}(\pi_{t},A_{t})-C,$ (2)
[…]
$_{\mathrm{early}}(\pi,A)>{}_{\mathrm{AGM}}(\pi,A)=0,\qquad\frac{\partial{}_{\mathrm{early}}}{\partial\pi}(\pi,A)>0,\qquad\frac{\partial{}_{\mathrm{early}}}{\partial A}(\pi,A)\geq 0.$ (3)
Hence, AGM influence is normalized to zero because most shareholders have already cast their votes, leaving little scope for NGOs to affect outcomes. By contrast, early campaigns have greater influence: higher credibility $\pi$ strengthens their effect since credible NGOs exert greater pressure on firms and investors, whereas awareness $A$ enhances influence indirectly by broadening the audience. The effect of awareness is therefore weak and may disappear once the issue is already widely known (i.e., $\frac{\partial^{2}{}_{\text{early}}}{\partial A^{2}}\leq 0$).
[…]
Success probability and realized outcome.
When a shareholder proposal is on the ballot, its likelihood of passing depends on three factors: how aware shareholders are of the issue, how credible they perceive the NGO to be, and whether the campaign occurs early enough to influence votes. We model the probability of a successful proposal for an NGO of type $\theta$ as a logistic function,
$S^{\theta}(A_{t},\pi_{t},d_{t})=\frac{\exp\big(h_{0}+h_{A}A_{t}+h_{\pi}\pi_{t}+h\,{}_{d_{t}}^{\theta}(A_{t},\pi_{t})\big)}{1+\exp\big(h_{0}+h_{A}A_{t}+h_{\pi}\pi_{t}+h\,{}_{d_{t}}^{\theta}(A_{t},\pi_{t})\big)},$ (4)
where $h_{0}$ is a baseline term, $h_{A}>0$ measures how greater awareness $A_{t}$ increases the baseline support among shareholders, $h_{\pi}>0$ captures the effect of credibility $\pi_{t}$ on the likelihood of success, and $h>0$ scales the influence due to the campaign’s timing.
[…]
The realized proposal outcome is the only random element in the model:
$y_{t}=\begin{cases}1,&\text{with probability }S(A_{t},\pi_{t},d_{t}),\\[3.0pt] 0,&\text{with probability }1-S(A_{t},\pi_{t},d_{t}).\end{cases}$ (5)
Thus, $y_{t}=1$ represents a successful proposal – one that either passes or achieves a favorable settlement before the AGM – and occurs more frequently when campaigns are early, credibility is high, or awareness is widespread. In contrast, unsuccessful proposals ($y_{t}=0$) provide weaker evidence of toughness and slow reputation accumulation.
Filing rule.
Before the AGM takes place, shareholders decide whether to file a proposal. Filing is worthwhile when the expected benefit of passage exceeds the fixed cost $f>0$. Let $V_{P}>0$ denote the value of a successful proposal to a shareholder, and let $\hat{d}_{t}$ denote the NGO’s anticipated timing implied by its policy. Because timing is a deterministic function of the state $(A_{t},\pi_{t})$, shareholders can forecast it. A proposal is filed if and only if
$V_{P}\cdot S(A_{t},\pi_{t},\hat{d}_{t})\geq f,$ (6)
where since the shareholder does not know the true type $\theta$, $S(\cdot)$ is the perceived success probability, a belief-weighted average of the type-specific probabilities $S^{C}(\cdot)$ and $S^{W}(\cdot)$ in (4), i.e., $S(\cdot)\ =\ \pi_{t}\cdot S^{C}(\cdot)+(1-\pi_{t})\cdot S^{W}(\cdot)$.
This rule makes filing completely determined by the current state: greater awareness and credibility increase $S(A_{t},\pi_{t},\hat{d}_{t})$ and therefore the likelihood that filing is worthwhile. AGM timing builds awareness and encourages future filing, whereas early timing boosts contemporaneous success without necessarily raising next-period awareness.
2.2 Belief Updating
After observing the NGO’s campaign timing $d_{t}$ and the outcome $y_{t}$ of the shareholder proposal, shareholders update their belief $\pi_{t+1}$ about the NGO’s credibility. The posterior depends on two sources of information: the timing choice itself, which is strategically informative, and the realized outcome, which is stochastic but type dependent.
Let $p^{C}(d_{t},y_{t}\mid\pi_{t},A_{t})$ and $p^{W}(d_{t},y_{t}\mid\pi_{t},A_{t})$ denote the likelihood of observing $(d_{t},y_{t})$ when the NGO is of the credible type ($C$) or weak type ($W$), respectively. By Bayes’ rule, the posterior belief that the NGO is credible is
$\pi_{t+1}=\frac{\pi_{t}\cdot p^{C}(d_{t},y_{t}\mid\pi_{t},A_{t})}{\pi_{t}\cdot p^{C}(d_{t},y_{t}\mid\pi_{t},A_{t})+(1-\pi_{t})\cdot p^{W}(d_{t},y_{t}\mid\pi_{t},A_{t})}.$
Dividing by $1-\pi_{t+1}$ and taking logs reveals the following belief updating rule, [^2] Note that the joint likelihood factors as $p^{\theta}(d_{t},y_{t}\mid\pi_{t},A_{t})=p^{\theta}(y_{t}\mid d_{t},\pi_{t},A_{t})\cdot p^{\theta}(d_{t}\mid\pi_{t},A_{t})$ for $\theta\in(W,C)$.
$\ln\left(\frac{\pi_{t+1}}{1-\pi_{t+1}}\right)-\ln\left(\frac{\pi_{t}}{1-\pi_{t}}\right)=\ln\frac{p^{C}(d_{t}\mid\pi_{t},A_{t})}{p^{W}(d_{t}\mid\pi_{t},A_{t})}+\ln\frac{p^{C}(y_{t}\mid d_{t},\pi_{t},A_{t})}{p^{W}(y_{t}\mid d_{t},\pi_{t},A_{t})}.$ (7)
The first term on the right-hand side captures what observers can infer from the NGO’s timing choice $d_{t}$ alone. Because the payoff advantage of early campaigning increases with credibility ($\partial{}_{\mathrm{early}}/\partial\pi>0$), a more credible NGO finds early timing relatively more attractive. In equilibrium, therefore, early campaigns occur more often when credibility is high, making timing itself an informative signal. We parameterize the informativeness of this timing signal as
[…]
The second likelihood term in (7) captures the information contained in the realized outcome. Conditional on $(A_{t},\pi_{t},d_{t})$, the probability of proposal success is given by (4). For notational convenience, write $S^{\theta}(A_{t},\pi_{t},d_{t})\equiv S(A_{t},\pi_{t},d_{t}\mid\theta)$ for the success probability under type $\theta$. In the underlying outcome process, early campaigning is more effective for credible NGOs, so that $S^{C}(A_{t},\pi_{t},\mathrm{early})>S^{W}(A_{t},\pi_{t},\mathrm{early})$, while at the AGM, where $_{\mathrm{AGM}}\equiv 0$, both types perform identically and $S^{C}=S^{W}$. Hence, outcomes are informative about credibility only when the campaign occurs early. Given this Bernoulli data-generating process, Appendix A.1 derives the outcome component of the likelihood ratio that drives belief updating as,
$\ln\frac{p^{C}(y_{t}\mid d_{t},A_{t},\pi_{t})}{p^{W}(y_{t}\mid d_{t},A_{t},\pi_{t})}=y_{t}\cdot\ln\frac{S^{C}(A_{t},\pi_{t},d_{t})}{S^{W}(A_{t},\pi_{t},d_{t})}+(1-y_{t})\cdot\ln\frac{1-S^{C}(A_{t},\pi_{t},d_{t})}{1-S^{W}(A_{t},\pi_{t},d_{t})}.$
Since $S^{C}=S^{W}$ at the AGM, outcomes then carry no additional information, and beliefs update only through the timing term. When the campaign is early, $S^{C}>S^{W}$, so a successful proposal ($y_{t}=1$) increases $\pi_{t+1}$, while failure ($y_{t}=0$) lowers it. Intuitively, early campaigns are reputation-building experiments: they can signal credibility, but they also expose the NGO to reputational risk if they fail.
Appendix A.2 shows that posterior beliefs are consistent under standard regularity conditions.
2.3 Dynamic Program
At each date $t$, the NGO chooses its campaign timing to balance the short-run benefit of visibility against the expected influence on shareholder behavior and the evolution of its reputation. Given beliefs $(\pi_{t},A_{t})$, the organization anticipates that an AGM campaign will provide the warm-glow payoff from visibility but no direct effect on voting, while an early campaign will generate a smaller visibility payoff today but can influence votes immediately and improve its reputation through a more informative outcome.
Let $V(\pi_{t},A_{t})$ denote the NGO’s continuation value. An NGO’s recursive problem is
$V(\pi_{t},A_{t})=\max_{d_{t}\in\{\mathrm{early},\mathrm{AGM}\}}\Big\{\alpha\,v(d_{t})+B\,{}_{d_{t}}(\pi_{t},A_{t})-C+\beta\,\mathbb{E}_{y}\big[V(\pi_{t+1},A_{t+1})\mid\pi_{t},A_{t},d_{t}\big]\Big\},$ (8)
where $\beta\in(0,1)$ is the discount factor, and expectations are taken with respect to the only stochastic element in the model – the realized proposal outcome $y_{t}$. Awareness follows (1), while beliefs evolve through Bayesian updating as explained below.
2.4 Equilibrium Timing and Cutoffs
The NGO’s equilibrium policy specifies, for every state $(A_{t},\pi_{t})$, whether it campaigns early or at the AGM. Define the action-value difference
[…]
$\quad+\beta\cdot\big(\mathbb{E}_{y}\big[V(\pi_{t+1},A_{t+1})\mid\pi_{t},A_{t},\mathrm{early}\big]-\mathbb{E}_{y}\big[V(\pi_{t+1},A_{t+1})\mid\pi_{t},A_{t},\mathrm{AGM}\big]\big),$
where expectations are taken over next-period beliefs and awareness determined by (1) and (7). The first term in brackets is the current payoff wedge: it weighs the incremental influence and lower visibility of early campaigning. The second term is the continuation-value wedge: early campaigning improves future payoffs by shifting beliefs and, through reputation, increasing expected influence in later periods.
[…]
Proposition 1 (Separation by Campaign Timing).
There exists a unique cutoff function $\bar{\pi}(A)\in(0,1)$ such that the NGO’s optimal campaign timing is
$d_{t}^{\star}=\begin{cases}\mathrm{AGM},&\pi_{t}<\bar{\pi}(A_{t}),\\ \mathrm{early},&\pi_{t}>\bar{\pi}(A_{t}).\end{cases}$
The cutoff is weakly decreasing in awareness, $\bar{\pi}^{\prime}(A)\leq 0$. Moreover, in equilibrium:
  1. (i)
    For $\pi_{t}<\bar{\pi}(A_{t})$, both credible and weak NGOs pool on AGM campaigns.
  2. (ii)
    For $\pi_{t}>\bar{\pi}(A_{t})$, the credible NGO separates by campaigning early, while the weak NGO continues to campaign at the AGM.
Proof. See Appendix A.3.
The intuition is straightforward. When credibility is low, early campaigning is too costly to be persuasive: stakeholders do not expect weak NGOs to follow through, and AGM visibility provides immediate warm-glow benefits for both types. As a result, both credible and weak NGOs pool on AGM timing. As credibility rises, however, early campaigning becomes informative. For the credible NGO, the influence and reputational benefits of early action outweigh the visibility advantage of AGM campaigns, leading it to switch to early timing. The weak NGO, which derives no direct influence payoff from early engagement ($^{W}_{\mathrm{early}}=0$), does not find this deviation profitable and remains at the AGM.
Importantly, although at very high beliefs early campaigning would be optimal for any NGO in a static sense, only the credible NGO reaches this region in equilibrium. Its reputation is built through successful early engagement, after which timing serves a purely instrumental role. The weak NGO, by contrast, remains trapped in the low-credibility region where visibility dominates and continues to rely on AGM campaigns. Thus, the equilibrium exhibits persistent separation: early campaigns are a credible signal of influence, not a pooling outcome.
2.5 Discussion and Empirical Implications
The model links the strategic timing of NGO campaigns to the joint behavior of NGOs and shareholders. The cutoff policy $\bar{\pi}(A)$ determines when an NGO switches from AGM to early campaigning: as credibility $\pi$ and awareness $A$ rise, the marginal value of early influence increases and the threshold $\bar{\pi}(A)$ declines. This dynamic generates a lifecycle pattern in which young NGOs focus on visibility and awareness building, while mature NGOs campaign earlier to exert direct influence.
The model yields several testable implications for both NGOs and shareholders. AGM campaigns, which maximize visibility, expand awareness ($A_{t}$) but do not directly affect contemporaneous success rates since $_{\mathrm{AGM}}(\pi,A)\equiv 0$. Their impact on future shareholder filings depends on both the awareness and the credibility channels as captured by the success function $S(A_{t},\pi_{t},d_{t})$ in (4). Because $S$ increases in both $A_{t}$ and $\pi_{t}$, higher awareness raises the baseline likelihood that shareholders notice and support related proposals, while greater credibility increases the perceived probability that the NGO’s engagement will be effective. The relative strength of these channels is governed by the parameter vector $h$, which is left unspecified. Hence, whether shareholder filings are primarily driven by awareness or by credibility remains an open question within the model, depending on how visibility and reputation interact to shape $S(A_{t},\pi_{t},d_{t})$.
Early campaigns, by contrast, work through the credibility channel. They generate informative outcomes and update beliefs via (7), raising $\pi_{t}$ when $y_{t}=1$ and thereby increasing expected proposal success in subsequent periods. Because the effect of timing enters success probabilities through $_{\mathrm{early}}(\pi,A)>0$, early campaigns also enhance contemporaneous influence: credible NGOs that campaign early obtain higher vote support and a greater likelihood of success within the same AGM cycle but do not affect future filing.
Over time, credibility accumulates endogenously. The average success rate of an NGO’s campaigns therefore rises with its reputation $\pi_{t}$, which reflects the history of its visible and early actions. New NGOs that have campaigned mostly at AGMs build awareness but remain less credible, whereas established NGOs that have undertaken successful early campaigns achieve higher influence on both current and future outcomes. As awareness and credibility jointly accumulate, the cutoff $\bar{\pi}(A)$ declines, prompting NGOs to shift gradually from AGM to early timing. This transition generates a lifecycle pattern in which young NGOs cluster around AGM dates to gain visibility, while mature NGOs increasingly campaign early to exert direct influence. Early successes accelerate this process by raising $\pi_{t}$, whereas early failures slow reputation building, producing persistent heterogeneity in influence and effectiveness across NGOs.
In this simple model, only credible NGOs switch to early campaigns. Our framework abstracts from a potential third region in which both credible and weak NGOs would eventually choose early campaigning. Such a region can arise in richer structures where weak NGOs derive some direct influence payoff or gain utility from being perceived as credible. Whether and when this occurs, however, depends entirely on functional-form assumptions about influence and visibility [^3] In richer specifications, weak NGOs could obtain a small influence payoff (e.g. $^{W}_{\mathrm{early}}(A)>0$) or benefit from being perceived as credible, implying that both types might eventually choose early timing. The location of such a region, however, depends on functional-form details we cannot discipline empirically. We therefore impose the simplifying assumption $^{W}=0$ and focus on the empirically relevant transition in which only credible NGOs switch to early campaigns. By imposing $^{W}=0$, we emphasize the empirically relevant transition from pooled AGM campaigning at low credibility to early campaigning by credible NGOs as reputation accumulates.
Interpretation.
The model follows the spirit of reputation-for-toughness models (kreps1982reputation; milgrom1982limit), where players of uncertain type use costly actions to build a reputation for persistence. Here, early campaigning is the costly signal: it forgoes short-term visibility and warm-glow utility (or donations) for long-term influence and reputational gains. AGM campaigns, by contrast, are cheap – they provide exposure but little information. Introducing the deterministic awareness stock extends this logic by separating visibility-driven exposure (awareness) from credibility-driven reputation. Both dimensions are necessary to reproduce the observed timing shift: with only warm glow and awareness, NGOs would campaign persistently at AGMs; adding reputation explains the empirical transition from AGM visibility to early influence as NGOs mature.
3 Data
Our dataset integrates four sources: NGO campaigns, online attention, shareholder proposals, and firm-level news.
First, we use data from Sigwatch, a consulting firm tracking global NGO campaigns against corporations from 2010 to 2021 (hatte2020geography). The data record the campaign’s public release date, cause (e.g., “GMOs in food”), the targeted company and its prominence, and sentiment (ranging from –2 to +2, typically negative). We focus on single-firm, negative-sentiment campaigns where the firm appears in the headline. [^4] Since our analysis links campaign timing with a firm’s AGM date, we exclude multi-firm campaigns as their AGM dates may differ. Single-firm campaigns represent 93% of the original sample. The final dataset spans 2,512 campaigns launched by 731 NGOs against 523 firms.
Second, we collect monthly Google Trends scores for all NGOs in our sample from 2010 to 2022. Scores are normalized on a 0–100 scale, either (i) relative to the NGO’s own peak searches or (ii) relative to a benchmark NGO in the same month. We use Greenpeace – the most searched NGO – as the benchmark for comparability across organizations.
Third, we use data from Institutional Shareholder Services (ISS) on proposals at the Annual General Meetings (AGMs) of the 3,000 largest US-listed firms. We select 4,483 environmental and social (E&S) proposals filed between 2010 and 2021 (837 firms in total). About one-third of these proposals are successful, either withdrawn before the AGM (after firm commitments) or passed in a vote. Both the number of E&S proposals and their success rate increased steadily over the period.
Fourth, we obtain firm-level news from RavenPack, which transforms unstructured news into structured data with sentiment and event tagging. [^5] RavenPack News Analytics assigns tone scores to news articles drawn from over 40,000 sources, such as Dow Jones Newswires, the Wall Street Journal, MarketWatch, Factiva, and Barron’s, and identifies articles linked to specific events. The validity of this score has been established in prior research (e.g., bushman2024influence). The dataset also differentiates between media-generated coverage and firm-issued press releases, enabling us to focus on the former. We count news items related to US-listed firms from 2010 to 2020, retaining only items with a relevance score of at least 75%. The final sample includes 5,212 firms.
Lastly, we collect balance sheet and income statement data for US NGOs in the sample from their Internal revenue Service form 990, available on the National Center for Charitable Statistics website. Form 990 is the annual information return that most tax-exempt nonprofit organizations in the United States must file with the IRS to disclose their finances, governance practices, and programmatic activities. Filing obligations are primarily determined by organizational size and type: tax-exempt entities with gross receipts over \$200,000 and assets over \$500,000 file the full Form 990. From our starting sample, we are able to link 115 NGOs to IRS filed forms.
We link Sigwatch, ISS, and RavenPack data using standard identifiers (CUSIP, ISIN, year). Table 1 reports descriptive statistics. Panel (a) shows that, on average, NGOs launch 3.2 campaigns during the sample, 24% of which are on the target AGM date, though over half of NGOs appear only once. Most campaigns focus on social rather than environmental issues, but the distribution is highly skewed: a small set of NGOs accounts for a disproportionate share of activity, while the median NGO runs only a single campaign. Panel (b) documents substantial variation across targeted firms in size and news coverage. Targets range from very large firms with extensive media presence to smaller firms with limited coverage, suggesting that NGOs are not exclusively focused on high-visibility corporations. 53% of the campaigns are on social causes, with the remaining being environmental. We return to the remaining panels of Table 1 in Section 5, after first providing background on AGM dates.
Table 1: Summary Statistics
Panel a: Variables varying by NGO
Mean 25% 50% 75% 95%
Number of campaigns by NGO 3.20 1 1 3 13
     – Campaigns on Env. topic 1.51 0 1 1 5
     – Campaigns on Soc. topic 1.70 0 1 1 6
     – On AGM date 0.24 0 0 0 1
Campaign sentiment -1.72 -2 -2 -1 -1
Form 990-filing U.S. NGO 0.15 0 0 0 1
     – Number of campaigns 10.19 1 3 14 33
     – Campaigns on Env. topic 4.36 0 1 4 19
     – Campaigns on Soc. topic 2.83 0 1 4 12
     – On AGM date 0.28 0 0 0 1
Campaign sentiment -1.75 -2 -2 -2 -1
NGO Total Assets (USD mn) 50.02 2.80 7.95 53.91 248.02
Yearly donations received (USD mn) 38.03 3.18 10.13 59.14 143.09
NGO age (years) 42.79 21 32 47 126
Panel b: Variables varying by target firm
Mean 25% 50% 75% SD
Market Cap (USD bn) 175.66 21.71 75.05 209.47 263.66
Total Assets (USD bn) 98.81 21.80 59.27 138.68 124.36
Yearly Sales (USD bn) 37.26 6.85 17.35 42.41 150.58
Yearly Number of News 1,183 328 616 1,046 1,695.25
Environmental score 5.66 4 5.58 7.3 2.22
Social score 4.44 3.30 4.70 5.60 1.58
Campaign on Env. topic 0.47 0 0 1 0.50
Campaign on Soc. topic 0.53 0 1 1 0.50
  • Note: This table reports summary statistics and robustness checks across four panels.
  • Panel (a). Summary statistics on NGO campaigns (Sigwatch); AGM dates (ISS).
  • Panel (b). Target-firm statistics: accounting (Compustat), ESG (MSCI), news counts (RavenPack).
4 Empirical Strategy
Our empirical strategy exploits the timing of NGO campaigns relative to AGMs, leveraging the fact that AGM dates are largely exogenous and stable within firms over time due to SEC regulations. Rule 14a-8 requires shareholder proposals to be filed at least 120 days before the proxy statement, effectively anchoring AGM dates across years (glac2014influence; fioretti2024shared). Figure 2 illustrates this stability by comparing each firm’s AGM month across consecutive years. Most observations lie along the diagonal, indicating strong persistence of the AGM month.
This predictability is central to our identification strategy. Because AGM dates are fixed well in advance and rarely change, NGOs face an exogenous schedule of heightened public attention: they can either react immediately to external events or strategically align their campaigns with these predetermined meetings, when firms are most exposed to shareholder and media scrutiny. In this section, we exploit this feature to study how campaigning around AGMs affects NGOs’ visibility (Section 5), stakeholder reactions (Section 5.1), and firms’ subsequent outcomes (Section 6.1).
Figure 2: Persistence of AGM month over time
Note: The figure plots the percentage of AGMs held in each month (y-axis) given the month of the previous year’s AGM (x-axis). Larger and darker circles indicate a higher proportion of meetings. Data cover 38,471 meetings of 5,745 U.S. companies from 2010–2020.
4.1 Target Selection
A potential concern is that NGOs might select target firms strategically to maximize either visibility or credibility rather than choosing campaign timing exogenously. For instance, NGOs could preferentially target firms that are already more visible, frequently covered in the media, or more responsive to shareholder pressure. Table 2 addresses this concern by testing for selection on observables o
Panels (a) and (b) examine whether NGOs with prior AGM campaigns systematically target different types of firms. Panel (a) compares average firm characteristics across NGOs with and without previous AGM campaigns. In addition, Appendix Figure B1 regresses an indicator for AGM-timed campaigns on firms’ financial and ESG characteristics and the campaign’s cause, controlling for year fixed effects and clustering standard errors at the firm level. [^6] The independent variables include the target firm’s market capitalization (Market Cap), yearly sales (Sales), total assets (Total Assets), number of news over the past 12 months (News count), environmental and social scores (Env. score, Soc. score), campaign sentiment, and an indicator for whether the campaign concerns environmental topics (Env. campaign). None of these variables significantly predicts AGM targeting. A one–standard-deviation increase in any firm characteristic shifts the likelihood of an AGM campaign by less than 5 percent of a standard deviation, indicating little selection on observables regarding on whether target firms of NGOs targeting AGM campaigns systematically differ from other NGOs.
Panel (b) further regresses target characteristics on the number of an NGO’s past campaigns (the results are similar for past AGM campaigns). We include standard balance-sheet variables (market capitalization, total assets, and sales) as well as media and ESG indicators (news coverage, environmental and social scores, and topic dummies). None of these coefficients is economically and statistically significant, except for a mild increase in environmental campaigns over time, consistent with a post–Paris Agreement shift in NGO focus toward environmental issues. Panels (c) and (d) turn to NGO-level selection. Panel (c) focuses on 541 U.S. NGOs with balance-sheet data and regresses firm characteristics on NGO size, proxied by deciles of total assets, while Panel (d) correlates NGO financial ratios – assets, donations-to-assets, and expenses-to-assets – with the number of past campaigns. In both analyses, coefficients are small and statistically insignificant, showing no systematic link between NGO scale or financial capacity and target-firm characteristics.
Taken together, these results indicate that neither NGO nor firm characteristics predict AGM-timed activism. The evidence supports our interpretation that campaign timing reflects strategic decisions by NGOs rather than endogenous target selection. However, while our analysis rules out selection on observables, unobserved factors may still affect campaign timing – a possibility we address in Section 8, where we exploit quasi-exogenous variation in campaign dates.
Table 2: Selection on Observables
Panel a: Avg. target characteristics’ based on whether the campaigning NGO campaigned on AGM dates in the past
NGO w/ Past AGM Campaigns NGO w/o Past AGM Campaigns Test Difference
Mean S.E. Mean S.E. p-value
Market Cap (decile) 9.23 0.10 9.07 0.04 0.13
Total Assets (decile) 9.26 0.08 9.23 0.04 0.72
Sales (decile) 9.31 0.09 9.14 0.04 0.07
Yearly Number of News (lag) 1,132.37 114.74 1,368.67 47.67 0.06
Environmental score 5.64 0.12 5.72 0.06 0.58
Social score 4.44 0.09 4.45 0.04 0.89
Campaign on Env. topic 0.46 0.03 0.5 0.01 0.15
Campaign on Soc. topic 0.54 0.03 0.49 0.01 0.12
Panel b: Target’s characteristics based on NGO’s past campaigns regressing $z_{ny}=\beta_{0}+\beta_{1}\cdot\textit{\# Past Campaign}_{g(n)y}+\varepsilon_{ny}$
$\hat{\beta}_{1}$ (SE) w/o FEs $\hat{\beta}_{1}$ (SE) w/ FEs N
Dependent Variable ($z_{ny}$): (1) (2) (3) (4)
Market Cap Decile -0.002 (0.003) 0.003 (0.016) 1,338
Total Assets Decile 0.003 (0.004) -0.002 (0.012) 1,338
Sales Decile 0.001 (0.004) -0.002 (0.03) 1,337
Yearly Number of News 0.996 (12.4) -4.12 (16.5) 1,323
Environmental score 0.004 (0.009) 0.041 (0.041) 1,338
Social score -0.0006 (0.006) 0.019 (0.032) 1,338
Campaign on Env. topic 0.006∗∗∗ (0.002) 0.021∗∗ (0.010) 1,338
Campaign on Soc. topic -0.006∗∗∗ (0.002) -0.018 (0.012) 1,338
Panel c: Target’s characteristics based on NGO size regressing $z_{ny}=\beta_{0}+\beta_{1}\cdot\textit{Size Decile}_{g(n)y}+\varepsilon_{ny}$
$\hat{\beta}_{1}$ (SE) w/o FEs $\hat{\beta}_{1}$ (SE) w/ FEs N
Dependent Variable ($z_{ny}$): (1) (2) (3) (4)
Market Cap Decile 0.015 (0.016) -0.006 (0.022) 541
Total Assets Decile 0.030 (0.023) 0.010 (0.025) 541
Sales Decile -0.008 (0.019) -0.002 (0.018) 541
Yearly Number of News -0.020 (0.036) 0.022 (0.028) 508
Environmental score 0.061 (0.068) -0.501∗∗ (0.189) 521
Social score 0.016 (0.039) 0.042 (0.120) 521
Campaign on Env. topic 0.026∗∗ (0.012) -0.006 (0.036) 541
Campaign on Soc. topic -0.024∗∗ (0.012) 0.016 (0.037) 541
Panel d: NGO characteristics based on past campaigns regressing $z_{gy}=\beta_{0}+\beta_{1}\cdot\textit{\# Past Campaign}_{gy}+\varepsilon_{gy}$
$\hat{\beta}_{1}$ (SE) w/o FEs $\hat{\beta}_{1}$ (SE) w/ FEs N
Dependent Variable ($z_{ny}$): (1) (2) (3) (4)
Total Assets (log) 0.046∗∗∗ (0.013) 0.003 (0.005) 1,079
Donations to assets -0.023 (0.025) -0.008 (0.016) 1,079
Expenses to assets -0.027 (0.025) -0.006 (0.015) 1,079
  • Note: This table reports summary statistics and robustness checks across four panels.
  • Panel (a). No-selection-on-observables test comparing targets where the NGO had previously campaigned on at least one of its targets’ AGM dates. Market Cap, Total Assets, and Sales are deciles using all U.S. Compustat firms in the campaign year.
[…]
  • Panel (d). Change in NGO characteristics as an NGO campaigns multiple times. Each row reports $\beta_{1}$ from the panel-header regression of a dependent variable (e.g., Total assets (log)) on the NGO’s number of past campaigns up to the previous year. Columns (3)–(4) include NGO and Year fixed effects. Standard errors clustered by NGO.
5 Campaign Timing and Exposure
The model predicts that campaigns timed on AGM dates generate a larger visibility shock than early campaigns – formally, that $v(\mathrm{AGM})>v(\mathrm{early})$ in the awareness law of motion (1). Because neither the latent visibility inputs $v(\cdot)$ nor the awareness stock $A_{t}$ are observed directly, we test this implication using observable proxies that capture different components of real-world awareness formation. Google search activity for the NGO reflects demand-based attention, measuring how many individuals actively seek information about the organization. Negative news coverage of the target firm reflects a supply-side media response, capturing how much information is produced and disseminated about the campaign. While the model remains agnostic about the microfoundations of awareness, it implies that AGM-timed campaigns should generate larger visibility responses in these variables than campaigns launched at other times.
To test this, we estimate the following stacked event-study specification:
$\text{Awareness}_{g,m,y}$ $=\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}^{AGM}\cdot\text{AGM Campaign}_{c}\cdot\mathbf{1}_{\{M_{c,m,y}=\tau\}}\right)$ (10)
$\quad+\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}^{Non-AGM}\cdot\mathbf{1}_{\{M_{c,m,y}=\tau\}}\right)+\alpha_{c,y}+\alpha_{m,y}+u_{g,m,y}.$
Because NGOs campaign multiple times, we follow baker2022much and stack all campaign events. Each event is denoted $c(i,n,g,y,m)$, where campaign $c(\cdot)$ is identified by cause $i$ (e.g., pollution), NGO $g$ (e.g., Greenpeace), target firm $n$ (e.g., Shell), calendar year $y$, and month $m$. The event-time indicator $\mathbf{1}\{M_{c,m,y}=\tau\}$ equals one if month $m$ is exactly $\tau$ months from the campaign, with $\tau=-1$ serving as the reference period. The coefficients $\beta_{\tau}^{AGM}$ capture the incremental visibility response to AGM-timed campaigns relative to the general effect of campaigning captured by $\beta_{\tau}^{Non\text{-}AGM}$. NGO-firm-year fixed effects ($\alpha_{c,y}$) absorb all time-invariant characteristics of each NGO-target pair within a year, while month-year fixed effects ($\alpha_{m,y}$) absorb seasonality and aggregate monthly shocks. Standard errors are clustered at the NGO level.
The model implies that the impact of visibility shocks on awareness depends on an NGO’s position in the $(A_{t},\pi_{t})$ state space. To capture this heterogeneity, we estimate (10) separately for younger and more mature NGOs. NGOs with fewer than four total campaigns are classified as “young,” while those with four or more campaigns are treated as “mature,” reflecting higher accumulated awareness and reputation. [^7] Appendix Figure B2 replicates the analysis on the full sample. Comparing its estimates with those in Figure 3 (discussed below) shows that the increases in Google Trends scores and in negative news coverage are largely driven by NGOs in their first few campaigns (i.e., “young” NGOs).
Panel (a) of Figure 3 reports the estimated $\beta^{AGM}_{\tau}$ using Google Trends as a proxy for demand-based attention. AGM campaigns by young NGOs generate a sizeable rise in search activity, though the estimates are imprecise. For mature NGOs, the AGM-related coefficients are small and statistically indistinguishable from zero, consistent with saturation of awareness: once $A_{t}$ is high, additional visibility shocks produce little incremental response.
[…]
Figure 3: Increased NGO visibility around an AGM campaign
(a) Campaign timing and campaigning NGO’s Google Trends scores
(b) Campaign timing and number of target firms’ negative news (standardized)
Note: Each panel plots the estimated $\beta^{AGM}$ coefficients from (10). In Panel (a), the dependent variable is the monthly Google Trends score for the NGO. In Panel (b), the dependent variable is the (normalized) number of monthly negative media coverage received by a target firm. Blue dots are estimated by restricting the sample to an NGO’s first three campaigns, while red triangles consider later campaigns. Error bars shows 95% confidence intervals using standard errors that are clustered by NGO.
The contrast between early and later campaigns also sheds light on the persistence of awareness. If awareness evolves according to $A_{t+1}=\rho A_{t}+(1-\rho)v(d_{t})$, then the first few campaigns produce the largest increases in $A_{t}$ unless $\rho$ is very close to one. The empirical pattern – strong AGM effects for young NGOs and no detectable response for mature NGOs – accordingly points to an awareness process that converges relatively quickly. [^8] To see this formally, suppose an NGO always chooses $d_{t}=\mathrm{AGM}$ and normalize $v(\mathrm{AGM})=1$. Then awareness follows $A_{t+1}=\rho A_{t}+(1-\rho)$, which solves to $A_{t}=1-\rho^{\,t}$. Awareness therefore converges quickly when $\rho$ is moderate: for example, with $\rho=0.5$, the first three campaigns raise $A_{t}$ to $1-0.5^{3}=0.875$, leaving little room for additional campaigns to generate detectable increases in visibility. This illustrates why strong AGM effects appear for early campaigns but not for later ones. Given that the average NGO conducts 3.2 campaigns, treating NGOs with four or more campaigns as “mature” is thus a reasonable way to proxy for having reached a high-awareness region of the state space.
Note that the estimates are obtained within NGOs: the sample split is not based on the total number of campaigns between 2010 and 2020, but on each NGO’s early versus later campaigns: that is, all NGOs appearing in the red regressions also appear in the blue regressions. Thus, early in their lifecycle, NGOs benefit more from AGM timing to build visibility, but the returns to this tactic decline as their public recognition grows.
Robustness
These results are robust to addressing three main concerns. First, the stacked event-study framework implicitly treats campaign timing as simultaneous across NGOs, whereas in reality campaigns are staggered across organizations and calendar months. This misalignment can distort estimates when treatment effects are heterogeneous across groups or over time. Second, some NGOs launch multiple campaigns within the same year, which may confound the analysis if two campaigns fall within five months of each other. To address both issues, Appendix C replicates the analysis using the staggered difference-in-differences estimator of sun2021estimating, which explicitly accounts for variation in treatment timing, as well as a distributed-lag specification that flexibly controls for past campaigns within a year. Both approaches yield similar results, with substantially larger gains for AGM campaigns by younger NGOs.
Third, although we find no evidence of pre-trends, NGOs might still select targets strategically to raise visibility. Section 4.1 examines this concern in detail. There we show that target-firm characteristics are unrelated to an NGO’s past campaigning behavior and that NGO size and financial capacity similarly fail to predict which firms are targeted. The only systematic pattern is a gradual shift toward environmental campaigns, consistent with broader post–Paris Agreement trends.
While these findings rule out selection on observables, selection on unobservables remains possible. To address this, Section 8 exploits quasi-exogenous variation in the timing of firms’ inclusion in public shaming lists, providing an external validation of our main identification strategy.
Next, we turn to the reaction of stakeholders to NGOs.
5.1 Stakeholders’ Reactions
We next examine how key stakeholders respond to NGO campaigns. In the model, shareholders react by filing proposals when awareness and credibility increase (Section 5.1.1), while donors respond through warm-glow motives that translate into higher contributions (Section 5.1.2). In addition, although consumer behavior lies outside the model, we explore whether campaign timing affects target firms’ revenues as a proxy for potential consumer backlash (Section 5.1.3).
5.1.1 Shareholders
We begin with shareholders. We test whether they are more likely to submit proposals on the same cause when an NGO campaigned around the firm’s AGM date in the previous year, as suggested by the filing rule in (6). A proposal is defined as “related” when it addresses the same cause as the NGO campaign. As detailed in Appendix D, we identify related proposals by matching each campaign topic in the Sigwatch data to the topic of an AGM proposal at the same target firm in the following year, using ISS topic definitions. Whenever such a match occurs, we classify the shareholder proposal as being on the same topic as the earlier NGO campaign.
To examine this, we recalendarize time so that month 0 corresponds to each firm’s AGM month, taking advantage of the stability of AGM dates over time. This creates a common event-time scale on which to compare campaigns launched near versus far from the AGM. We then estimate the following regression:
$\text{Proposal on Same Cause}_{i,n,y}$ $=\sumop\displaylimits^{\tau=-4}_{\begin{subarray}{c}\tau=-15\\ \tau\neq-4\end{subarray}}\left(\beta_{\tau}\cdot\mathbf{1}_{\{\text{Distance to AGM${}_{c(i,n,g,y,m)}$}=-\tau\}}\right)$ (11)
$+\alpha_{i,s(n),g,m,y}+\gamma_{n}+u_{i,n,y}$
For each campaign $c(i,n,g,y,m)$ launched by NGO $g$ against firm $n$ on cause $i$ in year $y$ and month $m$, we analyze the likelihood that the next AGM of firm $n$ includes a shareholder proposal on the same cause. Formally, $\text{Proposal on Same Cause}_{i,n,y}$ equals one if such a proposal is observed within 4 to 15 months after campaign $c(\cdot)$, and zero otherwise, with timing measured by the indicator $\mathbf{1}_{\{\text{Distance to AGM}_{c(i,n,g,y,m)}=-\tau\}}$, which is one if campaign $c(\cdot)$ occurs $\tau$ months from the the target’s AGM date, and zero otherwise. Since SEC 14a8(e)(2) rule prevents shareholders from including proposals in the proxy statement within 120 days of its release, we cap the analysis at four months before the previous year’s AGM and set the reference month to $\tau=-4$.
Although firm-by-time fixed effects are not possible in this setting as they would absorb the variation used to estimate $\beta_{\tau}$, this analysis purges all the variation from cause-by-industry-by-NGO-by-month-by-year through $\alpha_{i,s(n),g,m,y}$ and firm invariant characteristics through $\gamma_{n}$, which we view as the main confounding factor since issue salience varies systematically across industries, causes, and NGOs over time. We cluster standard errors by firm.
Panel (a) of Figure 4 compares an NGO’s first three campaigns (blue dots) with its later campaigns (red triangles). AGM-timed campaigns (identified at $\beta_{-12}$) by more established NGOs increase the probability of a related shareholder proposal by roughly two percentage points, whereas AGM campaigns by younger NGOs generate no significant response. This pattern is consistent with the success function $S(\cdot)$ in (6): for more mature NGOs, beliefs $\pi_{t}$ are higher, so the expected success probability $\pi_{t}S^{C}+(1-\pi_{t})S^{W}$ lies closer to the credible-type benchmark $S^{C}$, making shareholders more responsive to visibility shocks. Although fully credible NGOs would eventually shift entirely to early campaigns in the model, the mature NGOs in our data need not have reached that limiting region. AGM-timed campaigns by higher $\pi$ NGOs are therefore consistent with NGOs still transitioning toward higher credibility and awareness, for whom AGM visibility remains sufficiently informative to affect filing decisions. In contrast, the youngest NGOs are not credible enough to push shareholder to submit proposals. [^9] Appendix Figure 3(a) replicates the analysis on the full sample, suggesting that more “mature” NGOs drive the effect we observe in the full sample.
Panel (b) focuses on the subset of U.S. NGOs for which we observe balance-sheet data and incorporation year, allowing us to compare younger and more mature organizations based on age at incorporation. We classify as “young” (blue dots) those founded after 2000, corresponding to NGOs younger than ten years at the start of the sample. The pattern is similar to Panel (a): a clear spike in filing rates appears at t=12t=-12 for mature NGOs (red triangles), whereas effects for younger NGOs are smaller and imprecisely estimated. Note that these NGOs are generally larger than those in Panel (a), which explains why we see some actions around the AGM also for younger NGOs.
Figure 4: NGO campaigns and shareholder proposals
(a) By past experience
(b) By NGO age
Notes: Note: These figures plot the estimated coefficients of the regression of a dummy indicating the occurrence of a shareholder proposal on month $t$ at firm $f$, relative to an NGO campaign targeting firm $f$ on the same cause, $k$ months ahead. The regression also includes campaign and month-by-year-by-industry-by-topic fixed effects. Panel (a) regression is run on the full sample (1,227 NGO campaigns), while the one in Panel (b) is based on U.S. NGOs with balance sheet data (541 campaigns). Error bars shows 95% confidence intervals using standard errors that are clustered by firm.
5.1.2 Donors
Focusing on the smaller sample of U.S. NGOs for which balance-sheet data are available, we examine whether campaigns timed around firms’ AGMs translate into higher donations for the campaigning NGO, and whether this response differs across NGO maturity.
Because donations are observed only annually, we cannot align them precisely with the timing of individual campaigns, and many NGOs conduct multiple campaigns within the same year. Stacking event windows would therefore generate overlapping observations without adding variation on the left-hand side. Instead, we estimate a distributed-lag specification that relates the timing of campaigns relative to AGMs to annual financial outcomes:
$\frac{\text{Donations}_{g,y}}{\text{Assets}_{g,y-1}}=\sumop\displaylimits_{\tau=-4}^{4}\left(\beta_{\tau}\cdot\mathbf{1}_{\{\text{Distance to AGM}_{c(i,n,g,y,m)}=\tau\}}\right)+\alpha_{g}+\alpha_{y}+u_{g,y},$ (12)
where “Distance to AGM” measures how NGO $g$’s campaigns in year $y$ are distributed relative to the target firms’ AGMs. NGO fixed effects ($\alpha_{g}$) absorb persistent differences in size or donor base, while year fixed effects ($\alpha_{y}$) capture aggregate shocks. [^10] We treat total reported revenues as donations.
Panel (a) of Figure 5 shows that young NGOs (blue dots) experience a clear increase in donations when campaigning on the AGM date, while older NGOs (red triangles) show no meaningful response. Panel (b) presents similar patterns for return on assets, with sharp AGM effects only for young NGOs. Because ROA nets out expenditures, these findings suggest that AGM-day campaigns are relatively inexpensive yet effective in generating additional funding for less established NGOs.
These donation patterns should not be interpreted as estimates of the model’s warm-glow term $\alpha v(d_{t})$. Instead, they provide reduced-form evidence on how the exposure created by campaign timing translates into financial support. As shown in Figure 3, AGM campaigns generate substantially larger visibility shocks for young NGOs, making it natural that only these organizations exhibit a measurable donation response, while the same exposure increment has little effect for mature NGOs. [^11] Appendix Figure B4 finds a similar AGM-day effect on the full sample, highlighting the important need for visibility of “younger” NGOs.
Figure 5: NGO campaign timing and donations
(a) Donations
(b) Operating Margin
Notes: The figure plots the estimated ${\beta}_{\tau}$ from (12). Panel (a) focuses on donations over (lagged) assets, while Panel (b) focuses on donations minus expenditures over (lagged) assets. Error bars shows 95% confidence intervals using standard errors that are clustered by NGO.
5.1.3 Consumers
We next examine whether NGO campaigns trigger a consumer response, even though such behavior lies outside the model. Using quarterly Compustat data, we study how the target firm’s log of sales varies with the timing of NGO campaigns relative to the firm’s AGM in a specification similar to (10):
$\log\left(\text{Sales}_{n,q,y}\right)$ $=\sumop\displaylimits^{\tau=4}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}^{AGM}\cdot\text{AGM Campaign}_{c}\cdot\mathbf{1}_{\{M_{c,q,y}=\tau\}}\right)+\sumop\displaylimits^{\tau=4}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}^{Non-AGM}\cdot\mathbf{1}_{\{M_{c,q,y}=\tau\}}\right)$ (13)
$+\log\left(\text{Assets}_{n,q,y}\right)+\alpha_{c,y}+\alpha_{m,y}+u_{n,q,y}.$
where $q$ denotes the quarter, $\alpha_{c,y}$ are campaign fixed effects, $\alpha_{q}$ are industry-by-quarter fixed effects, and standard errors are clustered at the campaign level.
Figure 6 reports the estimated $\hat{\beta}_{\tau}$} coefficients, separately for campaigns launched on the AGM date (Panel a) and for all other campaigns (Panel b). Across both specifications, we find no meaningful change in sales around NGO campaigns, suggesting that consumers do not systematically adjust their purchasing behavior in response to NGO activism.
Taken together with earlier results, this evidence indicates that NGO campaigns primarily affect stakeholders other than consumers, especially shareholders and donors, consistent with the model’s focus on visibility and influence within the investor domain.
Figure 6: NGO campaign timing and sales
(a) NGO Campaigns on AGM dates
Notes: The figure plots the estimated ${\beta}_{\tau}$ from (13). The reference period is -1 quarter, consistent with the other analyses. The number of observations is 239,818 in Panel (a) and 239,925 in Panel (b). Error bars shows 95% confidence intervals using standard errors that are clustered by firm.
Overall, the evidence in this section indicates that AGM timing delivers meaningful visibility gains mainly for younger NGOs. We now turn to whether this timing advantage also translates into differences in shareholder proposal outcomes and firm responses.
6 Campaign Timing and Influence
The model predicts that early campaigns offer NGOs greater scope to influence contemporaneous AGM outcomes through the term $_{d_{t}}$ in the proposal success equation (4). Because most shareholders cast their votes by proxy before the meeting, campaigns launched on the AGM date have limited ability to affect voting outcomes (i.e., $_{AGM_{t}}=0$). For example, more than 85% of vote-eligible shares for Microsoft’s 2023 AGM were submitted in advance (microsoft2023), illustrating why visibility at the AGM does not necessarily translate into influence over the vote itself.
We classify a shareholder proposal as successful if it is withdrawn before the AGM or receives at least 50% + 1 of the votes at the meeting. To assess whether NGO campaigns influence the success of related proposals, we estimate the following specification on the full sample of shareholder proposals:
$\text{Success}_{i,n,m,y}=\sumop\displaylimits_{\tau=-7}^{0}\left(\beta_{\tau}\cdot\text{Campaign}_{i,n,m+\tau,y}\right)+\alpha_{n,y}+\alpha_{i,y}+\alpha_{m}+u_{i,n,m,y},$ (14)
where $\text{Success}_{i,n,m,y}$ is an indicator for the success of proposal $i$ at firm $n$, and $\text{Campaign}_{i,n,m+\tau,y}$ indicates whether an NGO campaign on the same cause occurred $\tau$ months before the AGM. Time is recentered so that the AGM month is coded as month 0 for each firm, and the reference period corresponds to months 11 to 7 months before the AGM, which we code as $\tau=-7$ in (14). Campaigns in this [11,7][-11,-7] window – roughly three to four quarters before the focal meeting – are more likely to reflect reactions to the previous year’s AGM rather than anticipatory campaigning. The regression includes firm-by-year, cause-by-year, and month fixed effects, thereby exploiting within–firm-year and within-cause variation in campaign timing. Standard errors are clustered at the firm level. [^12] The sample is not split by NGO maturity because matching campaigns and proposals by topic already reduces the number of observations.
Figure 7: NGOs’ campaigns before the AGM help E&S shareholder proposals succeed
(a) Proposal accepted or withdrawn
(b) Vote share in favor of proposal
Note: The figure plots the estimated coefficients from (14). The dependent variables are an indicator variable that is one if a shareholder proposal is withdrawn or successful if voted (911 proposals) in Panel (a), and the percentage of votes in favor of a proposal (2,078 proposals) in Panel (b). Standard errors are clustered at the firm level.
Panel (a) Figure 7 plots the estimated ${\beta}_{\tau}$. We find that NGO campaigns happening 6 months and 1 month before the AGM positively influence the outcome of related shareholder proposals. They increase, respectively, the probability of success by 2 percentage points and 3 percentage points more than a campaign launched more than six months earlier, corresponding to a relative increase between 6 and 8%.
To understand why campaigns launched six months before the AGM are also somewhat successful, we re-estimate (14) on the subset of proposals that went to a vote, using the percentage of votes in favor as the dependent variable. [^13] Since this analysis focuses on voted proposals, we also control for the voting recommendations of management and ISS (“For,” “Against,” or “Abstain”). Panel (a) of Figure 7 shows that only NGO campaigns launched one or three months before the AGM increase support for related shareholder proposals, with an increase of nearly 2 percentage points. Given the generally low baseline support for these proposals (24%), this corresponds to an 8% increase in shareholder support when an NGO campaign addresses the same cause, on average.
In contrast to Panel (a), Panel (b) detects no significant increase in support for campaigns initiated as early as six (or five) months before the AGM. This suggests that the difference between the two figures arises from shareholder proposals being withdrawn after reaching an agreement with the firm. Taken together, these results indicate that as a result of these early campaigns, the target firm may take actions to resolve the issue highlighted in both the campaign and the shareholder proposal.
Note that the subset of shareholder proposals going to a vote is not random; an argument raised, for instance, by flammer2015does. Given the generally low support of shareholder proposals going to vote, managers may let proposals that are the least likely to succeed go to a vote. However, those votes can still be helpful in disciplining firms as large shareholder support to a cause may be hard to conceal for managers (levit2011nonbinding). This reinforces the importance of analyzing all E&S proposals, including those withdrawn as in Panel (a).
6.1 Campaign Timing and Prosocial Changes
We next examine whether the timing of NGO campaigns relates to subsequent changes in firms’ environmental and social (E&S) performance. The model does not explicitly incorporate firm responses driven by public shaming, yet such a mechanism would imply that visibility generated during the AGM – when exposure is highest – could directly pressure firms, corresponding to $_{\mathrm{AGM}_{t}}>0$. By contrast, the model’s influence mechanism operates through stakeholder interactions before the AGM (see equation 3): early campaigns may foster constructive dialogue between shareholders and firms, shaping proposals or prompting managerial adjustments ahead of the meeting as illustrated in the previous section (Figure 7).
To investigate which mechanism is more consistent with the data, we estimate the following specification:
[…]
Figure 8: Campaign Timing and Changes in E&S Scores
Note: The figure plots the estimated coefficients $\beta_{l}$ from equation (15). The dependent variable is the yearly change in a firm’s average E&S scores. Shaded areas show 95% confidence intervals based on firm-level clustered standard errors. Data combine MSCI ESG scores, NGO campaigns from Sigwatch, firm characteristics from Compustat, and AGM dates from ISS.
Figure 8 reports the estimated $\beta_{l}$ coefficients from equation (15), distinguishing an NGO’s first three campaigns (blue dots) from its later campaigns (red triangles). Only early campaigns by more established NGOs – those launched between six and three months before the AGM – lead to measurable improvements in firms’ E&S performance. Following such campaigns, E&S scores rise by about 2% over the subsequent year, whereas campaigns at other times, including during the AGM month, have no statistically or economically significant effect. Combined with the results in Figure 7, this pattern suggests that established NGOs can influence firm behavior ahead of the AGM, potentially by shaping shareholder proposals or managerial decisions without requiring a formal vote.
For younger NGOs, positive effects arise only for campaigns launched close to the AGM, with estimates significant at the 90 percent level. Campaigns occurring earlier in the year show no beneficial impact and may even reduce subsequent E&S scores. This mirrors the spike in shareholder support observed one month before the AGM in Figure 7: NGOs still building credibility may be perceived as effective only when campaigning in the immediate run-up to the meeting, but lack the standing needed to engage constructively with firms at earlier stages. [^15] Appendix Figure B5 replicates the analysis on the full sample, showing that only early campaigns (about 6 months in advance) have a positive impact on ESG scores. The coefficients for the other months are close to zero, consistent with them being the average between the red triangles and blue dots in Figure 8, thereby masking important heterogeneity based on NGO maturity.
7 Exposure vs. Influence
[…]
7.1 Choosing when to Campaign based on Past Campaigns
We first examine whether NGOs shift from targeting AGMs to launching earlier campaigns as they campaign more. Specifically, we estimate
$\text{Distance to AGM}_{c(i,n,g,y,m)}$ $=\sumop\displaylimits_{\tau>1}\left(\beta_{\tau}\cdot\mathbf{1}_{\left\{\text{Campaign \#}_{g,y-1}=\tau\right\}}\right)$ (16)
$\quad+\alpha_{i,y}+\alpha_{g,y}+\alpha_{n,y}+u_{c(i,n,g,y,m)}.$
where the dependent variable measures the number of months between campaign $c(\cdot)$ and the targeted firm’s AGM, and the key explanatory variable is the sequence number of campaigns by NGO $g$ by year $y$. $\alpha_{i,y}$, $\alpha_{g,y}$, and $\alpha_{n,y}$ are topic-by-year, NGO-by-year, and firm-by-year fixed effects, respectively. Identification thus comes from NGOs launching multiple campaigns in the same year. [^16] Appendix Figure B6 shows similar results when controlling only for firm and NGO fixed effects, so that the estimation of $\hat{\beta}_{\tau}$ come from variation across both campaigns and years.
Panel (a) of Figure 9 shows that first campaigns tend to occur right at the AGM, while subsequent ones are launched progressively earlier. This pattern is consistent with NGOs initially using AGMs to gain visibility and later exploiting their visibility to influence voting outcomes.
These patterns are not driven by alternative explanations. First, they are unlikely to reflect NGOs learning that campaigning on the AGM date does not affect voting, since most shareholders have already voted by proxy. Indeed, the same downward trend in campaign timing holds when focusing on NGOs whose first campaign occurred after the AGM (Appendix Figure B7). Second, they are not driven by selection of more “influenceable” targets: neither target size (market capitalization) nor its ESG scores covary with NGO past AGM exposure (Appendix Figures B8 and B9).
7.2 Choosing When to Campaign Based on NGO Visibility
We next use Google Trends to proxy for an NGO’s visibility. For each NGO $g$, we compute a Relative Google Score equal to its average search score in year y1y-1 divided by the corresponding score for Greenpeace – the most searched NGO in our sample – and discretize this measure into quartiles. We then estimate:
$\text{Distance to AGM}_{c(i,n,g,y,m)}$ $=\sumop\displaylimits_{\tau=1}^{4}\left(\beta_{\tau}\cdot\mathbf{1}_{\left\{\text{Relative Google Score Quartile}_{g,y-1}=\tau\right\}}\right)$ (17)
$\quad+\alpha_{i,y}+\alpha_{g}+\alpha_{n,y}+u_{c(i,n,g,y,m)}$
which relates campaign timing to lagged NGO visibility. Because the key regressor varies only across NGOs and years, we cannot include NGO-by-year fixed effects as in earlier specifications. [^17] The Relative Google Score is constructed from the previous year’s Google Trends data, generating variation only across NGOs and years but not within NGO–year cells. Panel (b) of Figure 9 shows that only the most visible NGOs – those in the top quartiles of the relative search distribution – campaign several months before AGMs, whereas less visible NGOs continue to concentrate their actions on the AGM date.
Taken together, these results indicate a systematic shift: NGOs start by targeting AGMs to maximize exposure, and as their reputation grows, they move campaigns forward in time to increase their chances of influencing firm decisions.
Figure 9: NGO reputation and campaign timing
(a) NGOs tend to campaign before the target’s AGM date as they gain experience
(b) More visible NGOs campaign before the target’s AGM date
Note: This figure plots the estimated coefficients from (16). The dependent variable is the months between an NGO campaign and the targeted firm’s AGM date. Standard errors are clustered at the NGO level.
7.3 NGO Reputation and AGM Influence
We now examine how the evolution of NGO reputation shapes their ability to influence shareholder proposal outcomes, tying together the patterns documented in Sections 7.1 and 7.2. In the model, proposal success depends jointly on awareness and credibility through $S(A_{t},\pi_{t},d_{t})$ in (4). NGOs that rely heavily on AGM-timed campaigns accumulate visibility but gain little credibility, so repeated AGM activity eventually leads to lower beliefs $\pi_{t}$ and, with it, reduced persuasive power. By contrast, NGOs that shift toward earlier, influence-oriented campaigns build credibility over time. The model therefore implies that overusing visibility-driven tactics should diminish the effectiveness of AGM campaigns for otherwise credible NGOs, whereas NGOs with stronger reputations should be more effective at shaping proposal outcomes.
To test this implication, we estimate the following specification:
$\text{Success}_{i,n,y}$ $=\beta_{1}\,\text{Campaign on Same Cause}_{c(i,n,g,y,m)}+\beta_{2}\,\text{\# Past Campaigns}_{g,y-1}$ (18)
$\quad+\beta_{3}\,\text{Campaign on Same Cause}_{c(i,n,g,y,m)}\cdot\text{\# Past Campaigns}_{g,y-1}$
$\quad+\alpha_{n,y}+\alpha_{g,y}+\alpha_{i,y}+\alpha_{m(c(i,n,g))}+u_{i,n,y},$
where $\text{Success}_{i,n,y}$” indicates whether a proposal is withdrawn or passes a vote, as in equation (14). The key variable $\text{Campaign on Same Cause}_{i,n,y}$” equals one if NGO $g$ campaigned against firm $n$ on cause $i$ in the ten months preceding the AGM. “# $\text{\# Past Campaigns}_{g,y-1}$” counts NGO $g$’s past campaigns. We absorb firm-by-year, NGO-by-year, cause-by-year, and month variation through fixed effects, and cluster standard errors at the firm level.
Column (1) of Table 3 shows that campaigning on the same cause as a shareholder proposal has little effect on its own ($\beta_{1}\approx 0$). What matters is the NGO’s past activity on that issue: proposals are more likely to succeed when the NGO has previously campaigned, and a one–standard-deviation increase in past campaigns raises proposal success by about 8%. The squared term indicates diminishing marginal effects.
Table 3: NGO’s strategies and support for related shareholder proposals
Dependent Variable: Proposal Passes (0/1)
Model: (1) (2) (3) (4)
Campaign on Same Cause -0.0003 -0.013 -0.005 -0.001
(0.026) (0.028) (0.025) (0.030)
# Past Campaigns -0.034
Campaign on Same Cause $\times$ # Past Campaigns 0.078∗∗∗
# Past Campaigns squared 0.008∗∗
Campaign on Same Cause $\times$ # Past Campaigns squared -0.015∗∗
# Past Non-AGM Campaigns 0.0008
Campaign on Same Cause $\times$ # Past Non-AGM Campaigns -0.011
# Past AGM Campaigns 0.005 -0.034
(0.020) (0.023)
Campaign on Same Cause $\times$ # Past AGM Campaigns 0.048 0.091∗∗∗
(0.026) (0.032)
# Past AGM Campaigns squared 0.010∗∗
Campaign on Same Cause $\times$ # Past AGM Campaigns squared -0.013∗∗
Campaign on Same Cause $\times$ Google Search Popularity 0.071
Campaign on Same Cause $\times$ Google Search Popularity square -0.021
Fixed-effects
CUSIP $\times$ Year
NGO $\times$ Year
Cause $\times$ Month-Year
Observations 36,543 35,778 36,543 33,142
R2 0.61894 0.61698 0.61995 0.62078
*–p<0.1<0.1; **–p<0.05<0.05; ***–p<0.01<0.01.
  • Note: This table reports the estimated coefficients from (18) in Column 1. In Column 2, we also consider the number of campaigns the NGO has launched in the past which were not on the firm’s AGM day (# Past Non-AGM Campaigns) and the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (# Past AGM Campaigns), and their interaction with Campaign on Same Cause. In Column 3, we focus on the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (# Past AGM Campaigns), its square, and their interaction with Campaign on Same Cause. In Column 4, we replace # Past Campaigns from model 18 by NGO $g$’s average Google Trends Score relative to GreenPeace the year before the campaign. All regressions also include firm-by-year, NGO-by-year, Cause-by-month-by-year fixed effects. Clustered (firm) standard errors are reported in parentheses.
Column (2) separates past AGM and non-AGM campaigns. Only AGM campaigns contribute meaningfully: a one–standard-deviation increase in past AGM campaigns raises proposal success by roughly 4.8 percentage points, whereas non-AGM campaigns have no detectable impact. Column (3) confirms concavity for AGM campaigns. This pattern is consistent with the model’s belief dynamics: AGM campaigns generate visibility but convey limited new information, so repeated AGM activity initially raises beliefs $\pi_{t}$ but eventually adds little, reducing its influence on shareholder decisions.
Column (4) uses an alternative proxy for reputation – NGO visibility from Google Trends, measured relative to Greenpeace in year y1y\!-\!1. The interaction between this visibility measure and campaigning on the same cause is positive but concave, indicating that the effectiveness of AGM campaigns increases with NGO visibility but at a decreasing rate. This aligns with the idea that once awareness and credibility are high, additional exposure contributes little to reputation formation.
Different measures of reputation
We next examine whether broader measures of NGO scale show similar patterns using data on U.S. NGOs. Table 4 replaces Google visibility with lagged assets (Column 1), lagged donations (Column 2), and lagged expenditures (Column 3). We view these variables as manifestations of an NGO’s reputation: financially larger organizations are more established, have accumulated greater public visibility, and possess the operational capacity to sustain campaigns, all of which are consistent with higher positions in the $(A_{t},\pi_{t})$ state space. [^18] A large literature documents that NGO reputation is closely tied to organizational scale, resources, and operational capacity. Larger and better-resourced NGOs are viewed as more credible, legitimate, and effective by external stakeholders (e.g., baur2011moral; cooley2002ngo; mitchell2017reputations). Across all specifications, proposal success initially rises with NGO size but exhibits diminishing returns. This suggests that while more established NGOs are better positioned to influence proposal outcomes, additional scale eventually contributes little new reputational information and may attenuate the marginal effect of further AGM-driven exposure.
Table 4: NGO’s strategies and support for related shareholder proposals
Dependent Variable: Proposal Passes (0/1)
Model: (1) (2) (3)
Campaign on Same Cause -0.299∗∗ -0.167 -0.208
(0.145) (0.119) (0.143)
Campaign on Same Cause $\times$ NGO Size 0.132∗∗∗ 0.073∗∗ 0.089∗∗
(0.049) (0.032) (0.040)
Campaign on Same Cause $\times$ Squared NGO Size -0.011∗∗∗ -0.006∗∗ -0.007∗∗
(0.004) (0.003) (0.003)
Size is defined by Total Assets Donations Expenses
Fixed-effects
CUSIP $\times$ Year
[…]
Observations 12,127 12,367 12,127
R2 0.59598 0.59441 0.59474
*–p<0.1<0.1; **–p<0.05<0.05; ***–p<0.01<0.01.
  • Note: This table reports the estimated coefficients from (18) where, instead of Past Campaign, we use NGO size as measured in terms of its Total Assets (Column 1), Donations (Column 2), and Expenses (Column 3). The dataset includes only campaigns run by U.S. NGOs for which we observe their balance sheet data. These variables are lagged. All regressions also include firm-by-year, NGO-by-year, Cause-by-month-by-year fixed effects. Clustered (firm) standard errors are reported in parentheses.
7.4 Summary of Results
These findings provide clear evidence for the model’s dynamic trade-off between exposure and influence. NGOs initially rely on AGM-timed campaigns to build visibility, but as they accumulate reputation and credibility, they shift toward earlier, influence-oriented actions. Consistent with the model, repeated AGM campaigning eventually generates diminishing reputational returns: past exposure raises the effectiveness of current campaigns only up to a point, after which additional AGM visibility adds little new information and may even weaken influence. Across all empirical measures of reputation – past AGM activity, Google search visibility, and organizational size – we observe the same concave pattern, consistent with credibility saturation and declining informational value of purely visible actions.
8 Accounting for Selection
A remaining concern is that NGOs may choose campaign dates strategically, potentially confounding the timing effects documented above. To address this, we turn to a setting in which both campaign timing and target identities are anchored to an external event and thus provides quasi-exogenous variation.
Institutional background
[…]
Data
We collect daily stock prices (form LSEG), AGM dates, and data on environmental and social (E&S) shareholder proposals (from ISS) for 97 listed firms included in the Index between 2016 and 2019.191919The first report included 40 brands globally, and its coverage has increased to 200 by 2020, located in 16 different countries. We focus on the firms that have been listed between 2016 and 2019.
Stock market reactions
For each firm, we download daily returns on its national stock index (e.g., FTSE 100 for U.K. companies, CAC40 for French companies) and compute its daily excess return over the index. We then cumulate those returns and plot the results in Figure 10, depending on the report’s publication timing relative to the company’s AGM (Panel a), and depending on the company’s initial FTI score (below or above the median that year, in Panel b). The patterns are consistent with the model’s visibility channel: when firms are most exposed to scrutiny, credible activism amplifies the financial costs from a deteriorating reputation (kolbel2017media).
Figure 10: Stock returns around inclusion
(a) Stock returns depending on timing
(b) Stock returns depending on first score
Note: This figure plots the Cumulative Excess Returns of stocks around the publication of the first Fashion Transparency Index covering the company. The excess returns are computed as the difference between the stock’s daily return and the return of its national stock index. The sample covers 97 listed stocks between 2016 and 2019.
Shareholder outcomes
Figure 11 shows that proposals on the same issues as the Index are 7 percentage points more likely to succeed when the Index is published before the AGM and receive about 8 percentage points more support if voted. These magnitudes closely match those estimated in Section 7.3, reinforcing that early, credibility-driven actions enhance influence through coordinated shareholder responses.
Figure 11: Index publication and support for E&S shareholder proposals
Note: This figure reports the average success of shareholder proposals at firms targeted by the Fasion Transparency Index, depending on whether the report is published before (blue bars) or after (gray bars) the firm’s AGM. The sample covers 40 E&S shareholder proposals, filed at 23 firms between 2016 and 2019.
Shareholder proposals
Finally, Figure 12 replicates the timing tests from earlier sections. Publication in the month leading up to the AGM raises the probability of a related proposal at the next AGM, whereas publication two months before the AGM – analogous to early campaigning – leads to roughly 10-point improvements in firms’ FTI scores over the following year.
Figure 12: Publication timing and firm outcomes
(a) Publication timing and related shareholder proposals in the next AGM
(b) Publication timing and FTI score improvement
Note: This Figure plots the estimated coefficients from:
$y_{n,y}=\sumop\displaylimits_{\tau=-5}^{2}\left(\beta_{\tau}\cdot\mathbf{1}_{\{\text{Distance to AGM = }\tau\}}\cdot\text{Treated}_{n,y}\right)+\mathbf{X}_{n,y}\gamma+\alpha_{n}+\alpha_{y}+u_{n,y},$
where $y_{n,y}$ is either an indicator for having a shareholder proposal at the AGM in year $y+1$ (panel (a)) or the year-on-year Fashion Transparency Index score improvement (panel (b)), $\mathbf{1}_{\{\text{Distance to AGM = }\tau\}}$ is an indicator for the report being published $\tau$ months away from firm $n$’s closest AGM, and $\text{Treated}_{n,y}$ is an indicator for firm $n$ being included in the index in year $y$. $\mathbf{X}_{n,y}$ includes the direct effect and controls for firm size. Standard errors are clustered at the firm level. Regression in panel (a) has 284 observations and regression in panel (b) has 162 observations.
These results mirror the model’s main mechanism while accounting for target selection: AGM-timed actions mobilize shareholders, while earlier actions directly influence corporate behavior.
9 Discussion
Our findings show that when NGOs campaign can be as important as whether they campaign. Early in their life cycle, NGOs concentrate actions on AGM dates, when firms are most visible to investors and the media, echoing work on private politics and focal corporate events (baron2001private; egorov2017private; fioretti2024shared). These AGM‑timed campaigns raise media coverage, Google searches, and donations, and increase the likelihood that related shareholder proposals are filed at the next AGM, but they leave current votes and firms’ environmental and social (E&S) performance largely unchanged, aligning with a visibility‑oriented view of NGO activity (yaziji2009ngos).
As NGOs gain reputation, their timing shifts. Campaigns in the months before the AGM raise the success and vote share of related shareholder proposals and are followed by improvements in firms’ E&S scores, consistent with influence operating through informed shareholder coordination and managerial concessions rather than immediate consumer reactions. The transition from AGM‑day exposure to pre‑AGM engagement therefore traces a lifecycle in which NGOs first invest in awareness and then leverage credibility to change corporate behavior (e.g., cabral2010dynamics; baur2011moral).
Conceptually, the paper opens the black box of NGO activism by exploiting one of the few actions that must be public to be effective: campaigns. NGOs are otherwise highly opaque organizations, with limited standardized reporting and scarce data on internal decisions (maxwell2012economic). We show that the timing of these campaigns relative to the target’s corporate calendar reveals both (i) the NGO’s motive – whether it is primarily seeking visibility or trying to exert influence – and (ii) the main audience of the campaign – donors and the broader public versus shareholders and managers. Different audiences imply different mechanisms and map into the heterogeneous firm‑ and shareholder‑level responses we document across sections.
To interpret these patterns, our model compares two mechanisms: an awareness stock $A_{t}$, built by highly visible actions such as AGM‑day campaigns, and a credibility state $\pi_{t}$, shaped by earlier, costlier engagement that moves beliefs and proposal success. Together, these channels rationalize NGOs’ transition from exposure‑oriented tactics to influence‑oriented activism and connect our evidence to work on shareholder voice, exit, and ESG engagement (e.g., cunat2016say; broccardo2020exit; hart2022private).
Beyond these implications, our results also speak to future research on firm strategy and stakeholder salience. We still know relatively little about how firms decide which stakeholders matter in different contexts, despite a large theoretical literature on stakeholder prioritization and corporate objectives that is often qualitative or case-study based (e.g., freeman1984strategic; mitchell1997toward). By providing a novel, quasi-exogenous shift in NGO campaign timing – driven by NGOs’ own incentives to campaign early or late – we offer a quantitative framework to study how firms form and update their objective functions and whose welfare is put on the agenda, complementing recent empirical work on firms’ objectives and stakeholder weights (e.g., scott2002love; romer2006firms; duarte2025competitive).
A further implication of our findings is that firms’ objectives – here, the objectives of NGOs as strategic actors – are not monolithic. Classic approaches often tie organizational objectives to ownership structure (e.g., grossman1977value; grossman1979theory), implicitly assuming that preferences flow from residual claimants downward. Our results instead show that objectives evolve with the views and expectations of other stakeholders.
In the case of NGOs, it is not only donors and shareholders who influence their actions, but also the credibility concerns that arise from how these stakeholders interpret the timing of campaigns. NGOs shift away from AGM-day campaigning not simply to influence firms more effectively today, but because doing so is instrumental to building credibility, which expands their influence tomorrow. In this sense, organizational objectives are shaped endogenously by the incentives created by stakeholder attention and beliefs, rather than solely by ownership or mission. This perspective aligns with recent work emphasizing that firms’ (and nonprofits’) objectives emerge from their strategic interaction with multiple stakeholders rather than from a single principal (e.g., fioretti2022caring).
These results have policy implications for boards, investors, and regulators in the corporate finance tradition. For boards, ignoring early, credibility‑backed campaigns is costly: these are precisely the actions that raise proposal success and are associated with subsequent E&S improvements, suggesting that constructive engagement with reputable NGOs can substitute for more confrontational AGM events (dimson2015active; kolbel2017media).
For investors and stewardship codes, credible NGOs supply information that helps coordinate dispersed shareholders on E&S issues, complementing traditional voice and exit mechanisms. For regulators, rules on AGM timing, proposal access, and disclosure shape the returns to NGO campaigns and thereby how far private activism can help firms internalize social and environmental externalities alongside, or in place of, formal regulation.
A natural question left open by our theory and evidence is how NGOs compete with one another. Prior work treats competition largely as a battle for funding—typically among large, global actors—and measures success by donations (e.g., aldashev2009ngos). Our framework suggests a broader view: NGOs also compete on visibility and credibility, and the form of competition depends on where organizations sit in that visibility–credibility space. When visibility is scarce, NGOs crowd focal moments (AGMs) to attract attention; when reputation matters, competition shifts toward earlier, costlier engagement that builds influence. This raises testable questions about the interaction between attention markets and reputation dynamics: How does competing for visibility affect reputation formation? Do multiple NGOs targeting the same firm amplify or dilute stakeholder coordination? Evidence from related contexts (e.g., microfinance NGOs) indicates competition can change outcomes (ly2012competition); studying its interactions with awareness and credibility is a promising avenue for future work.
10 Conclusion
This paper opens the black box of NGO behavior by showing how campaign timing reflects NGOs’ underlying objectives. Empirically, we use data on roughly 2,500 campaigns to show that AGM-day actions generate large visibility shocks, raising media coverage, search activity, donations, and the likelihood of related proposals at the next AGM, while having little effect on contemporaneous votes or firms’ E&S performance. By contrast, campaigns launched in the months before the AGM significantly increase proposal success, raise vote shares, and are followed by improvements in E&S outcomes.
A simple model in which NGOs trade off two forces explains these patterns. The model features an awareness channel, whereby AGM visibility builds issue salience, and a credibility channel, whereby earlier, costlier actions update beliefs and directly influence shareholder behavior.
Together, the model and evidence reveal NGOs as strategic actors whose objectives evolve with external incentives, vary across contexts, and reflect the views of multiple stakeholders who shape NGOs’ reputational payoffs. These insights matter for policy, showing that NGO campaigns mobilize different classes of investors with distinct implications for firms’ outcomes, whether by exerting direct influence through support for shareholder votes or indirectly, by prompting new proposals at future AGMs. Overall, the results demonstrate how civil-society pressure can help firms internalize social and environmental externalities.
Appendix A Proofs
A.1 Outcome Likelihood Ratio with Bernoulli Success
Let $\theta\in\{C,W\}$ denote the NGO’s type. Conditional on state $(A_{t},\pi_{t})$ and timing $d_{t}$, the proposal outcome $y_{t}\in\{0,1\}$ is Bernoulli with type–specific success probabilities
$S^{\theta}(A_{t},\pi_{t},d_{t})\;\equiv\;\Pr(y_{t}=1\mid A_{t},\pi_{t},d_{t},\theta).$
Hence the conditional likelihood under type $\theta$ is
$p^{\theta}(y_{t}\mid d_{t},A_{t},\pi_{t})=\big[S^{\theta}(A_{t},\pi_{t},d_{t})\big]^{y_{t}}\,\big[1-S^{\theta}(A_{t},\pi_{t},d_{t})\big]^{1-y_{t}}.$
Lemma A1 (Bernoulli outcome log–likelihood ratio).
For the two types $C$ and $W$, the outcome component of the log–likelihood ratio satisfies
$\ln\frac{p^{C}(y_{t}\mid d_{t},A_{t},\pi_{t})}{p^{W}(y_{t}\mid d_{t},A_{t},\pi_{t})}=y_{t}\,\ln\frac{S^{C}(A_{t},\pi_{t},d_{t})}{S^{W}(A_{t},\pi_{t},d_{t})}\;+\;(1-y_{t})\,\ln\frac{1-S^{C}(A_{t},\pi_{t},d_{t})}{1-S^{W}(A_{t},\pi_{t},d_{t})}.$
Proof.
By the Bernoulli likelihoods above,
$\frac{p^{C}(y_{t}\mid d_{t},A_{t},\pi_{t})}{p^{W}(y_{t}\mid d_{t},A_{t},\pi_{t})}=\frac{\big[S^{C}\big]^{y_{t}}\big[1-S^{C}\big]^{1-y_{t}}}{\big[S^{W}\big]^{y_{t}}\big[1-S^{W}\big]^{1-y_{t}}},$
where for brevity we write $S^{\theta}=S^{\theta}(A_{t},\pi_{t},d_{t})$. Taking logs,
$\ln\frac{p^{C}}{p^{W}}=y_{t}\big(\ln S^{C}-\ln S^{W}\big)+(1-y_{t})\big(\ln(1-S^{C})-\ln(1-S^{W})\big),$
which is the stated expression. ∎
Remarks.
(i) If $S^{C}(A_{t},\pi_{t},d_{t})=S^{W}(A_{t},\pi_{t},d_{t})$ (e.g., at $d_{t}=\mathrm{AGM}$ when $_{\mathrm{AGM}}\equiv 0$), the outcome term is 0 and outcomes do not update beliefs. (ii) When $S^{C}>S^{W}$ (e.g., at $d_{t}=\mathrm{early}$ under (4) and (3)), a success $y_{t}=1$ raises the posterior odds of type $C$, while a failure $y_{t}=0$ lowers them. (iii) The expected value of the log–likelihood ratio under type $C$ equals the Kullback–Leibler divergence $\mathrm{KL}(\mathrm{Bern}(S^{C})\,\|\,\mathrm{Bern}(S^{W}))\geq 0$, with equality iff $S^{C}=S^{W}$, formalizing informativeness.
A.2 Belief Consistency
Iterating (7) gives
$\operatorname{logit}(\pi_{t})=\operatorname{logit}(\pi_{0})+\sumop\displaylimits_{\tau=0}^{t-1}\Bigg[\ln\frac{p^{C}(d_{\tau}\mid\pi_{\tau},A_{\tau})}{p^{W}(d_{\tau}\mid\pi_{\tau},A_{\tau})}+y_{\tau}\ln\frac{S^{C}}{S^{W}}+(1-y_{\tau})\ln\frac{1-S^{C}}{1-S^{W}}\Bigg].$ (A1)
Assume (i) the prior is non-degenerate, $\pi_{0}\in(0,1)$; (ii) signals are observed each period; (iii) early campaigns – where $S^{C}\neq S^{W}$ – occur infinitely often with positive probability; and (iv) the expected log-likelihood increments are bounded. Then, applying a law of large numbers to the subsequence of informative (early) periods, the cumulative log-likelihood ratio diverges almost surely with sign depending on the true type:
[…]
Hence posterior beliefs are consistent: $\pi_{t}\to 1$ almost surely under the credible type and $\pi_{t}\to 0$ under the weak type. Even though the timing rule is deterministic, the stochastic nature of proposal outcomes $y_{t}$ guarantees that belief trajectories and the transition from AGM to early campaigning remain random and path-dependent.
A.3 Existence and Uniqueness of the Cutoff Policy and Equilibrium
[…]
Step 1: Existence and uniqueness of the value function
[…]
$(\mathcal{T}V)(\pi,A)=\max_{d\in\{\text{early},\text{AGM}\}}\Big\{u(d;\pi,A)+\beta\,\mathbb{E}[V(\pi^{\prime},A^{\prime})\mid\pi,A,d]\Big\}.$ (A2)
Because $u$ is bounded and continuous, $0<\beta<1$, and the transition kernel $(\pi^{\prime},A^{\prime})$ is well defined and independent of $V$, $\mathcal{T}$ is a contraction on $\mathcal{B}$ with modulus $\beta$. Because $u(d;\pi,A)$ is increasing in $(\pi,A)$ by (3), and the transition kernel is monotone in $(\pi,A)$ (awareness increases in $A$, and expected posteriors are nondecreasing in $(\pi,A)$; cf. Step 2), $\mathcal{T}$ is monotone, so the fixed point $V^{\star}$ is increasing in each argument. This establishes existence and uniqueness of the value function.
Step 2: Single-crossing in $\pi$ and monotonicity in AA
Lemma A2 (Single-Crossing and Monotonicity).
[…]
Proof.
Preliminaries. By standard contraction arguments, the value function $V(\pi,A)$ solving the Bellman equation is bounded, continuous, and increasing in each argument. (The one-period utility is increasing in $(\pi,A)$ and $0<\beta<1$.)
(i) Current wedge. The current payoff wedge equals
[…]
By (3), early is strictly increasing in $\pi$ and weakly increasing in $A$. Hence the current wedge is strictly increasing in $\pi$ and weakly increasing in $A$.
(ii) Continuation wedge. Write
$^{\mathrm{cont}}(\pi,A)\;\equiv\;\mathbb{E}\big[V(\pi_{t+1},A_{t+1})\mid\pi,A,\mathrm{early}\big]-\mathbb{E}\big[V(\pi_{t+1},A_{t+1})\mid\pi,A,\mathrm{AGM}\big],$
with $(\pi_{t+1},A_{t+1})$ evolving by (7) and (1).
Awareness transition. Equation (1) implies that $A_{t+1}^{\mathrm{early}}-A_{t+1}^{\mathrm{AGM}}=(1-\rho)\,\big[v(\mathrm{early})-v(\mathrm{AGM})\big]<0$, a constant gap that is independent of both $\pi$ and $A$. As $A$ increases, both $A_{t+1}^{\mathrm{early}}$ and $A_{t+1}^{\mathrm{AGM}}$ rise with the same slope $\rho$, so the awareness gap between timings is unchanged by $A$ or $\pi$. Therefore, any dependence of cont on $(\pi,A)$ must come from the belief channel described next.
Belief transition. In log-odds form,
[…]
$\quad+\;\underbrace{\mathbb{E}[\text{outcome LLR}\mid\mathrm{early}]}_{\geq 0},$
[…]
Using the add–subtract device and using superscripts $E$ and $G$ for variables computed for early and AGM timings respectively,
[…]
Step 3: Existence and shape of the cutoff curve
At this point we are ready to state the following proposition.
Proposition A1 (Existence and Shape of the Cutoff Curve).
For each level of awareness $A\in[0,1]$, there exists a unique continuous cutoff $\bar{\pi}(A)\in(0,1)$ such that the NGO campaigns early if and only if $\pi>\bar{\pi}(A)$. The cutoff is weakly decreasing in $A$.
[…]
The cutoff $\bar{\pi}(A)$ therefore separates timing choices: for $\pi<\bar{\pi}(A)$, both types optimally choose AGM campaigns, while for $\pi>\bar{\pi}(A)$ only the credible type finds early campaigning profitable. ∎
Step 4: Uniqueness of the Markov-Perfect equilibrium
Because the Bellman operator (A2) is a contraction, the value function $V^{\star}$ is unique. The cutoff $\bar{\pi}(A)$ is also unique for each $A$, and the corresponding optimal policy $d^{\star}(\pi,A)$ is unique except on a measure-zero indifference set $\{\pi=\bar{\pi}(A)\}$. Hence the pair $(V^{\star},d^{\star})$ constitutes a unique pure-strategy MPE.
Proposition A2 (Existence and Uniqueness of the Markov-Perfect Equilibrium).
Under the maintained assumptions, the model admits a unique bounded, continuous value function $V^{\star}(\pi,A)$ solving the Bellman equation (A2), and a unique cutoff policy $\bar{\pi}(A)$ characterized by Proposition A1. The resulting pure-strategy Markov-perfect equilibrium is unique up to a measure-zero indifference set.
Conclusion of the proof.
Steps 2–4 jointly establish Proposition 1 in the main text. Lemma A2 delivers single crossing of the action–value difference, ensuring that timing incentives are monotone in credibility and awareness. Proposition A1 then yields a unique cutoff policy $\bar{\pi}(A)$, under which both types pool on AGM campaigns at low beliefs, while only the credible NGO switches to early campaigning once beliefs exceed the cutoff. Finally, Proposition A2 guarantees that this cutoff policy is part of a unique pure-strategy Markov-perfect equilibrium. Together, these results formalize the pooling–separating structure described in Proposition 1.
Appendix B Omitted Figures
Figure B1: How does target characteristics affect the decision to campaign on AGM dates?
Note: The figure reports the estimated coefficients of the regression of an indicator that an NGO campaign was launched on the target firm’s AGM date on the displayed (standardized) variables, and year fixed effects. Error bars shows 90% (shaded) and 95% confidence intervals using standard errors that are clustered by firm.
Figure B2: Increased NGO visibility around an AGM campaign – All sample
(a) Campaign timing and campaigning NGO’s Google Trends scores
(b) Campaign timing and number of target firms’ negative news (standardized)
Note: Each panel plots the estimated $\beta^{AGM}$ coefficients from (10). In Panel (a), the dependent variable is the monthly Google Trends score for the NGO. In Panel (b), the dependent variable is the (normalized) number of monthly negative media coverage received by a target firm. Error bars shows 95% confidence intervals using standard errors that are clustered by NGO.
Figure B3: NGO campaigns and shareholder proposals – All Sample
(a) All NGOs
Notes: Note: This figure plots the estimated coefficients of the regression of a dummy indicating the occurrence of a shareholder proposal on month $t$ at firm $f$, relative to an NGO campaign targeting firm $f$ on the same cause, $k$ months ahead. The regression also includes campaign and month-by-year-by-industry-by-topic fixed effects. Error bars shows 95% confidence intervals using standard errors that are clustered by firm.
Figure B4: NGO campaign timing and donations – All sample
(a) Donations
(b) Operating Margin
Notes: The figure plots the estimated ${\beta}_{\tau}$ from (12). Panel (a) focuses on donations over (lagged) assets, while Panel (b) focuses on donations minus expenditures over (lagged) assets. Error bars shows 95% confidence intervals using standard errors that are clustered by NGO.
Figure B5: Campaign Timing and Changes in E&S Scores – All Sample
Note: The figure plots the estimated coefficients $\beta_{l}$ from equation (15). The dependent variable is the yearly change in a firm’s average E&S scores. Shaded areas show 95% confidence intervals based on firm-level clustered standard errors. Data combine MSCI ESG scores, NGO campaigns from Sigwatch, firm characteristics from Compustat, and AGM dates from ISS.
Figure B6: NGO past campaigns and the timing of current NGO campaigns – Alternative set of Fixed Effects
Note: This figure plots the estimated coefficients from the regressions of the distance between NGO campaigns and the targeted firm’s closest AGM (in months) on dummies indicating the number of campaigns the focal NGO had already launched. The regression includes replaces NGO-by-year fixed effects with separate fixed effects for each NGO and each year. Hence the coefficients $\{\hat{\beta}_{\tau}\}_{\tau}$ in (16) is also identified from variation in the numbers of NGO campaigns over years. Standard errors are clustered at the NGO level.
Figure B7: Campaign timing for NGOs whose first campaign is after the AGM
Note: This figure plots the estimated coefficients from the regressions (16) of the distance between NGO campaigns and the targeted firm’s closest AGM (in months) on dummies indicating the number of past campaigns of the focal NGO. The panel focuses on NGOs whose first campaign was within the two weeks following the AGM. Standard errors are clustered at the NGO level.
Figure B8: Past campaigns and target firm’s size
Note: This figure plots the estimated coefficients from regressing the target firm’s market cap (in deciles to control for the fact that the tails of the market cap distribution have thickened in recent years) on dummies indicating the number of campaigns the focal NGO had launched in previous years as in (16). Standard errors are clustered at the firm level.
Figure B9: Past campaigns and target firm’s E&S characteristics
(a) Target firm’s E score
(b) Target firm’s S score
Note: This figure plots the estimated coefficients from regressing either a target firm’s environmental score (top panel) or its social score (bottom panel) on dummies indicating either the number of environmental (top panel) or social (bottom panel) campaigns the focal NGO had launched in previous years as in (16). Standard errors are clustered at the firm level.
Appendix C Robustness of Google Trends Results
This section discusses the robustness of the results presented in Figure 3 to alternative specifications. In particular, we replicate the analysis using (i) a staggered difference-in-differences estimator (sun2021estimating), which accounts for heterogeneous campaign timing across NGOs, and (ii) a distributed lags specification, which controls for the possibility of overlapping campaigns within the same year.
C.1 Staggered Difference-in-Differences Specification
We first replicate Figure 3, applying the staggered difference-in-differences approach to account for heterogeneous campaign timing across NGOs (sun2021estimating).
Estimated coefficients $\beta_{\tau}^{AGM}$ are reported in Figure C1. The results are consistent with our baseline findings: AGM-timed campaigns (blue dots) generate sizeable visibility gains for young NGOs (first four campaigns), raising Google search intensity by about 58%. By contrast, NGOs see a more modest increase (33%) in search intensity for later campaigns, and only in the month of the campaign.
Figure C1: Staggered difference-in-differences specification
Note: The Figure plots the estimated $\beta^{AGM}$ coefficients from (10) using a staggered difference-in-differences estimator (sun2021estimating). The dependent variable is the monthly Google Trends score for the NGO. Blue dots are estimated by restricting the sample to an NGO’s first three campaigns, while red triangles consider later campaigns. Standard errors are clustered at the campaign level.
C.2 Distributed Lags Specification
In the baseline event-study, if an NGO launches several campaigns in close succession, their effects may overlap in calendar time. For example, if one campaign takes place at $t=0$ and another at $t=2$, the months around $t=2$ capture the lingering impact of both events. This overlap can bias the estimated effects in the previous analyses. To address this, we estimate a distributed lags specification that explicitly includes leads and lags of both AGM and non-AGM campaigns:
$\text{Google Trends Score}_{g,m,y}$ $=\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-5\end{subarray}}\left(\beta_{\tau}^{AGM}\cdot\text{AGM Campaign}_{g,m+\tau}\right)$ (C1)
$\quad+\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-5\end{subarray}}\left(\beta_{\tau}^{Non-AGM}\cdot\text{Non-AGM Campaign}_{g,m+\tau}\right)$
$\quad+\alpha_{m,y}+\alpha_{g,y}+u_{g,m,y},$
By doing so, the regression attributes observed variation in search activity in each month to the full sequence of campaigns occurring in that window, rather than assigning the entire effect to whichever campaign happens to be closest in event time. Put differently, the distributed lags approach disentangles the impact of overlapping campaigns by allowing each campaign to contribute separately to visibility over several months. This ensures that the estimated $\beta^{AGM}_{\tau}$ captures the marginal visibility effect of a single campaign, net of any other campaigns that may have occurred nearby.
Figure C2: Distributed lags specification
(a) First three campaigns
(b) Later campaigns
Note: The two plots present the estimated $\beta^{AGM}$ with blue circles and $\beta^{Non-AGM}$ with red triangles coefficients from (C1) on two different samples. The first panel subsets the data to each NGO’s first three campaigns, while the second panel examines later campaings. The dependent variable is the monthly Google Trends score for the NGO. Standard errors are clustered at the campaign level.
Figure C2 reports the estimated coefficients. When restricting attention to an NGO’s first three campaigns, the results confirm that AGM-timed campaigns lead to significant short-run increases in visibility (blue dots), while later campaigns generate little or no effect. Whether early or late, non-AGM campaings have a limited effect on an NGO’s visibility.
Overall, these robustness checks confirm that our main results are not driven by campaign timing heterogeneity across NGOs or by overlapping campaigns within a year.
Appendix D Matching Proposals And NGO Campaigns
This section outlines the process of matching each shareholder proposal in the ISS dataset with a relevant NGO campaign cause.
Sigwatch assigns a“Generic Topic” to each campaign in their dataset. To streamline the analysis, we combine certain causes when they are difficult to distinguish or when one category has few campaigns and can reasonably fit within a broader category. Specifically, we merge “Labor Rights” with “Human & Labor Rights,” “Children” with ”Human & Labor Rights,” “CSR / ESG” with “CSR,” and “Biodiversity” and “Wildlife” with ”Environment.” For campaigns under the “Safety” cause, which includes only 12 entries, we examine each campaign’s specific purpose and classify it as either “Product Safety” or “Food Safety.”
Using these consolidated “Generic Cause Names”, we then categorize each shareholder proposal based on its short description. When a shareholder proposal is submitted, it includes a brief, one-sentence description, which is provided in the ISS dataset and serves as the foundation for cause classification. We start by assigning proposals to causes based on keywords in their descriptions and then manually classify any remaining proposals.
Table D1: Proposal Causes
NGO Campaign Category Example of Shareholder Proposals
Topic Classification (E or S)
Climate Change E Encourage and Support Carbon Tax to Address Climate Change
Environment E Stop financing mountain top removal coal mining
Pollution E Report on Efforts to Reduce Pollution From Products and Operations
Rainforest E Report on impact on rainforest
Sustainability E Establish board committee on sustainability
Waste Generation E Do not increase radioactive waste production
Water E Report on supply chain water impacts
Animal Rights & Welfare S Review suppliers’ animal slaughter methods
CSR S Disclose Charitable Contributions
Food Safety S Label Genetically Modified Ingredients
Health S Reduce Tobacco Harm to Health
Human & Labor Rights S End Sri Lanka Trade Partnerships Until Human Rights Improve
Indigenous People S Adopt Global Policy Regarding the Rights of Indigenous People
Intellectual Property Rights S Implement a Process to Fully Vet Whether a
Supervisor is a Bona Fide Inventor Before a Patent is Filed
Politics S Affirm Political Nonpartisanship
Product Safety S Report on Nanomaterial Product Safety
Worker Safety S Report on Reducing Risks of Accidents
  • Note: This Table reports the 17 topics assigned to each shareholder proposal in the sample. Shareholder proposals data from ISS. The topics are from Sigwatch’s categorization of campaigns.